### v2

# Estimating the FDR of variable selection

Abstract
We introduce a generic estimator for the false discovery rate of any model selection procedure, in common statistical modeling settings including the Gaussian linear model, Gaussian graphical model, and model-X setting. We prove that our method has a conservative (non-negative) bias in finite samples under standard statistical assumptions, and provide a bootstrap method for assessing its standard error. For methods like the Lasso, forward-stepwise regression, and the graphical Lasso, our estimator serves as a valuable companion to cross-validation, illuminating the tradeoff between prediction error and variable selection accuracy as a function of the model complexity parameter.
1 Introduction
When deciding which variables to include in a statistical model, a data analyst commonly wants to balance multiple goals including predictive accuracy, accurate detection of signal variables, parsimony, and interpretability. Perhaps the best-known paradigm is to use a variable-selecting method such as the Lasso (Tibshirani, 1996), forward stepwise regression (Hocking, 1976), or the graphical Lasso (Friedman et al., 2008), with a tuning parameter that controls the model complexity, and select the value of the tuning parameter by minimizing the cross-validation (CV) estimate of out-of-sample prediction error. This paradigm is versatile but prioritizes prediction error above all else, while ignoring the other goals mentioned above. In particular, CV can be quite liberal in including irrelevant explanatory variables (see e.g. Baumann, 2003), an undesirable feature in applications where accurate variable selection is a priority.
In such applications, we could abandon estimation methods like the Lasso and instead apply multiple testing methods such as knockoffs (Barber and Candès, 2016; Candès et al., 2018), or the Benjamini–Hochberg procedure or its dependence-corrected variants (Benjamini and Hochberg, 1995; Benjamini and Yekutieli, 2001; Fithian and Lei, 2022), which guarantee control of the false discovery rate (FDR) at a user-specified level. However, multiple testing methods do not yield fitted models, and the model defined by including only the selected variables can have very poor fit to the data.
In this work we introduce a new approach for assessing variable selection accuracy of estimation methods like the Lasso: an estimator for the FDR of any variable selection procedure, which can be evaluated alongside the cross-validation curve to inform the analyst about the tradeoff between variable selection accuracy and prediction accuracy over the entire regularization path. Our estimator can be applied in common statistical settings including the Gaussian linear model, the Gaussian graphical model, and the nonparametric model-X framework of Candès et al. (2018), requiring no additional statistical assumptions in these settings beyond those required for standard hypothesis tests. Its bias is non-negative in finite samples and typically small, and its standard error can be assessed using the bootstrap. In the special case of thresholding independent $p$-values for multiple hypothesis tests, our method is almost identical to the FDR estimation method advocated by Storey (2002), but slightly less conservative.
(a) Easy variable selection scenario
(b) Difficult variable selection scenario
Figure 1: Cross-validation MSE (blue), true FDR (black), and our FDR estimator (red), for Lasso regression in two scenarios with explanatory variables that are (a) independent, and (b) highly correlated. In scenario (a) the minimum-CV model (solid blue vertical line) has high FDR, while the one-standard-error rule (dashed vertical line) achieves a reasonably low FDR. In scenario (b) there is no model that simultaneously achieves good predictive performance and low FDR. In both scenarios, our FDR estimator successfully captures important information about variable selection performance that is not evident from the CV curve.
Figure 1 illustrates why our method is useful by contrasting two Lasso regression scenarios with very different tradeoffs between prediction error and FDR. In both scenarios, we generate $n=600$ data points from a Gaussian linear model with $d=200$ standard Gaussian explanatory variables, of which only $d_{1}=20$ are true signal variables with the same moderate signal strength. The only difference is that the explanatory variables are independent in the first scenario, but correlated in the second (specifically, they follow $\text{AR}(0.8)$ correlation). This difference dramatically changes the variable selection performance: When the explanatory variables are independent, it is possible to select a good predictive model if we use the one-standard-error rule (but not if we minimize CV error). When the explanatory variables are correlated, however, all models with reasonably good predictive error have FDR exceeding $40\%$. While the CV curve gives no indication of the variable selection performance in either scenario, our estimator is accurate in both. The next section explains how we calculate it.
1.1 FDR estimation in the Gaussian linear model
Although our method applies in a more general setting, the special case of the Gaussian linear model with fixed design provides a concrete and familiar context in which to introduce our key ideas. In this model, we assume
$\boldsymbol{Y}=\sum_{j=1}^{d}\boldsymbol{X}_{j}\theta_{j}+\boldsymbol{\varepsilon}=\boldsymbol{X}\boldsymbol{\theta}+\boldsymbol{\varepsilon},\quad\boldsymbol{\varepsilon}\sim\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I}_{n}),$ (1)
where $\boldsymbol{X}_{j}=(x_{1,j},\ldots,x_{n,j})^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{Y}=(y_{1},\ldots,y_{n})^{\mkern-1.5mu\mathsf{T}}$ denote the vectors of $n$ observed values of the explanatory variables $X_{j}$ and outcome variable $Y$, respectively, and $\boldsymbol{X}$ denotes the $n$-by-$d$ matrix of all observed explanatory variables. We call $X_{j}$ a noise variable if the hypothesis $H_{j}:\;\theta_{j}=0$ is true and a signal variable otherwise. $\boldsymbol{X}$ is assumed to be fixed with $d<n$, while $\boldsymbol{\theta}\in\mathbb{R}^{d}$ and $\sigma^{2}>0$ are assumed to be unknown.
Any variable selection procedure yields an index set $\mathcal{R}$, which we can regard as an estimator for the set of signal variables. For example, the Lasso estimator for a fixed regularization parameter $\lambda>0$ selects $\mathcal{R}(\lambda)=\left\{j:\;\widehat{\boldsymbol{\theta}}^{\textnormal{Lasso}}_{j}(\lambda)\neq 0\right\},$ where
$\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}}(\lambda)=\operatorname*{argmin}_{\boldsymbol{\theta}}\;\frac{1}{2}\,\left\|\boldsymbol{Y}-\boldsymbol{X}\boldsymbol{\theta}\right\|_{2}^{2}+\lambda\left\|\boldsymbol{\theta}\right\|_{1}.$ (2)
By contrast, forward stepwise regression (FS) selects variables by starting from the null model and adding one variable at a time, greedily minimizing at each step the next model’s residual sum of squares. After $k=1,2,\ldots$ steps, the estimator $\boldsymbol{\widehat{\theta}^{\textnormal{FS}}}(k)$ is given by estimating an ordinary least squares (OLS) regression for the $k$ selected variables.
Benjamini and Hochberg (1995) define the false discovery rate (FDR) as the expectation of the false discovery proportion (FDP), defined here as the fraction of selected variables that are actually irrelevant. If $\mathcal{H}_{0}$ is the index set of all noise variables and $R=|\mathcal{R}|$ is the total number of selections, we have
$\textnormal{FDR}=\mathbb{E}[\textnormal{FDP}]\quad\textnormal{and}\quad\textnormal{FDP}=\frac{|\mathcal{R}\cap\mathcal{H}_{0}|}{R},$ (3)
using the convention that $/0=0$ 0; i.e., $\textnormal{FDP}=0$ when no variables are selected.
Note that we can linearly decompose the FDR as
$\textnormal{FDR}=\sum_{j=1}^{d}\textnormal{FDR}_{j},\quad\textnormal{FDR}_{j}=\mathbb{E}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{|\mathcal{R}|}\right]\cdot\mathbf{1}\left\{j\in\mathcal{H}_{0}\right\}.$ (4)
Using this decomposition, a promising estimator of the FDR is
$\widehat{\textnormal{FDR}}=\sum_{j=1}^{d}\widehat{\textnormal{FDR}}_{j},\quad\widehat{\textnormal{FDR}}_{j}=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{|\mathcal{R}|}\mid\boldsymbol{S}_{j}\right]\cdot\frac{\mathbf{1}\left\{p_{j}>0.1\right\}}{0.9},$ (5)
where $p_{j}$ is the $p$-value for the usual two-sided $t$-test, and $\boldsymbol{S}_{j}=(\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\,\left\|\boldsymbol{Y}\right\|^{2})$ is a sufficient statistic under the null model $H_{j}$. Under $H_{j}$, the first factor in $\widehat{\textnormal{FDR}}_{j}$ is the uniform minimum-variance unbiased (UMVU) estimator of $\textnormal{FDR}_{j}$, obtained by Rao-Blackwellization, while the second factor attempts to zero out non-null terms.
Because $p_{j}$ and $\boldsymbol{S}_{j}$ are independent under $H_{j}$, the product $\widehat{\textnormal{FDR}}_{j}$ is an unbiased estimator for its counterpart $\textnormal{FDR}_{j}$ whenever $j\in\mathcal{H}_{0}$; otherwise $\widehat{\textnormal{FDR}}_{j}\geq\textnormal{FDR}_{j}=0$. As a direct consequence, we have
$\mathbb{E}[\widehat{\textnormal{FDR}}]\geq\textnormal{FDR}.$ (6)
The bias of $\mathbb{E}[\widehat{\textnormal{FDR}}]$ is small if most signal variables have $p_{j}\leq 0.1$ with high probability.
The definition of our estimator $\widehat{\textnormal{FDR}}$ in (5) is highly general and makes no assumptions about the form of the variable selection procedure $\mathcal{R}$. As a result, it applies to any selection procedure including the Lasso, forward stepwise regression, LARS (Efron et al., 2004), the elastic net (Zou and Hastie, 2005), and many others.
Our method requires no further assumptions beyond the standard linear model assumptions in (1) required to justify the marginal $t$-tests for each variable. Indeed, in the common case where the method $\mathcal{R}$ is a function of $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}$ and $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ only (which is true of all methods referenced above), the only statistical assumption required to justify the inequality (6) is that the $t$-statistic follows a $t_{n-d}$ distribution under $H_{j}$, conditional on $\boldsymbol{S}_{j}$. This assumption may hold approximately even when the parametric assumptions of the linear model fail.
As we see next, our method generalizes far beyond the Gaussian linear model to other parametric and nonparametric contexts, provided that our modeling assumptions supply an appropriate sufficient statistic $\boldsymbol{S}_{j}$ for each null hypothesis $H_{j}$. Section 2 defines our method generically and discusses additional applications including the graphical Lasso (Friedman et al., 2008) and nonparametric model-X assumptions introduced by (Candès et al., 2018). Section 3 discusses theoretical analysis of the standard error in stylized settings, as well as our bootstrap scheme for estimating the standard error, which is used to generate the error bars in all figures, and Sections 4 and 5 illustrate our method’s behavior in real data and simulation examples, respectively. Section 6 discusses computational issues, and Section 7 concludes.
2 General formulation of our estimator
2.1 Notation and statistical assumptions
Throughout the paper, we will use upper-case letters to represent variables and lower-case letters to denote individual observations. For example, $y_{i}$ is the $i$th observation of variable $Y$ and $x_{i,j}$ is the $i$th observation of variable $X_{j}$. We use the upper-case bold-face letter to denote a vector of all $n$ observations, e.g. $\boldsymbol{X}_{j}=(x_{1,j},\ldots,x_{n,j})^{\mkern-1.5mu\mathsf{T}}$. We use $[d]$ to denote the set $\{1,\ldots,d\}$. For $B\subseteq[d]$, we denote as $\boldsymbol{X}_{B}$ the block of matrix $\boldsymbol{X}$ including only the variables in $B$, and we use $\boldsymbol{X}_{-j}$ as a shorthand for $\boldsymbol{X}_{[d]\setminus\{j\}}$ and $\boldsymbol{X}_{-B}$ for $\boldsymbol{X}_{[d]\setminus B}$.
We denote by $\boldsymbol{D}$ a generic random object that encodes the entire data set, whose distribution $P$ we will assume lies in some known family $\mathcal{P}$. If $\mathcal{P}$ is a parametric family, we generically use $\boldsymbol{\theta}$ to denote its parameters, and $P_{\boldsymbol{\theta}}$ to denote the corresponding distribution. In the Gaussian linear model in Section 1.1, irrelevance of explanatory variable $X_{j}$ is represented by the null hypothesis $H_{j}:\boldsymbol{\theta}_{j}=0$, but other problem settings will involve other null hypotheses.
To define our estimator in a generic modeling context and prove that its bias is non-negative, the sole statistical assumption we will rely on is the availability of a sufficient statistic $\boldsymbol{S}_{j}$ for the submodel corresponding to each null hypothesis $H_{j}$. In one sense this assumption is vacuous, because $\boldsymbol{D}$ itself is a sufficient statistic for any model, but we will only obtain a useful estimator if $\boldsymbol{S}_{j}$ encodes less information than $\boldsymbol{D}$. While not strictly required, it is also convenient to assume that there is a standard choice of $p$-value $p_{j}$ for each $H_{j}$, which is valid conditional on $\boldsymbol{S}_{j}$.
Section 2.4 discusses appropriate choices of $\boldsymbol{S}_{j}$ and $p_{j}$ in two additional modeling settings where our method can be applied: the nonparametric model-X setting, and the Gaussian graphical model. In both examples, as in the Gaussian linear modeling example, $\boldsymbol{S}_{j}$ is a complete sufficient statistic for $H_{j}$ and there exists a natural choice of $p_{j}$ which is independent of $\boldsymbol{S}_{j}$ under $H_{j}$.
2.2 FDR estimation
We now present a generic construction of our $\widehat{\textnormal{FDR}}$, assuming only that a sufficient statistic $\boldsymbol{S}_{j}$ is available for each $H_{j}$. Recall the decomposition (4):
$\textnormal{FDR}=\sum_{j=1}^{d}\textnormal{FDR}_{j},\quad\text{where }\textnormal{FDR}_{j}=\mathbb{E}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\right]\cdot\mathbf{1}\left\{j\in\mathcal{H}_{0}\right\}.$
To estimate each variable’s contribution $\textnormal{FDR}_{j}$, we proceed by estimating each of the two factors in turn.
Step 1: Estimating the first factor:
Crucially, note that the factor $\mathbb{E}[\mathbf{1}\{j\in\mathcal{R}\}/R]$ is only relevant when $j\in\mathcal{H}_{0}$. As a result, we can assume when estimating it that $H_{j}$ holds. Since $\boldsymbol{S}_{j}$ is sufficient for the submodel $H_{j}$, we can do so by Rao-Blackwellizing the integrand (Rao, 1945; Blackwell, 1947), obtaining
$\widehat{\textnormal{FDR}}^{*}_{j}(\boldsymbol{S}_{j}):=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right].$ (7)
For $j\in\mathcal{H}_{0}$, $\widehat{\textnormal{FDR}}_{j}^{*}$ is unbiased by construction, and it is the UMVU estimator for the first factor of $\textnormal{FDR}_{j}$ whenever $\boldsymbol{S}_{j}$ is complete (Lehmann and Scheffé, 1950).
Step 2: Estimating the second factor:
We could obtain a conservative estimator by simply upper-bounding the indicator $\mathbf{1}\{j\in\mathcal{H}_{0}\}\leq 1$, resulting in the estimator $\widehat{\textnormal{FDR}}^{*}=\sum_{j}\widehat{\textnormal{FDR}}_{j}^{*}$. However this estimator’s bias, while still non-negative, could be quite large. In particular, if some of the signal variables are selected with high conditional probability given $\boldsymbol{S}_{j}$, their excess contributions may be large.
A better estimator of the indicator $\mathbf{1}\{j\in\mathcal{H}_{0}\}$ should reliably eliminate terms for strong signal variables without introducing bias in the terms for noise variables. For any estimator $\psi_{j}(\boldsymbol{D})\in[0,1]$, define its normalized version:
$\phi_{j}(\boldsymbol{D})=\frac{\psi_{j}(\boldsymbol{D})}{\mathbb{E}_{H_{j}}[\psi_{j}(\boldsymbol{D})\mid\boldsymbol{S}_{j}]},$ (8)
or $\phi_{j}(\boldsymbol{D})=1$ if $\mathbb{E}_{H_{j}}[\psi_{j}|\boldsymbol{S}_{j}]=0$. Note we have $\mathbb{E}_{H_{j}}[\phi_{j}\mid\boldsymbol{S}_{j}]=1$ almost surely.
Commonly there is a standard test for $H_{j}$, whose $p$-value $p_{j}$ is uniform and independent of $\boldsymbol{S}_{j}$ under $H_{j}$. Then, we can take $\psi_{j}=\mathbf{1}\{p_{j}>\zeta\}$ for $\zeta\in(0,1)$, and
$\phi_{j}(p_{j})=\frac{\mathbf{1}\{p_{j}>\zeta\}}{1-\zeta}.$ (9)
If the signal is strong enough for the test to reject $H_{j}$ with high probability at level $\zeta$, then variable $j$’s contribution will almost always be zero. We adopt (9) as the canonical form of $\phi_{j}$, using $\zeta=0.1$ for all simulations.
Combining (7) and (8), we obtain the estimator
$\widehat{\textnormal{FDR}}=\sum_{j=1}^{d}\widehat{\textnormal{FDR}}_{j},\quad\text{where }\widehat{\textnormal{FDR}}_{j}(\boldsymbol{D})=\widehat{\textnormal{FDR}}_{j}^{*}(\boldsymbol{S}_{j})\cdot\phi_{j}(\boldsymbol{D}).$ (10)
We can write the canonical version of $\widehat{\textnormal{FDR}}_{j}$ more concretely as
$\widehat{\textnormal{FDR}}_{j}(\boldsymbol{D})=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]\cdot\frac{\mathbf{1}\{p_{j}>\zeta\}}{1-\zeta}.$ (11)
These two factors are independent if $p_{j}\perp\!\!\!\perp\boldsymbol{S}_{j}$, as is commonly the case.
The estimator derived above, in its generic form (10), is always conservatively biased:
Theorem 2.1.
For any selection procedure $\mathcal{R}$, and for any estimator $\psi_{j}:\;\boldsymbol{D}\to[0,1]$ of $\mathbf{1}\{j\in\mathcal{H}_{0}\}$, the estimator $\widehat{\textnormal{FDR}}$ defined in (710) has non-negative bias:
$\mathbb{E}[\widehat{\textnormal{FDR}}]\geq\textnormal{FDR}.$
Proof.
It suffices to show $\mathbb{E}[\widehat{\textnormal{FDR}}_{j}]\geq\textnormal{FDR}_{j}$ for all $j$. For $j\in\mathcal{H}_{0}$, $\widehat{\textnormal{FDR}}_{j}$ is unbiased:
$\mathbb{E}[\widehat{\textnormal{FDR}}_{j}]$ $=\mathbb{E}\left[\mathbb{E}_{H_{j}}\left[\widehat{\textnormal{FDR}}_{j}^{*}(\boldsymbol{S}_{j})\cdot\phi_{j}(\boldsymbol{D})\mid\boldsymbol{S}_{j}\right]\right]$
$=\mathbb{E}\left[\widehat{\textnormal{FDR}}_{j}^{*}(\boldsymbol{S}_{j})\cdot\mathbb{E}_{H_{j}}[\phi_{j}(\boldsymbol{D})\mid\boldsymbol{S}_{j}]\right]$
$=\mathbb{E}\left[\widehat{\textnormal{FDR}}_{j}^{*}(\boldsymbol{S}_{j})\right]$
$=\textnormal{FDR}_{j}.$
For $j\not\in\mathcal{H}_{0}$, $\widehat{\textnormal{FDR}}_{j}\geq\textnormal{FDR}_{j}=0$ almost surely. ∎
Remark 2.1.
While we might have initially held out hope for an unbiased estimator, attaining one is typically impossible in practice. We can see this in the Gaussian linear model (1), where the FDR of any nontrivial procedure $\mathcal{R}$ is bounded and discontinuous in $\boldsymbol{\theta}$, but the distribution of the response $\boldsymbol{Y}$ depends on $\boldsymbol{\theta}$ in a continuous way. As a result, no statistic $T(\boldsymbol{Y})$ can have $\mathbb{E}_{\boldsymbol{\theta}}[T(\boldsymbol{Y})]=\textnormal{FDR}_{\boldsymbol{\theta}}(\mathcal{R})$ for all $\boldsymbol{\theta}$.
Remark 2.2.
Extending the previous remark, we might view the smoothness of $\mathbb{E}[\widehat{\textnormal{FDR}}]$ with respect to $\boldsymbol{\theta}$, and the corresponding conservatism of $\widehat{\textnormal{FDR}}$, as a robustness feature of our estimator, if we have conceptual concerns about the use of point nulls $H_{j}:\;\theta_{j}=0$ in defining the FDR. If many of the $\theta_{j}$ parameters are minuscule, but none are exactly zero, then the formal FDR of any procedure is exactly zero, but a more practical perspective would consider these “near-nulls” to be false discoveries whenever they are selected. Because the distribution of $\widehat{\textnormal{FDR}}$ is continuous in $\boldsymbol{\theta}$, it effectively treats the near-nulls as nulls. That is, while it may be a severely biased estimator for the formal FDR (which is zero), it may be nearly unbiased for the “practical FDR” that counts near-null selections as Type I errors.
Remark 2.3.
As discussed in Section 2.1, our sole statistical assumption — availability of a sufficient statistic $\boldsymbol{S}_{j}$ for the submodel defined by $H_{j}$ — is formally vacuous, because $\boldsymbol{D}$ is itself a sufficient statistic. However, if we take $\boldsymbol{S}_{j}=\boldsymbol{D}$, we have
$\widehat{\textnormal{FDR}}_{j}(\boldsymbol{D})=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{D}\right]\cdot\frac{\psi_{j}(\boldsymbol{D})}{\mathbb{E}_{H_{j}}[\psi_{j}(\boldsymbol{D})\mid\boldsymbol{D}]}=\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R},$
and consequently $\widehat{\textnormal{FDR}}=1$ almost surely. Theorem 2.1 still applies to this estimator, which trivially has non-negative bias for the true FDR, but the estimator is useless.
Remark 2.4.
For variable selection by thresholding independent p-values $p_{j}$ at threshold $c\in(0,1)$, Storey (2002) proposed a conservatively biased FDR estimator
$\widehat{\textnormal{FDR}}^{\textnormal{Storey}}=\frac{c}{R(c)}\cdot\frac{\#\{p_{j}>\zeta\}}{1-\zeta},\quad\textnormal{ where }\mathcal{R}(c)=\left\{j:\;p_{j}\leq c\right\}.$
To apply our approach, let $\boldsymbol{S}_{j}=p_{-j}$ and assume $c\leq\zeta$. Then we have
$\widehat{\textnormal{FDR}}=\sum_{j=1}^{d}\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{p_{j}\leq c\right\}}{\#\left\{j:\;p_{j}\leq c\right\}}\mid p_{-j}\right]\cdot\frac{\mathbf{1}\{p_{j}>\zeta\}}{1-\zeta}=\frac{c}{R(c)+1}\cdot\frac{\#\{p_{j}>\zeta\}}{1-\zeta}.$
Thus, our $\widehat{\textnormal{FDR}}$ is nearly identical to Storey’s estimator but slightly less conservative.
The next section elaborates on our method’s definition to help readers build intuition for how it works.
2.3 Understanding our method
To better understand our method, it is helpful to consider a simpler but closely related method that we could use to estimate a different Type I error rate, called the per-family error rate (PFER)
PFER $=\mathbb{E}\left[|\mathcal{R}\cap\mathcal{H}_{0}|\right]$
$=\sum_{j=1}^{d}\mathbb{P}\left[j\in\mathcal{R}\right]\cdot\mathbf{1}\left\{j\in\mathcal{H}_{0}\right\}$
$=\sum_{j\in\mathcal{H}_{0}}\mathbb{P}\left[j\in\mathcal{R}\right].$
We can follow the same approach as in Section 2.2 to obtain the PFER estimator
$\widehat{\textnormal{PFER}}$ $=\sum_{j=1}^{d}\mathbb{P}_{H_{j}}\left[j\in\mathcal{R}\mid\boldsymbol{S}_{j}\right]\cdot\frac{\mathbf{1}\{p_{j}>\zeta\}}{1-\zeta}$
$=\frac{1}{1-\zeta}\cdot\sum_{j:\,p_{j}>\zeta}\mathbb{P}_{H_{j}}\left[j\in\mathcal{R}\mid\boldsymbol{S}_{j}\right].$ (12)
The $j$th term in the sum (12) is an estimate of variable $j$’s chance of being selected, calculated under the assumption that $H_{j}$ holds. The estimator $\widehat{\textnormal{PFER}}$ adds up these estimated selection probabilities over all variables with unimpressive $p$-values, then multiplies the sum by an inflation factor $(1-\zeta)^{-1}$ to account for the null terms that were eliminated by chance. If $\zeta=0.1$, the inflation factor is $10/9\approx 1.1$.
In problems where the number $R$ of selections is relatively large and stable, such that $1/R$ is roughly constant conditional on $\boldsymbol{S}_{j}$, we have
$\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]\approx\frac{\mathbb{P}_{H_{j}}\left[j\in\mathcal{R}\mid\boldsymbol{S}_{j}\right]}{R},$
and as a result
$\widehat{\textnormal{FDR}}\approx\frac{\widehat{\textnormal{PFER}}}{R}=\frac{1}{(1-\zeta)R}\cdot\sum_{j:\,p_{j}>\zeta}\mathbb{P}_{H_{j}}\left[j\in\mathcal{R}\mid\boldsymbol{S}_{j}\right].$ (13)
For variable selection methods where $R$ is deterministic, including forward stepwise regression with a fixed number of steps, this approximation is exact.
This approximation aids our intuition by telling us that $\widehat{\textnormal{FDR}}$ is well-approximated by a sum of estimated inclusion probabilities, divided by the number $R$ of selected variables, and inflated by $(1-\zeta)^{-1}$. Roughly speaking, $\widehat{\textnormal{FDR}}$ is large when some unimpressive variables stood a good chance of being selected (or actually were selected), and otherwise $\widehat{\textnormal{FDR}}$ is small.
Remark 2.5.
The PFER is interesting in its own right as an upper bound for the family-wise error rate (FWER), defined as
$\textnormal{FWER}=\mathbb{P}\left[|\mathcal{R}\cap\mathcal{H}_{0}|\geq 1\right]\leq\textnormal{PFER}.$
As a result, the estimator $\widehat{\textnormal{PFER}}$ in (12) can also be understood as a conservatively biased estimator for FWER. We leave exploration of this extension to future work.
2.4 Example model assumptions
In this section we give additional examples beyond the Gaussian linear model and demonstrate the choices of $\boldsymbol{S}_{j}$ and the ways to compute a valid $p$-value.
Example 2.1 (Nonparametric “model-X” setting).
Candès et al. (2018) introduced novel nonparametric modeling assumptions for multiple testing in supervised learning problems, which allow the analyst to test for conditional independence between the response and each explanatory variable, given the other variables. Assume that we observe $n$ independent realizations of the $(d+1)$-variate random vector $(X,Y)\sim P$, where the joint distribution $P_{X}$ of explanatory variables is fully known, but nothing is assumed about the conditional distribution $P_{Y\mid X}$. [^4] These assumptions can be relaxed to allow for $P_{X}$ to be known up to some unknown parameters; see Huang and Janson (2020).
Under these modeling assumptions, $\boldsymbol{S}_{j}=(\boldsymbol{X}_{-j},\boldsymbol{Y})$ is a complete sufficient statistic for the $j$th null hypothesis $H_{j}:\;Y\perp\!\!\!\perp X_{j}\mid X_{-j}$. Under $H_{j}$, after fixing the observed $\boldsymbol{S}_{j}$ we can sample from the conditional distribution of $\boldsymbol{D}$ by sampling a new $\boldsymbol{X}_{j}$ vector according to its known distribution given $\boldsymbol{X}_{-j}$. For any test statistic $T_{j}(\boldsymbol{D})$, we can compare the observed value to its simulated distribution to obtain a $p$-value that is independent of $\boldsymbol{S}_{j}$; the corresponding test is called the conditional randomization test (CRT). In our simulations, we use the marginal covariance $\textnormal{cov}(\boldsymbol{X}_{j},\boldsymbol{Y})$ as our CRT test statistic and perform a two-sided test, but other choices may be preferred depending on the problem setting and the manner in which the analyst expects $Y$ to depend on the signal variables.
Example 2.2 (Gaussian graphical model).
Assume we observe $n$ independent realizations of a $d$-variate Gaussian random vector $X$ ($n>d$), with mean zero and positive-definite covariance matrix $\boldsymbol{\Sigma}$. In the corresponding precision matrix $\boldsymbol{\Theta}=\boldsymbol{\Sigma}^{-1}$, the entry $\boldsymbol{\Theta}_{jk}$ is zero if and only if $X_{j}$ and $X_{k}$ are conditionally independent given the other entries. In other words, the sparsity pattern of the undirected dependence graph $\mathcal{G}$ for $X$ is identical to the sparsity pattern of $\boldsymbol{\Theta}$. In lieu of estimating all $d(d-1)/2$ off-diagonal entries of $\boldsymbol{\Theta}$, a common modeling assumption is that the dependence graph is sparse, and by extension that $\boldsymbol{\Theta}$ is sparse as well (Meinshausen and Bühlmann, 2006; Yuan and Lin, 2007; Friedman et al., 2008).
We associate to each pair $j\neq k$ the null hypothesis that $(j,k)$ is absent from the true dependence graph $\mathcal{G}$, or equivalently $H_{jk}:\;\boldsymbol{\Theta}_{jk}=0$. The Gaussian graphical model is a parametric exponential family model whose natural parameters are the entries of $\boldsymbol{\Theta}$, with corresponding sufficient statistics given by the entries of $\widehat{\boldsymbol{\Sigma}}={\boldsymbol{X}}^{\mkern-1.5mu\mathsf{T}}{\boldsymbol{X}}$, which follows a Wishart distribution. As a result, a complete sufficient statistic for $H_{jk}$ is given by observing every entry of $\widehat{\boldsymbol{\Sigma}}$ except $\widehat{\boldsymbol{\Sigma}}_{jk}$ and $\widehat{\boldsymbol{\Sigma}}_{kj}$, or equivalently by
$\boldsymbol{S}_{jk}=\left(\boldsymbol{X}_{-k}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{-k},\;\boldsymbol{X}_{-\{j,k\}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{k},\;\|\boldsymbol{X}_{k}\|^{2}\right).$
This sufficient statistic is strongly reminiscent of the sufficient statistic for $H_{j}$ in the Gaussian linear model, with $\boldsymbol{X}_{-k}$ playing the role of the design matrix and $\boldsymbol{X}_{k}$ playing the role of the response, and the standard test for $H_{jk}$ is indeed identical to the $t$-test for $H_{j}$ in that regression problem. In our simulations, we use the corresponding two-sided $t$-test $p$-value $p_{jk}$. See Drton and Perlman (2007) for additional details.
In this modeling context, we will study the FDR of the graphical Lasso algorithm of Friedman et al. (2008), which is obtained by maximizing the log-likelihood with a Lasso penalty on the off-diagonal entries of $\boldsymbol{\Theta}$:
$\boldsymbol{\widehat{\Theta}}(\lambda)=\operatorname*{argmax}_{\boldsymbol{\Theta}\succ\boldsymbol{0}}\;\log\det\boldsymbol{\Theta}-\textnormal{trace}(\boldsymbol{\widehat{\Sigma}}\boldsymbol{\Theta})-\lambda\sum_{j<k}|\Theta_{jk}|.$ (14)
3 Standard error of $\widehat{\textnormal{FDR}}$
In this section, we discuss the standard error of our estimator $\widehat{\textnormal{FDR}}$. Since our general method can be adapted for many different selection algorithms in many different statistical modeling contexts, it is very difficult to provide a general analysis that covers all contexts in which our method might be applied, but we can offer theoretical results in a narrower context, as well as a practical estimation strategy that we have found to be reasonably reliable in our simulation studies. Section 3.1 develops a theoretical finite-sample bound for our estimator under stylized assumptions, leading to conditions under which the variance of $\widehat{\textnormal{FDR}}$ is vanishing, while Section 3.2 discusses a specialized bootstrap method for standard error estimation, which we use to produce all figures in this work.
3.1 Theoretical bound for standard error
[…]
Assumption 3.1 (block-orthogonal variables).
Let $\boldsymbol{X}$ be an $n\times d$ matrix, and let $B_{1},\ldots,B_{K}\subseteq[d]$ represent a partition of $[d]$ into $K$ disjoint subsets (that is, $\bigcup_{k}B_{k}=[d]$). We say the matrix $\boldsymbol{X}$ has block orthogonal variables with respect to the partition $B$ if $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{B_{\ell}}=\boldsymbol{0}_{|B_{k}|\times|B_{\ell}|}$, for all $k\neq\ell$.
Assumption 3.1 allows us to treat the Lasso on the entire matrix $\boldsymbol{X}$ as a separable optimization problem on the $K$ different blocks of explanatory variables. When this assumption is in effect, we will denote the size of the largest block as $b=\max_{k}|B_{k}|$. We will assume throughout this section that every explanatory variable has mean zero and unit norm: $\sum_{i=1}^{n}x_{i,j}=0$ and $\|\boldsymbol{X}_{j}\|^{2}=1$, for all $j=1,\ldots,d$. This last assumption comes with little practical loss of generality since conventional implementations of the Lasso begin by normalizing all variables in this way.
For our results below, it will be important to be able to provide a probabilistic lower bound on the average number $q$ of selections per block:
Condition 3.1 (enough selections).
We have $\mathbb{P}(R\leq Kq)\leq\rho$ for some $q,\rho\geq 0$.
Because Condition 3.1 always holds for some $q$ and $\rho$, it has force only after we show that it holds for specific values of $q$ and $\rho$, defined in terms of other problem quantities. Proposition 3.1 gives examples of such bounds for Lasso problems.
Proposition 3.1.
In the Gaussian linear model with block-orthogonal variables (Assumption 3.1), the following statements hold for the Lasso with tuning parameter $\lambda>0$:
  1. 1.
    Let $\delta(z):=2(1-\Phi(z))$ denote the probability that a Gaussian random variable with mean 00 and variance 1 exceeds $z$ in absolute value. Then $R$ is stochastically larger than a $\textnormal{Binom}(K,\delta(\lambda/\sigma))$ random variable, and for any $q<\delta(\lambda/\sigma)$, the Lasso procedure satisfies Condition 3.1 with
    $\rho=\exp\left(-\;\frac{(\delta(\lambda/\sigma)-q)^{2}}{2\delta}\cdot K\right).$
  2. 2.
    Assume no fewer than $K\xi$ blocks contain at least one variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$ for some $\xi\in(0,1]$, where $\boldsymbol{\mu}=\boldsymbol{X}\boldsymbol{\theta}$ is the mean of $\boldsymbol{Y}$. Then $R$ is stochastically larger than a $\textnormal{Binom}(K,\xi/2)$ random variable, and for any $q<\xi/2$, the Lasso procedure satisfies Condition 3.1 with
    $\rho=\exp\left(-\;\frac{(\xi-2q)^{2}}{4\xi}\cdot K\right).$
A finite-sample upper bound for $\text{Var}(\widehat{\textnormal{FDR}})$ for Lasso with block-orthogonal variables then follows.
Theorem 3.1 (finite sample variance bound).
In the Gaussian linear model with block-orthogonal variables (Assumption 3.1), suppose we apply Lasso for variable selection at $\lambda>0$, and in the form of $\widehat{\textnormal{FDR}}$, $\phi_{j}=\phi(p_{j})$ only depends on $p_{j}$, the two-sided t-test p-value for $H_{j}$, where $\phi$ is a $L_{\phi}$-Lipschitz continuous function upper bounded by $c_{\phi}\geq 1$.
If Condition 3.1 holds for given values of $q$ and $\rho$, we have
$\textnormal{Var}(\widehat{\textnormal{FDR}})\leq O\left(\left(\frac{b^{4}}{q^{4}}+\frac{b^{6}}{Kq^{4}}\right)\cdot\frac{1}{K}+\frac{b^{2}}{q^{2}}\cdot\frac{1}{n-d}+b^{4}K^{2}(\rho\vee e^{-(n-d)/32})\right),$
where the big-O notation only hides constants that depend on $c_{\phi}$, $L_{\phi}$, and $\sigma^{2}$.
Remark 3.1.
An explicit formula of the variance bound that contains $c_{\phi}$, $L_{\phi}$, and $\sigma^{2}$ is provided in Appendix D.
We briefly introduce the core idea of the proof which relies on the concentration laws. Interested readers may refer to Appendix D for the full proof. Assumption 3.1 has three key consequences:
  1. 1.
    Conditioning on $\boldsymbol{S}_{j}$ fixes $\boldsymbol{X}_{k}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ for every $k$ that is not in the same block as $j$. As a result, resampling the data conditional on $\boldsymbol{S}_{j}$ would not change the selection in all other blocks. Hence, the number of selections is conditionally stable: with resampling conditional on $\boldsymbol{S}_{j}$ it is $R(\boldsymbol{D})+O(b)$.
  2. 2.
    For any $j\in B_{k}$, let $X_{B_{k}-j}$ denotes the variables in block $B_{k}$ except $X_{j}$. The selection of variable $X_{j}$ depends on $\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. Formally, we have
    $E_{j}:=\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}]=\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}]\approx\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{S}_{j}].$
    Note these $E_{j}$ are block-independent. [^5] Alert readers might have noticed that conditioning on $\boldsymbol{S}_{j}$ is not the same as conditioning on $\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$: there is an additional $\left\|\boldsymbol{Y}\right\|$ term. The analysis of the effects of $\left\|\boldsymbol{Y}\right\|$ is rather technical and it is the source of the $O(1/(n-d))$ term in the variance bound.
  3. 3.
    Likewise, the p-value $p_{j}$ depends on $\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. So $\phi_{j}(p_{j})$ are block-independent.
As a result, we have
$\widehat{\textnormal{FDR}}_{j}\;=\;\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]\cdot\phi_{j}\;\approx\;\frac{E_{j}\cdot\phi_{j}}{R+O(b)}\;\approx\;\frac{E_{j}\cdot\phi_{j}}{R}.$
The last approximation holds when $R>Kq$ with high probability and $K$ is substantially larger than $b$. Then
$\widehat{\textnormal{FDR}}\;=\;\sum_{j}\widehat{\textnormal{FDR}}_{j}\;\approx\;\frac{\sum_{k}\left(\sum_{j\in B_{k}}E_{j}\cdot\phi_{j}\right)}{R}.$
This is an average of $K$ independent random variables. Therefore, the variance is of order $O(1/K)$ by the concentration laws. A formal asymptotic result is shown in Corollary 3.1.
Corollary 3.1 (Asymptotically vanishing variance).
Under the same settings as in Theorem 3.1, consider a series of problems where we hold $c_{\phi}$, $b$, $L_{\psi}$ and $\sigma$ fixed but allow $\boldsymbol{\theta}$, $\lambda$, and $n$ to vary as $K$ increases. Suppose that at least $s(K)$ blocks contain a variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$, where $\boldsymbol{\mu}=\boldsymbol{X}\boldsymbol{\theta}$ is the mean of $\boldsymbol{Y}$.
  1. 1.
    For $40<\varepsilon<1/$ 4, suppose $s(K)\geq K^{1-4\varepsilon}$, and let $K\to\infty$ with $n-d\geq K^{1-2\varepsilon}$. Then
    $\textnormal{Var}(\widehat{\textnormal{FDR}})=O(K^{-1+4\varepsilon})\to 0.$
  2. 2.
    For $10<\xi<$ 1, suppose $s(K)\geq K\xi$, and let $K\to\infty$ with $n-d\geq 64\log(K)$. Then
    $\textnormal{Var}(\widehat{\textnormal{FDR}})=O(K^{-1})\to 0.$
[…]
3.2 Standard error estimation by bootstrap
[…]
To ameliorate this issue, we try to construct $\widehat{P}$ in such a way that most of the null variables in $P$ remain null in $\widehat{P}$, but the signal variables are estimated with little bias. Formally, let $\widehat{\mathcal{H}_{0}}$ denote an estimate of the index set of noise variables. In the parametric setting $\boldsymbol{D}\sim P_{\boldsymbol{\theta}}$, we find a sparse MLE of $\boldsymbol{\theta}$, denoted as $\hat{\boldsymbol{\theta}}$, under the constraints that $\theta_{j}=0$ for all $j\in\widehat{\mathcal{H}_{0}}$. Then we set $\widehat{P}=P_{\hat{\boldsymbol{\theta}}}$ and draw bootstrap samples to calculate $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$. We formally describe our bootstrap scheme in the parametric setting in Algorithm 1.
Algorithm 1 $\widehat{\textnormal{FDR}}$ standard error estimation (parametric)
Input: The observed data $\boldsymbol{D}$ that follows parametric distribution $P_{\boldsymbol{\theta}}$; an estimator $\widehat{\mathcal{H}_{0}}$ for the indices with $H_{j}:\,\theta_{j}=0$ is true for $j\in\widehat{\mathcal{H}_{0}}$.
Compute: the MLE under the intersection null model with $\theta_{j}=0$ for all $j\in\widehat{\mathcal{H}_{0}}$
$\hat{\boldsymbol{\theta}}(\widehat{\mathcal{H}_{0}};\boldsymbol{D})=\underset{\theta_{j}=0,\,j\in\widehat{\mathcal{H}_{0}}}{\operatorname*{argmax}}\;P_{\boldsymbol{\theta}}(\boldsymbol{D})$
Loop: for $m=1,2,\ldots,M$ do
  generate sample $\hat{\boldsymbol{D}}^{(m)}\sim P_{\hat{\boldsymbol{\theta}}}$;  compute $\widehat{\textnormal{FDR}}^{(m)}=\widehat{\textnormal{FDR}}(\hat{\boldsymbol{D}}^{(m)})$
Output: Sample standard error
$\widehat{\textnormal{s.e.}}=\sqrt{\frac{1}{M-1}\sum_{m=1}^{M}\left(\widehat{\textnormal{FDR}}^{(m)}-\frac{1}{M}\sum_{l=1}^{M}\widehat{\textnormal{FDR}}^{(l)}\right)^{2}}$
In our implementation, we construct $\widehat{\mathcal{H}_{0}}$ as the set of variables neglected by the best model in cross-validation. More precisely, our cross-validation scheme is built to select $\widehat{\mathcal{H}_{0}}$ such that the constrained MLE model $P_{\hat{\boldsymbol{\theta}}(\widehat{\mathcal{H}_{0}})}$ is close to the true distribution $P$. It estimates the constrained MLE with the candidate $\widehat{\mathcal{H}_{0}}$ on the training set and employs the negative log-likelihood on the validation set as the error measurement. A formal description of our cross-validation algorithm is in Appendix A.2. In Section 1, we demonstrated that the model that minimizes the cross-validation error tends to over-select variables moderately: it contains most of the signal variables and a bunch of noise variables. This is exactly what we need here – to include most of the signals while roughly keeping the signal sparsity level. We give examples of the constrained MLE problem next.
Example 3.1 (Gaussian linear model).
In the Gaussian linear model, the negative log-likelihood $-\log P_{\theta}(\boldsymbol{D})$ is proportional to the mean-squared error. Therefore, $\boldsymbol{\hat{\theta}}(\widehat{\mathcal{H}_{0}})$ is the least square coefficients regressing $Y$ on variables $X_{j}$ for $j\in\widehat{\mathcal{H}_{0}}^{\mathsf{c}}$.
Example 3.2 (Gaussian graphical model).
In the Gaussian graphical model, we parameterize the model using the symmetric positive definite (SPD) precision matrix $\boldsymbol{\Theta}$. Let $\widehat{\boldsymbol{\Sigma}}$ denote the sample covariance matrix. The log-likelihood is
$\log P_{\boldsymbol{\Theta}}(\boldsymbol{D})=\log\det\boldsymbol{\Theta}-\textnormal{trace}(\widehat{\boldsymbol{\Sigma}}\boldsymbol{\Theta}).$
The constrained MLE $\boldsymbol{\hat{\Theta}}$ is obtained by maximizing $P_{\boldsymbol{\Theta}}$ under the constraint that $\boldsymbol{\Theta}_{jk}=0$ for pairs $(j,k)$ in $\widehat{\mathcal{H}_{0}}$. While there are many algorithms that can solve this problem (e.g. Yuan and Lin, 2007), we apply a simple heuristic algorithm described in Appendix A.3 to save computation.
For the model-X setting, we can employ an analogous method by using a variant of the i.i.d. bootstrap that constrains $H_{j}:\;Y\perp\!\!\!\perp X_{j}\mid X_{-j}$ to be true for all $j\in\widehat{\mathcal{H}_{0}}$. Specifically, we first draw bootstrap samples of $(Y,X_{-\widehat{\mathcal{H}_{0}}})$ from the empirical distribution of the data. Then, we sample $X_{\widehat{\mathcal{H}_{0}}}$ from the conditional distribution $X_{\widehat{\mathcal{H}_{0}}}\mid(Y,X_{-\widehat{\mathcal{H}_{0}}})$, which is equal to $X_{\widehat{\mathcal{H}_{0}}}\mid X_{-\widehat{\mathcal{H}_{0}}}$ under the intersection null $\bigcap_{j\in\widehat{\mathcal{H}_{0}}}H_{j}$. We formally describe our bootstrap scheme in the model-X nonparametric setting in Algorithm 2.
If the researchers make model assumptions, the set $\widehat{\mathcal{H}_{0}}$ can be constructed in the same way as in the parametric settings. If no model assumption is made and the likelihood is unavailable, researchers can choose $\widehat{\mathcal{H}_{0}}$ via p-value-based approaches. For example, we can calculate the conditional randomization test p-values (Candès et al., 2018) for each variable and define $\widehat{\mathcal{H}_{0}}$ as all variables whose p-values are larger than $0.1$. This choice lets the majority of null variables remain null in $\widehat{P}$ while retaining the signal for strong signal variables.
Algorithm 2 $\widehat{\textnormal{FDR}}$ standard error estimation (non-parametric, model-X)
Input: The observed data matrix $\boldsymbol{D}$ whose $n$ rows are independent observations $(y_{i},x_{i,1},\ldots,x_{i,d})$; an index set $\widehat{\mathcal{H}_{0}}$ that we believe $H_{j}:\,\theta_{j}=0$ is true for $j\in\widehat{\mathcal{H}_{0}}$.
Preparation: For simplicity, dentoe $x_{i,\widehat{\mathcal{H}_{0}}}=(x_{i,j},\;j\in\widehat{\mathcal{H}_{0}})$ as the $i$th observation of $X_{\widehat{\mathcal{H}_{0}}}$. Let $\widehat{P}_{-}$ denote the empirical distribution of $(Y,X_{-\widehat{\mathcal{H}_{0}}})$
$\widehat{P}_{-}(\,\cdot\,)=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{(y_{i},x_{i,-\widehat{\mathcal{H}_{0}}})\in\,\cdot\,\right\}.$
Loop: for $m=1,2,\ldots,M$ do
  for $i=1,\ldots,n$ do: Generate sample $(\hat{y}_{i},\hat{x}_{i,-\widehat{\mathcal{H}_{0}}})\sim\widehat{P}_{-}$ and $\hat{x}_{i,\widehat{\mathcal{H}_{0}}}\sim X_{\widehat{\mathcal{H}_{0}}}\mid X_{-\widehat{\mathcal{H}_{0}}}=\hat{x}_{i,-\widehat{\mathcal{H}_{0}}}$. Let $\hat{\boldsymbol{D}}^{(m)}$ be a data matrix whose $i$th row is $(\hat{y}_{i},\hat{x}_{i,1},\ldots,\hat{x}_{i,d})$. Compute $\widehat{\textnormal{FDR}}^{(m)}=\widehat{\textnormal{FDR}}(\hat{\boldsymbol{D}}^{(m)})$
Output: Sample standard error
$\widehat{\textnormal{s.e.}}=\sqrt{\frac{1}{M-1}\sum_{m=1}^{M}\left(\widehat{\textnormal{FDR}}^{(m)}-\frac{1}{M}\sum_{l=1}^{M}\widehat{\textnormal{FDR}}^{(l)}\right)^{2}}$
4 Real world examples
In this section we illustrate our method on real scientific data in two different modeling settings. In Section 4.1, we analyze an HIV drug resistance data set using Lasso regression, and in Section 4.2 we analyze a protein signaling network using the graphical Lasso.
4.1 HIV drug resistance studies
We first re-analyze the Human Immunodeficiency Virus (HIV) drug resistance data set of Rhee et al. (2006), who found that regularized regression algorithms were effective at predicting which in vitro virus samples would be resistant to each of 16 antiretroviral drugs, based on which genetic mutations were present in the samples. In this problem it is of scientific interest not only to predict from the virus’s genetic signature whether it would be resistant to various drugs — the focus of the original study — but also to identify which mutations best explain drug resistance, which was the focus of the re-analysis in Barber and Candès (2016). As a result, both model fit and variable selection accuracy are highly relevant, so it may be scientifically attractive to tackle both problems with a unified analysis.
Our initial processing of the data follows that of Barber and Candès (2016), yielding a different regression data set for each of 16 drugs in three different drug classes. In the data set for a given drug, the response $y_{i}$ is a measure of sample $i$’s resistance to that drug, and the explanatory variables are binary indicators of whether each of several hundred mutations was present or absent in that sample.
As a measure of replicability, Barber and Candès (2016) assessed whether detected genes were also present in a list of treatment-selected mutations (TSM), which a previous study had found were enriched in patients who had been treated by drugs in each of the three classes (Rhee et al., 2005). Rhee et al. (2006) also found that excluding a priori all but the TSM list of mutations was an effective form of variable selection. Following Barber and Candès (2016), we treat the TSM list, which comes from an independent data set, as an imperfect approximation to the “ground truth” set of signal mutations. [^6] In addition to the TSM list, we include a small number of variables with extreme $t$-statistics as “ground truth” signals in each regression. Specifically, we consider all variables rejected by the Bonferroni procedure at level $\alpha=0.01$ to be “ground truth” signal variables. Having an estimate of the ground truth from another data set allows us to assess our estimator’s effectiveness in measuring the true FDR in this example; in most scientific problems, the ground truth is unavailable and we would have to rely on our method to measure FDR on its own.
(a) NFV drug, $d=207$, $n=842$.
(b) ATV drug, $d=147$, $n=328$.
(c) APV drug, $d=201$, $n=767$.
(d) DDI drug, $d=283$, $n=628$.
Figure 2: $\widehat{\textnormal{FDR}}$ along with estimated one-standard-error bars by bootstrap, for four representative experiments selected from among the 16 HIV experiments in Rhee et al. (2006). The vertical lines show the locations of the complexity parameters selected by CV and the one-standard-error rule.
Figure 2 shows our FDR estimates alongside the CV estimate of MSE, for four selected experiments representing a variety of statistical scenarios. Results for all drugs are available in Appendix E. The error bars for both CV and $\widehat{\textnormal{FDR}}$ our estimator show $\pm$±one standard error of $\widehat{\textnormal{FDR}}$ estimated by bootstrap. The $\widehat{\textnormal{FDP}}$ curves in the figures are estimated using the TSM signals as an approximation to the ground truth.
The four panels of Figure 2 illustrate how different variable selection problems can have very different trade-offs between variable selection accuracy and prediction accuracy. These scenarios would be impossible to distinguish from inspecting the CV curves, but our $\widehat{\textnormal{FDR}}$ estimator is quite effective in distinguishing them from each other. Panel (a) shows an experiment in which variable selection appears to be easy: the one-standard-error rule simultaneously achieves good model fit and low FDR. Panel (b), by contrast, shows a regime in which no lasso estimator can simultaneously achieve competitive model fit and low FDR. Panel (c) shows an experiment in which the estimated FDR is non-monotone in the complexity parameter $\lambda$: both our estimator and the TSM $\widehat{\textnormal{FDP}}$ judge the fifth variable to enter the model, gene P71.V, to be a likely noise variable, because its $t$-statistic is $-0.38$, corresponding to a $p$-value of $0.70$, and it is also absent from the TSM list (the $t$-statistics for the first four variables to enter all exceed 5). Panel (d) shows one of the three experiments in which our $\widehat{\textnormal{FDR}}$ departs substantially from the TSM $\widehat{\textnormal{FDP}}$. In this experiment, our $\widehat{\textnormal{FDR}}$ estimator is large mainly because three of the first five selected variables have $p$-values equal to $0.77$, $0.53$, and $0.18$, but the TSM $\widehat{\textnormal{FDP}}$ curve is small because all three variables appear in the TSM list.
4.2 Protein-signaling network
To test our method in a realistic Gaussian graphical model setting, we use the flow cytometry data from Sachs et al. (2005), which Friedman et al. (2008) employed as an illustrative example in their graphical Lasso paper. The data set records levels of $11$ proteins in individual cells under $14$ different perturbation conditions, in order to estimate relationships in the protein signaling network. For each experimental condition, a sample of cells received a different biological stimulation, and the cells’ protein levels were observed.
(a) Protein causal DAG.
(b) Induced undirected graph.
Figure 3: Causal DAG supplied by Sachs et al. (2005) and its induced undirected graph, which we use for independent validation of our FDR estimator. A pair of proteins have an edge in the undirected graph if they are not conditionally independent given the other proteins. There are $22$ edges in the induced undirected graph, out of $55$ possible edges.
Sachs et al. (2005) supplied a causal directed acyclic graph (DAG) on the proteins, shown here in Figure 3(a). The causal graph was based on theoretical and empirical knowledge from previous literature and validation experiments that were independent from the flow cytometry dataset analyzed by Friedman et al. (2008). After converting the directed causal graph to its induced undirected conditional dependence graph, shown in Figure 3(b), we obtain an estimate of the set of true signals. As in the HIV data set, we also include as ground truth signals in each experiment a few additional pairs for which the $p$-values pass the Bonferroni threshold at level $\alpha=0.01$; i.e., those whose $t$-statistic $p$-values are below $1.8\times 10^{-4}$. Similarly to the TSM list in the HIV data set, the set of edges obtained above functions as an approximation to the ground truth, giving us a point of reference to help evaluate our FDR estimator.
(a) Experiment 1: $d=11$, $n=853$.
(b) Experiment 11: $d=11$, $n=799$.
Figure 4: $\widehat{\textnormal{FDR}}$ along with estimated one-standard-error bars by bootstrap. The $\widehat{\textnormal{FDP}}$ curves are calculated by with a “ground truth” based on the dependence graph in Figure 3(b). Cross-validation alone tells little about the right way to select variables (pairs) while our $\widehat{\textnormal{FDR}}$ is informative.
Figure 4 shows two representative examples of the results from running the graphical Lasso on the $14$ experimental settings. Results for all experiments are available in Appendix E. As in the HIV study, we show our $\widehat{\textnormal{FDR}}$ and its estimated standard error, alongside an estimate of FDP calculated using the causal DAG supplied by Sachs et al. (2005) (denoted as the DAG $\widehat{\textnormal{FDP}}$ in Figure 4), and the CV estimate of the predictive log-likelihood. In both panels of Figure 4, the models that we would select by cross-validation are nearly useless: one selects almost no edges, while the other selects every edge. Friedman et al. (2008) analyzed the data somewhat differently than we did, pooling all $14$ experiments together, but likewise reported that the CV curve continuously decreased until all edges were present in the model. By contrast, our $\widehat{\textnormal{FDR}}$ gives useful information about the selection accuracy, and suggests in each experiment a natural stopping point.
The main way in which the $\widehat{\textnormal{FDR}}$ curves in all $14$ experiments depart from the DAG $\widehat{\textnormal{FDP}}$ estimates is that our method consistently estimates the FDR to be a bit larger across the board. The reason for this departure is that the evidence for a number of the edges identified in the DAG appears to be very weak in the data sets we analyze. That is, some of the “ground truth” edges identified by Sachs et al. (2005) appear to be absent in the actual data. This fact does not imply that those edges are truly absent in the data set, but if they are present their signals are quite weak, so our estimator treats them as likely nulls and thereby “overestimates” the FDR.
5 Simulation studies
5.1 Gaussian linear model
In this section, we evaluate the performance of $\widehat{\textnormal{FDR}}$ in simulation environments where the ground truth is known. We begin with the Gaussian linear model and employ Lasso as the variable selection procedure.
The explanatory variables are generated under four primary scenarios, which cover independent, positively correlated, and negatively correlated explanatory variables as well as sparse $\boldsymbol{X}$ matrices:
  1. 1.
[…]
  2. 2.
    Auto-regression (X-AR): $X_{j}$ ($j=1,\ldots,d)$ is an AR(1) Gaussian random process with $\text{cov}(X_{j},X_{j+1})=0.5$.
  3. 3.
    Regression coefficients auto-regression (Coef-AR): The OLS coefficients $\hat{\theta}_{j}$ ($j=1,\ldots,d)$ follow an AR(1) Gaussian random process with $\text{cov}(\hat{\theta}_{j},\hat{\theta}_{j+1})=0.5$. That is, Cov(X) is banded with negative off-diagonal entries.
  4. 4.
[…]
The signal variable set $\mathcal{H}_{0}^{\mathsf{c}}$ is a random subset of $[d]$ that has cardinality $d_{1}$, uniformly distributed among all such subsets. To mimic realistic scenarios where there is a mix between weak and strong signal variables, we generate nonzero signals with random signal strength
[…]
for all $j\in\mathcal{H}_{0}^{\mathsf{c}}$, where $\textnormal{Exp}(1)$ represents a rate-$1$ exponentially distributed random variable and $\theta^{*}$ is a fixed value controlling the average signal strength. We calibrate a moderate signal strength $\theta^{*}$ such that the selection procedure will have type II error $\textnormal{FPR}=0.2$ when $\textnormal{FDR}=0.2$, where FPR stands for the False Positive Rate
$\textnormal{FPR}=1-\mathbb{E}\left[\frac{|\mathcal{R}\cap\mathcal{H}_{0}^{\mathsf{c}}|}{d_{1}}\right].$
[…]
Figure 5: Performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ in the Gaussian linear model. The error bars show the $5\%$ to $95\%$ quantiles of the empirical distributions and the solid lines show the sample average. $\widehat{\textnormal{FDR}}$ shows a small non-negative bias in estimating FDR and $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$ successively captures the magnitude of uncertainty of $\widehat{\textnormal{FDR}}$.
Next, we stress-test the performance of $\widehat{\textnormal{FDR}}$ in several variations on the IID normal scenario above. The other settings remain the same except for those described in each of the four scenarios below. The first three scenarios are intended to illustrate cases where $\widehat{\textnormal{FDR}}$ may have a substantial bias and the last scenario to illustrate a case where $\widehat{\textnormal{FDR}}$ could have a high variance.
  1. 1.
    Weaker Signals: A weaker signal strength $\theta^{*}$ such that Lasso has type II error $\textnormal{FPR}=0.7$ when $\textnormal{FDR}=0.2$. In this setting, only about $50\%$ of the signal variables have their $\widehat{\textnormal{FDR}}_{j}$ contribution killed by the bias correction $\phi$, which could potentially lead to significant bias.
  2. 2.
    Dense Signals: We have $d=60$ variables and half of them are signals ($d_{1}=30$). The number of observations is $n=180$ (same aspect ratio). This could lead to a higher bias because signal variables’ contribution is the source of bias.
  3. 3.
    Aspect Ratio Close to 1: We have $n=600$ observations for $d=500$ variables. The IID Gaussian observations are thus closer to collinear. This could lead to higher bias because the $p$-values we use in the bias correction are weak.
  4. 4.
    Very Sparse: We have only $d_{1}=10$ signal variables, so typical model sizes are small, leading to high variance in FDP and potentially also in $\widehat{\textnormal{FDR}}$.
Figure 6: Performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ in additional settings as an edge-case test for bias (column 1-3) and variance (column 4).
In Figure 6, $\widehat{\textnormal{FDR}}$ does not show a substantial bias in the Weaker Signals setting, but does in the Dense Signals and Large Aspect Ratio settings when $\lambda$ gets small. In the Weaker Signals setting, although many signal variables have $\phi_{j}>0$, each of them contributes very little to $\widehat{\textnormal{FDR}}$ since their probability of being selected is small under the null $H_{j}:\boldsymbol{\theta}_{j}=0$; moreover, because there are many more noise variables than signal variables, the total contribution from the former is still much larger than the contribution from the latter. However, these excess contributions do cause a bias in the Dense Signals setting, because the signal variables’ total contributions is not offset by the noise variables’ total contribution. The story is a bit different in the Large Aspect Ratio setting. In this scenario, even though the signals are strong enough to be reliably detected by the Lasso, the $p$-values are nearly powerless because the variables are nearly collinear. Hence many signal variables have $\phi_{j}>0$. Unlike in the Weaker Signals setting, a signal variable still has a substantial chance to be selected by Lasso under $H_{j}$ due to the strong correlation structure. Therefore they make a substantial contribution to the bias. Finally, in the Very Sparse setting, we see a higher variance for $\widehat{\textnormal{FDR}}$ owing to the typically small size of $\mathcal{R}$, but the FDP also has a larger variance.
5.2 Model-X setting
Next, we test the model-X version of our procedure in two linear modeling scenarios. The response $Y$ is either a continuous response with Gaussian errors, where we employ Lasso, or a binary response that follows a logistic regression model:
$\ln\left(\frac{\mathbb{P}(Y=1)}{1-\mathbb{P}(Y=1)}\right)=\sum_{j=1}^{d}X_{j}\theta_{j},$
where we employ $\ell_{1}$-penalized logistic regression for variable selection. The other settings are the same as in Section 5.1.
Figure 7: Performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ in the model-X settings. The error bars show the $5\%$ to $95\%$ quantiles of the empirical distributions and the solid lines show the sample average. $\widehat{\textnormal{FDR}}$ shows a small non-negative bias in estimating FDR and $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$ successively captures the magnitude of uncertainty of $\widehat{\textnormal{FDR}}$.
Figure 7 shows the performance of our estimators where the regression problem (left panel) is high-dimensional with $d=500$ and $n=400<d$ and the classification problem (right panel) has $d=500$ and $n=1500$. As in Section 5.1, our estimators perform well.
5.3 Gaussian graphical model
Next, we test our method in the problem of detecting a sparse dependence graph as described in Example 2.2. We generate independent samples of $d$ variables
$(X_{1},\ldots,X_{d})\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma}),\quad i=1,\ldots,n,$
[…]
Figure 8: Performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ in the Gaussian graphical model. The error bars show the $5\%$ to $95\%$ quantiles of the empirical distributions and the solid lines show the sample average. $\widehat{\textnormal{FDR}}$ shows a small non-negative bias in estimating FDR and $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$ successively captures the magnitude of uncertainty of $\widehat{\textnormal{FDR}}$.
5.4 Unfavorable case: equicorrelated Gaussian
In Section 3.1, we show that $\widehat{\textnormal{FDR}}$ has a small standard error if the variables have a block-independent structure. In this section, we examine the performance of $\widehat{\textnormal{FDR}}$ on a practical problem with an unfavorable structure – the Multiple Comparisons to Control (MCC) problem (Dunnett, 1955), where we estimate the effects of $d$ treatments by comparing treatment group for each one to a common control group. We assume all $d+1$ groups have the same sample size. Mathematically, the problem is equivalent to considering $d$-dimensional Gaussian vector $\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{\theta},\boldsymbol{\Sigma})$, where $\boldsymbol{\Sigma}\in\mathbb{R}^{d\times d}$ has diagonal $1$ and off-diagonal $\rho=0.5$. So all $Z_{j}$ variables are positively equicorrelated: $\textnormal{cov}(Z_{j},Z_{k})=\rho$ for $j\neq k$. We are interested in selecting the signal variables with $\theta_{j}\neq 0$. As discussed in Section 3.1, this is an unfavorable case for $\widehat{\textnormal{FDR}}$.
The problem can be equivalently formalized as variable selection in a Gaussian linear model whose OLS coefficients are $\boldsymbol{Z}$. We employ the Lasso selection procedure and test $\widehat{\textnormal{FDR}}$ in this unfavorable case. Figure 9 shows the performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ on the MCC problem with $d=500$, $n=1500$, and $d_{1}=30$ signal variables. As expected, $\widehat{\textnormal{FDR}}$ shows a nonnegative bias and an uncertainty much larger than the oracle FDP. But like in previous simulations, the $\widehat{\textnormal{s.e.}}$ tells the magnitude of uncertainty of $\widehat{\textnormal{FDR}}$ and can warn the researchers that $\widehat{\textnormal{FDR}}$ is not very reliable.
Figure 9: Performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ on the MCC problem. The error bars show the $5\%$ to $95\%$ quantiles of the empirical distributions and the solid lines show the sample average. $\widehat{\textnormal{FDR}}$ shows a large variance but $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$ successively captures the magnitude of uncertainty of $\widehat{\textnormal{FDR}}$.
5.5 Robustness to model misspecification
The robustness of $\widehat{\textnormal{FDR}}$ is closely related to the robustness of the conditional test on $H_{j}$. Take the Gaussian linear model, for example. Figure 10 demonstrates the calculation of
$\widehat{\textnormal{FDR}}^{*}_{j}=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]$
on the Lasso selection procedure. Conditioning on $\boldsymbol{S}_{j}$, there is a one-to-one map between the t-statistic for $H_{j}$ and the sufficient statistic $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{y}$. So this conditional expectation can be calculated as a one-dimensional integral of ${\mathbf{1}\left\{j\in\mathcal{R}\right\}}/{R}$ (shown in red curve) against the $t$-distribution density (shown in blue curve). And the calculated $\widehat{\textnormal{FDR}}^{*}_{j}$ is the area in purple. Now consider the Gaussian linear model is misspecified. For example, the noise $\varepsilon_{i}$ is not Gaussian but follows a t-distribution with $5$ degrees of freedom. The true conditional density of the t-statistic, assuming $\theta_{j}=0$, is shown by the orange dashed curve. As expected, the conditional t-test is robust: the true density of the $t$-statistic is not far from the theoretical $t$-distribution density. As a result, the calculated $\widehat{\textnormal{FDR}}^{*}_{j}$ does not change much with our misspecified model. $\phi_{j}$ is robust for the same reason if it is based on the t-test p-value. Therefore, we expect $\widehat{\textnormal{FDR}}$ to be robust as long as we have confidence in the robustness of the conditional t-test on each null hypothesis.
Figure 10: A demonstration of computing $\widehat{\textnormal{FDR}}^{*}_{j}$. The condition expectation is essentially a one-dimensional integral of ${\mathbf{1}\left\{j\in\mathcal{R}\right\}}/{R}$ (shown in red) against the $t$-distribution density (shown in blue). And the calculated $\widehat{\textnormal{FDR}}^{*}_{j}$ is the area in purple. With model misspecification, if the true conditional density (shown in dashed orange) is not far from the misspecified $t$-distribution density, then the calculation of $\widehat{\textnormal{FDR}}^{*}_{j}$ is approximately correct.
The reasoning can be extended to more general scenarios. For example, in the Gaussian graphical model with graphical Lasso selection, we expect $\widehat{\textnormal{FDR}}$ to be robust to model misspecification if the conditional t-test on $H_{jk}$ is robust. In the model-X settings, consider the selection procedures where $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ is sufficient. These include Lasso and penalized logistic regression selection. We expect $\widehat{\textnormal{FDR}}$ to be robust if the conditional distribution of $\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ is robust.
Figure 11: Performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ with model misspecification. The error bars show the $5\%$ to $95\%$ quantiles of the empirical distributions and the solid lines show the sample average.
In Section 4, we have seen $\widehat{\textnormal{FDR}}$ works well on real-world problems where the model assumptions are never exactly true. Here in Figure 11, we test the robustness of $\widehat{\textnormal{FDR}}$ in a clean simulation study environment. We consider the same settings as previous simulations but with different types of model misspecifications. In the “Heteroscedasticity” setting, we introduce heteroscedasticity to the Gaussian linear model. That is, we generate
$Y=\sum_{j=1}^{d}X_{j}\theta_{j}+\exp\left(\sum_{j=1}^{d}|X_{j}\theta_{j}|\right)\cdot\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,1).$
[…]
6 Computational details
The major computational task in our method is to calculate conditional expectations
$\widehat{\textnormal{FDR}}^{*}_{j}=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]$
for $j=1,\ldots,d$ with $\psi_{j}\neq 0$. These calculations can be easily parallelized. In a generic setting, we can use Monte-Carlo integration to approximate $\widehat{\textnormal{FDR}}^{*}_{j}$. The worst-case time complexity for computing $\widehat{\textnormal{FDR}}$ is $O(d\cdot M\cdot T)$, where $M$ is the number of Monte-Carlo samples and $T$ is the running time of the selection procedure. We note that calculating the integrand of $\widehat{\textnormal{FDR}}_{j}^{*}$ for each Monte-Carlo sample has the same computation complexity as calculating one placebo statistic in the conditional randomization test (CRT, Candès et al. (2018)) with the same selection procedure for variable $j$. In our implementation, we use $20$ Monte-Carlo samples for computing each $\widehat{\textnormal{FDR}}^{*}_{j}$ and $10$ Bootstrap samples.
While the Monte-Carlo procedure can work with any selection procedure, for commonly used ones like Lasso and FS under the Gaussian linear model, we can develop exact and fast algorithms by leveraging the special structures. In particular, $\widehat{\textnormal{FDR}}^{*}_{j}$ reduces to a one-dimensional integral. We briefly describe the algorithms below and leave details in Appendix C.
Example 6.1 (fast and exact calculation of $\widehat{\textnormal{FDR}}_{j}^{*}$ for Lasso).
Lei (2019) proposed a fast homotopy algorithm that computes the entire Lasso solution path when the sufficient statistic $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ varies in a given direction in the context of conformal prediction. With simple modifications, we apply their algorithm to obtain the Lasso selection set $\mathcal{R}$ for all possible values of $\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ given $\boldsymbol{S}_{j}$. This provides a full characterization of the piecewise constant function $\mathbf{1}\left\{j\in\mathcal{R}\right\}/{R}$ and hence yields the exact value of $\widehat{\textnormal{FDR}}_{j}^{*}$.
The runtime of the homotopy algorithm scales linearly in the number of knots along the homotopy path. At each knot, the algorithm applies a handful of $d$-dimensional matrix-vector multiplications. As a result, the algorithm is faster for a reasonably large $\lambda$ where $\mathcal{R}$ varies less over the range of $\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ given $\boldsymbol{S}_{j}$.
Example 6.2 (fast and exact calculation of $\widehat{\textnormal{FDR}}_{j}^{*}$ for FS).
At the $k$-th step, the FS procedure for the Gaussian linear model selects a variable $j_{k}$ that has the largest $|\boldsymbol{X}_{j}^{(k)}\cdot\boldsymbol{Y}^{(k)}|$ where $\boldsymbol{X}^{(k)},\boldsymbol{Y}^{(k)}$ can be obtained by applying a sequence of normalization and orthogonalization steps. When the number of selected variables $k$ is fixed, $\widehat{\textnormal{FDR}}^{*}_{j}$ simplifies into $\mathbb{P}_{H_{j}}(j\in\mathcal{R}\mid S_{j})/k$. The numerator can be decomposed into $\sum_{\ell=1}^{k}\mathbb{P}_{H_{j}}(j\text{ is selected at the }\ell\text{-th step}\mid S_{j})$. We can calculate each term by using an auxiliary FS procedure that has the same computation cost as the original one. Again, each step only involves a handful of $d$-dimensional matrix-vector multiplications. Thus, the algorithm is faster when the number of selections $k$ is small.
The running time of evaluating $\widehat{\textnormal{FDR}}$ (without bootstrap s.e. estimation) at all $10$ different tuning parameter values in the simulation problems in Section 5 is typically several minutes on a single-cored computer. We report the exact running time in Appendix F.
7 Summary
In this work we present a new approach for estimating FDR of any variable selection procedure. Our estimator can be applied in common statistical settings including the Gaussian linear model, the Gaussian graphical model, and the nonparametric model-X framework, requiring no additional statistical assumptions in these settings beyond those required for standard hypothesis tests. Our estimator can be evaluated alongside the cross-validation curve to inform the analyst about the tradeoff between variable selection accuracy and prediction accuracy over the entire regularization path. The bias of the estimator is non-negative in finite samples and typically small. The standard error can be assessed using the specialized bootstrap and it is typically vanishing when there are many variables with only local dependency. Examples with real-world and simulation data demonstrate our estimator works well in both bias and standard error. R package is available online at the Github repository https://github.com/yixiangLuo/hFDR.
Reproducibility
Our FDR estimator is implemented in an R package available online at the Github repository
github.com/yixiangLuo/hFDR. The R code to reproduce all simulations and figures is available at github.com/yixiangLuo/hFDR_expr.
Appendix A Additional details in experiments and methodologies
A.1 Simulation settings for the illustrative examples
[…]
In Figure 1(b), the explanatory variables $X_{j}$ ($j=1,\ldots,d)$ form an AR(1) Gaussian random process with $\text{cov}(X_{j},X_{j+1})=0.8$. The other settings are the same as above.
A.2 Cross-validation for constructing $\widehat{\mathcal{H}_{0}}$ in $\widehat{\textnormal{s.e.}}$
In this section, we describe the cross-validation algorithm we use to construct the set $\widehat{\mathcal{H}_{0}}$ in computing $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$. Formally, let $\boldsymbol{D}_{\text{trn},k}$ and $\boldsymbol{D}_{\text{val},k}$ be the $k$th training and validation data set in a $K$-folds cross-validation, respectively. Throughout the section, we write $\boldsymbol{\hat{\theta}}(\widehat{\mathcal{H}_{0}}(\lambda);\boldsymbol{D}_{\text{trn},k})$ with $\widehat{\mathcal{H}_{0}}(\lambda)=\mathcal{R}^{\mathsf{c}}(\lambda;\boldsymbol{D}_{\text{trn},k})$ as $\boldsymbol{\hat{\theta}}(\lambda;\boldsymbol{D}_{\text{trn},k})$ for simplicity of notation. We have
$\boldsymbol{\hat{\theta}}(\lambda;\boldsymbol{D}_{\text{trn},k})=\underset{\theta_{j}=0,\,j\not\in\mathcal{R}(\lambda;\boldsymbol{D}_{\text{trn},k})}{\operatorname*{argmax}}\;P_{\boldsymbol{\theta}}(\boldsymbol{D}_{\text{trn},k}).$
Then we find the best $\lambda$ that achieves the largest averaged likelihood on the validation sets
[…]
and define
$\widehat{\mathcal{H}_{0}}=\left\{j:\;j\not\in\mathcal{R}(\lambda_{\text{cv}};\boldsymbol{D})\right\}.$
A.3 Sparse model estimation with graphical Lasso
We start with the unconstrained MLE
$\boldsymbol{\tilde{\Theta}}:=\underset{\boldsymbol{\Theta}}{\operatorname*{argmax}}\;\log P_{\boldsymbol{\Theta}}(\boldsymbol{D})=\log\det\boldsymbol{\Theta}-\textnormal{trace}(\widehat{\boldsymbol{\Sigma}}\boldsymbol{\Theta})$
and then enforce the entries in $\widehat{\mathcal{H}_{0}}$ to be zero
$\boldsymbol{\hat{\Theta}}_{jk}\leftarrow\begin{cases}0,&\text{if }(j,k)\in\widehat{\mathcal{H}_{0}}\text{ or }(k,j)\in\widehat{\mathcal{H}_{0}}\\ \boldsymbol{\tilde{\Theta}}_{jk},&\text{otherwise}\end{cases}.$
It is possible that the resulting $\boldsymbol{\hat{\Theta}}$ is not positive definite. If that is the case, we increase the eigenvalues of $\boldsymbol{\hat{\Theta}}$ by the same amount until it is positive definite.
Appendix B Example where $\widehat{\textnormal{FDR}}$ has non-vanishing variance
We describe an example where $\widehat{\textnormal{FDR}}$ has a non-vanishing standard error as $d$ increases.
Consider a $d$-dimensional Gaussian vector $\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{\theta},\boldsymbol{\Sigma})$, where $\boldsymbol{\Sigma}\in\mathbb{R}^{d\times d}$ has diagonal $1$ and off-diagonal $\rho=0.8$. So all $Z_{j}$ variables are positively equicorrelated: $\textnormal{cov}(Z_{j},Z_{k})=\rho$ for $j\neq k$. To select the signal variables with $\theta_{j}\neq 0$, we do a simple one-sided thresholding,
=$\mathcal{R}=\left\{j:\;Z_{j}\geq 0.5\cdot\sqrt{2\log d}\right\},$
and set $\psi_{j}=\mathbf{1}\left\{p_{j}>0.5\right\}$, where $p_{j}$ is the one-sided marginal z-test p-value. Let $\boldsymbol{S}_{j}=\boldsymbol{Z}_{-j}-\boldsymbol{\Sigma}_{-j,j}\cdot Z_{j}$ and $\boldsymbol{\theta}=\boldsymbol{0}$. Figure 12 shows the standard deviation of $\widehat{\textnormal{FDR}}$ does not converge to 00 as the number of variables $d$ approach infinity.
Figure 12: For variable selection on strongly positively equicorrelated $Z$-variables, standard deviation of $\widehat{\textnormal{FDR}}$ does not converge to 00 as the number of variables approach infinity.
The noisy behavior of $\widehat{\textnormal{FDR}}$ is a consequence of the strong equicorrelation between the $Z$ variables. Due to the correlation, perturbing $Z_{j}$ strongly affects all other $Z$ variables and hence the entire selection set. As a result, $\widehat{\textnormal{FDR}}_{j}$ are highly positively correlated. The backbone idea for the concentration laws, which builds upon the cancellation of uncorrelated/independent noise, does not apply.
Appendix C Fast algorithms for computing $\widehat{\textnormal{FDR}}^{*}_{j}$
In this section, we describe the fast and exact algorithms for computing $\widehat{\textnormal{FDR}}^{*}_{j}$ for the Lasso (Example 6.1) and FS (Example 6.2) procedures in the Gaussian linear model. Our goal is to compute
$\widehat{\textnormal{FDR}}^{*}_{j}=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right].$
To avoid confusion, we use $j$ to refer to the particular index for which we want to compute $\widehat{\textnormal{FDR}}^{*}_{j}$ and use $i$ to refer to a generic variable index
C.1 fast and exact calculation for Lasso
The idea is inspired by Lei (2019). Recall the Lasso coefficients are
$\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}}(\lambda)=\operatorname*{argmin}_{\boldsymbol{t}}\;\frac{1}{2}\,\left\|\boldsymbol{Y}-\boldsymbol{X}\boldsymbol{t}\right\|_{2}^{2}+\lambda\left\|\boldsymbol{t}\right\|_{1}.$
The first-order optimal condition is
$-\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}(\boldsymbol{Y}-\boldsymbol{X}\cdot\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}})+\boldsymbol{\nu}=\boldsymbol{0},$
where $\boldsymbol{\nu}\in\mathbb{R}^{d}$ is the dual variable satisfying $\nu_{i}={\textnormal{sign}}(\widehat{\theta}^{\textnormal{Lasso}}_{i})\cdot\lambda$ for $\widehat{\theta}^{\textnormal{Lasso}}_{i}\neq 0$ and $\nu_{i}\in[-\lambda,\lambda]$ otherwise. Suppose we have computed the selection set $\mathcal{R}$. Then by the optimal condition, $\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}}$ has a explicit expression
$\widehat{\boldsymbol{\theta}}_{\mathcal{R}}^{\textnormal{lasso}}=\left(\boldsymbol{X}_{\mathcal{R}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{\mathcal{R}}\right)^{-1}\left(\boldsymbol{X}_{\mathcal{R}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}-{\textnormal{sign}}(\widehat{\boldsymbol{\theta}}_{\mathcal{R}}^{\textnormal{lasso}})\cdot\lambda\right),\quad\widehat{\boldsymbol{\theta}}_{-\mathcal{R}}^{\textnormal{lasso}}=\boldsymbol{0}.$ (15)
Now condition on $\boldsymbol{S}_{j}$ and let the value of $\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ vary. We can compute $\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}}$ by (15) as long as $\mathcal{R}$ does not change. The change happens when the computed $\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}}$ becomes critical:
  1. 1.
    an already selected variable $i\in\mathcal{R}$ gets kicked out of the selection set if and only if its computed coefficient $\widehat{\theta}^{\textnormal{Lasso}}_{i}$ become zero;
  2. 2.
    an unselected variable $i\not\in\mathcal{R}$ gets selected if and only if the absolute value of its dual variable $\nu_{i}=\boldsymbol{X}_{i}^{\mkern-1.5mu\mathsf{T}}(\boldsymbol{Y}-\boldsymbol{X}\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}})$ touches $\lambda$.
Readers may refer to Lei (2019) for a more detailed discussion on these two conditions.
Combine all the above. We start from a computed $\mathcal{R}$, let the value of $\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ vary (say increase), and calculate $\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}}$ by (15). Once either of the two conditions is satisfied, we update $\mathcal{R}$ and move forward. We perform a low-rank update on $\left(\boldsymbol{X}_{\mathcal{R}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{\mathcal{R}}\right)^{-1}$ in (15) by the Sherman–Morrison–Woodbury formula. This algorithm reveals the exact Lasso solution path. To compute $\widehat{\textnormal{FDR}}^{*}_{j}$, we only care about the points of change of $\mathcal{R}$, which can be found by solving simple linear equations derived from the two conditions. The calculation of $\widehat{\textnormal{FDR}}^{*}_{j}$ then becomes a one-dimensional integral of the known piecewise constant function $\mathbf{1}\left\{j\in\mathcal{R}\right\}/{R}$ on a known probability measure.
C.2 fast and exact calculation for FS
In the Gaussian linear model, FS first centers and normalizes the observed variables
$\boldsymbol{Y}^{(0)}\leftarrow\boldsymbol{Y}-\frac{1}{n}(\boldsymbol{Y}\cdot\boldsymbol{1})\cdot\boldsymbol{1},\quad\boldsymbol{X}_{j}^{(0)}\leftarrow\boldsymbol{X}_{j}-\frac{1}{n}(\boldsymbol{X}_{j}\cdot\boldsymbol{1})\cdot\boldsymbol{1},\quad\boldsymbol{X}_{j}^{(0)}\leftarrow\boldsymbol{X}_{j}^{(0)}/\left\|\boldsymbol{X}_{j}^{(0)}\right\|_{2},\quad\forall j.$
Then do the following operations in each step.
  1. 1.
    Selection: $J^{(\lambda)}={\operatorname*{argmax}}_{j\not\in\mathcal{R}(\lambda-1)}\;|\boldsymbol{X}_{j}^{(\lambda-1)}\cdot\boldsymbol{Y}^{(\lambda-1)}|$;
  2. 2.
    Gram-Schmidt orthogonalization: $\boldsymbol{Y}^{(\lambda)}\leftarrow\boldsymbol{Y}^{(\lambda-1)}-(\boldsymbol{Y}^{(\lambda-1)}\cdot\boldsymbol{X}_{j^{(\lambda)}}^{(\lambda-1)})\cdot\boldsymbol{X}_{j^{(\lambda)}}^{(\lambda-1)}$ and $\boldsymbol{X}_{j}^{(\lambda)}\leftarrow\boldsymbol{X}_{j}^{(\lambda-1)}-(\boldsymbol{X}_{j}^{(\lambda-1)}\cdot\boldsymbol{X}_{j^{(\lambda)}}^{(\lambda-1)})\cdot\boldsymbol{X}_{j^{(\lambda)}}^{(\lambda-1)}$ for all $j\not\in\mathcal{R}(\lambda)$.
  3. 3.
    renormalization: $\boldsymbol{X}_{j}^{(\lambda)}\leftarrow\boldsymbol{X}_{j}^{(\lambda)}/\left\|\boldsymbol{X}_{j}^{(\lambda)}\right\|_{2}$ for all $j\not\in\mathcal{R}(\lambda)$.
Since we select one variable at each step, the number of selections $R$ is known. Therefore, to compute $\widehat{\textnormal{FDR}}^{*}_{j}$, it suffices to calculate $\mathbb{P}_{H_{j}}(j\in\mathcal{R}\mid S_{j})$.
Now condition on $\boldsymbol{S}_{j}$ and let the value of $\boldsymbol{X}_{j}\cdot\boldsymbol{Y}$ vary. Consider an auxiliary FS procedure where we enforce that variable $X_{j}$ is not to be selected:
$J^{(\lambda)}={\operatorname*{argmax}}_{l\not\in\mathcal{R}(\lambda-1)\cup\left\{j\right\}}\;|\boldsymbol{X}_{l}^{(\lambda-1)}\cdot\boldsymbol{Y}^{(\lambda-1)}|.$
This auxiliary procedure is determined by the value of $\boldsymbol{S}_{j}$ without looking at $\boldsymbol{X}_{j}\cdot\boldsymbol{Y}$. At each step, if
$|\boldsymbol{X}_{j}^{(\lambda-1)}\cdot\boldsymbol{Y}^{(\lambda-1)}|>|\boldsymbol{X}_{J^{(\lambda)}}^{(\lambda-1)}\cdot\boldsymbol{Y}^{(\lambda-1)}|.$
then the variable with index $j$, instead of $J^{(\lambda)}$, will be selected. By this inequality and certain linear transforms, we obtain the set of values of $\boldsymbol{X}_{j}\cdot\boldsymbol{Y}$ such that variable $X_{j}$ is selected at each step. Taking a union of them and using the known distribution of $\boldsymbol{X}_{j}\cdot\boldsymbol{Y}$, we can compute $\mathbb{P}_{H_{j}}(j\in\mathcal{R}\mid S_{j})$. The algorithm requires one auxiliary FS process (shared among many different $\widehat{\textnormal{FDR}}^{*}_{j}$) and produces a fast and exact calculation. Similar ideas apply to the model-X settings.
Appendix D Deferred proofs
In this section, we assume Assumption 3.1 and prove the propositions in Section 3.1. Throughout the section, we use $\mathcal{R}(\lambda)$ to denote the selection set of Lasso. Define $\mathcal{R}_{k}=\mathcal{R}\cap B_{k}$ as the selected variables in block $k$ and $R_{k}=|\mathcal{R}_{k}|$ as its cardinality. Throughout the section, we consider any pair of $(j,k)$ such that $j\in B_{k}$. We start with showing two technical lemmas.
Lemma D.1.
$\mathcal{R}_{k}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. The p-value $p_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$, where $\textnormal{RSS}=\left\|\boldsymbol{Y}\right\|^{2}-\left\|\boldsymbol{X}(\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}\right\|^{2}$.
Proof.
Let $\boldsymbol{Y}^{(k)}$ be the projection of $\boldsymbol{Y}$ onto the subspace spanned by the columns of $\boldsymbol{X}_{B_{k}}$, and denote the projection of $\boldsymbol{Y}$ onto the columns of $\boldsymbol{X}$ as $\widehat{\boldsymbol{Y}}=\boldsymbol{X}(\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. Then
$\widehat{\boldsymbol{Y}}=\sum_{k=1}^{K}\boldsymbol{Y}^{(k)}$
and the $\boldsymbol{Y}^{(k)}$s are orthogonal to each other under Assumption 3.1. We rewrite the Lasso coefficients
$\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}}(\lambda)$ $=\operatorname*{argmin}_{\boldsymbol{t}}\;\frac{1}{2}\,\left\|\boldsymbol{Y}-\boldsymbol{X}\boldsymbol{t}\right\|_{2}^{2}+\lambda\left\|\boldsymbol{t}\right\|_{1}$
$=\operatorname*{argmin}_{\boldsymbol{t}}\;\frac{1}{2}\,\left\|\widehat{\boldsymbol{Y}}-\boldsymbol{X}\boldsymbol{t}\right\|_{2}^{2}+\lambda\left\|\boldsymbol{t}\right\|_{1}$
$=\operatorname*{argmin}_{\boldsymbol{t}}\;\frac{1}{2}\,\left\|\sum_{k=1}^{K}\boldsymbol{Y}^{(k)}-\boldsymbol{X}_{B_{k}}\boldsymbol{t}_{B_{k}}\right\|_{2}^{2}+\lambda\sum_{k=1}^{K}\left\|\boldsymbol{t}_{B_{k}}\right\|_{1}$
$=\operatorname*{argmin}_{\boldsymbol{t}}\;\sum_{k=1}^{K}\frac{1}{2}\,\left\|\boldsymbol{Y}^{(k)}-\boldsymbol{X}_{B_{k}}\boldsymbol{t}_{B_{k}}\right\|_{2}^{2}+\lambda\left\|\boldsymbol{t}_{B_{k}}\right\|_{1}.$
We see solving for $\widehat{\theta}^{\textnormal{Lasso}}$ is equivalent to solving for Lasso coefficients for $K$ subproblems of regressing $\boldsymbol{Y}$ on $\boldsymbol{X}_{B_{k}}$. So $\widehat{\theta}^{\textnormal{Lasso}}_{B_{k}}(\lambda)$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}^{(k)}$, which is equal to $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. Therefore, $\mathcal{R}_{k}=\left\{j\in B_{k}:\;\widehat{\theta}^{\textnormal{Lasso}}_{j}(\lambda)\neq 0\right\}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
Recall the p-value $p_{j}$ is a function of the t-statistic
$\frac{\left((\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\right)_{j}\cdot\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}}{\sqrt{\left((\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\right)_{jj}\cdot\textnormal{RSS}/(n-d)}}.$
The denominator depends on $\boldsymbol{Y}$ only through RSS. For the numerator, note $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}$ is block-diagonal by the orthogonality in Assumption 3.1. We have
$\left((\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\right)_{j}\cdot\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}=\left((\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{B_{k}})^{-1}\right)_{i(j,k)}\cdot\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},$
where $i(j,k)$ is the index of variable $X_{j}$ in block $B_{k}$. Therefore $p_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$. ∎
Lemma D.2 (mean value).
Suppose $f(x)\in[a,b]\subseteq\mathbb{R}^{+}$ for $x\in D$ and $1/f$ is integrable with respect to the measure $\mu$. Then there exists some $c\in[a,b]$ such that
$\int_{x\in D}1/f(x)\,\mu(dx)=1/c\cdot\mu(D).$
Proof.
Define
$c=\frac{\mu(D)}{\int_{x\in D}1/f(x)\,\mu(dx)}.$
Since $f(x)\in[a,b]\subseteq\mathbb{R}^{+}$, we have $c\in[a,b]$ satisfies the conditions. ∎
Now we are ready to show the propositions in Section 3.1.
Proof of Proposition 3.1.
Recall in the Gaussian linear model, we conditional on the explanatory variables $X$ and treat them as fixed. By Lemma D.1 and the block-orthogonality, $R_{k}$ are independent for $k=1,\ldots,K$. Our goal is to show $R=\sum_{k}R_{k}$ is stochastically larger than the sum of certain independent Bernoulli random variables. Then the proposition follows by the concentration laws.
Part 1. By Lemma D.1 and the first order condition, $R_{k}\geq 1$ if and only if there exist some $j\in B_{k}$ such that $|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda$. We have
$\mathbb{P}(R_{k}\geq 1)=\mathbb{P}(\bigcup_{j\in B_{k}}\left\{|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda\right\})\geq\mathbb{P}(|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda)\geq 2(1-\Phi(\lambda/\sigma)).$
The last inequality is because that $\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ is a Gaussian random variable with variance $\sigma^{2}$ (note $\boldsymbol{X}$ is normalized). Direct calculation shows the probability achieves minimum when $\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ has mean zero.
As a consequence, $R=\sum_{k=1}^{K}R_{k}$ is stochastically larger than $\sum_{k=1}^{K}A_{k}$, where $A_{k}$ are $K$ independent Bernoulli random variables with success probability $\delta$. By the multiplicative Chernoff bound, we have
$\mathbb{P}(R\leq Kq)\;\leq\;\mathbb{P}\left(\sum_{k=1}^{K}A_{k}\leq\left(1-\frac{\delta-q}{\delta}\right)K\delta\right)\;\leq\;\exp\left(-\frac{(\delta-q)^{2}}{2\delta}\cdot K\right).$
Part 2. To avoid triviality, assume $K\xi$ is an integer. Let $k$ denote any one of the $K\xi$ blocks that contains a variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$. Note $\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ is Gaussian with mean $\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}$. By a similar reasoning as in Part 1,
$\mathbb{P}(R_{k}\geq 1)\;\geq\;\mathbb{P}(|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda)\;\geq\;\frac{1}{2}.$
As a consequence, $R=\sum_{k=1}^{K}R_{k}$ is stochastically larger than $\sum_{k=1}^{K\xi}A_{k}$, where $A_{k}$ are $K\xi$ independent Bernoulli random variables with success probability $1/2$. By the multiplicative Chernoff bound, we have
$\mathbb{P}(R\leq Kq)\;\leq\;\mathbb{P}\left(\sum_{k=1}^{K\xi}A_{k}\leq\left(1-\frac{\xi-2q}{\xi}\right)\frac{K\xi}{2}\right)\;\leq\;\exp\left(-\frac{(\xi-2q)^{2}}{4\xi}\cdot K\right).$
Proof of Theorem 3.1.
Notice in (2), if we scale both $\boldsymbol{Y}$ and $\lambda$ by $1/\sigma$, the optimal solution $\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}}$ will be scaled by $1/\sigma$ as well, but the selected variables do not change. Hence in the proof, we assume $\sigma^{2}=1$ without loss of generality.
Let $\boldsymbol{X}_{B_{k}}$ be the columns of $\boldsymbol{X}$ whose indices are in $B_{k}$ and $\boldsymbol{X}_{-B_{k}}$ be its complement. Let $\boldsymbol{X}_{B_{k}-j}$ be the same as $\boldsymbol{X}_{B_{k}}$ but with $\boldsymbol{X}_{j}$ excluded.
Under Assumption 3.1,
$\widehat{\textnormal{FDR}}^{*}_{j}$ $=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{|\mathcal{R}\cup\left\{j\right\}|}\mid\boldsymbol{S}_{j}\right]=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}_{k}\right\}}{|\mathcal{R}_{k}\cup\left\{j\right\}|+\sum_{l\neq k}R_{l}}\mid\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2}\right],$
where $r_{j}^{2}=\textnormal{RSS}+u_{j}^{2}$ and $u_{j}$ is the norm of the projection of $\boldsymbol{Y}$ on to the direction of $(\boldsymbol{I}-\boldsymbol{X}_{-j}(\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{-j})^{-1}\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}})\boldsymbol{X}_{j}$. By Lemma D.1, $\sum_{l\neq k}R_{l}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ and $\mathcal{R}_{k}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$. The conditional expectation $\widehat{\textnormal{FDR}}^{*}_{j}$ is an one-dimensional integration over $u_{j}$. To avoid confusion, we use $u_{j}$ to denote the value corresponding to the observed $\boldsymbol{Y}$ and $v_{j}$ to denote a placeholder for $u_{j}$ in the integral. We write
$\widehat{\textnormal{FDR}}^{*}_{j}=\int_{\left\{v_{j}:\;j\in\mathcal{R}_{k}\right\}}\frac{1}{f(v_{j};\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})}\,\mu_{j}(dv_{j}\mid\boldsymbol{S}_{j}),$
where
$f(v_{j};\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})=|\mathcal{R}_{k}\cup\left\{j\right\}|+\sum_{l\neq k}R_{l}$
and
$\mu_{j}(A\mid\boldsymbol{S}_{j})=\mathbb{P}_{H_{j}}(v_{j}\in A\mid\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})=\mathbb{P}_{H_{j}}(v_{j}\in A\mid\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2}).$
We suppress $\boldsymbol{S}_{j}$ and write $\mu_{j}(A)$ for simplicity of notation. Note the measure $\mu_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})$. Note
$f(v_{j})\in\left[1+\sum_{l\neq k}R_{l},\;b+\sum_{l\neq k}R_{l}\right]\subseteq\left[1\vee\left(-b+\sum_{k=1}^{K}R_{k}\right),\;b+\sum_{k=1}^{K}R_{k}\right].$
By Lemma D.2, for some $\gamma_{j}\in[-b,b]$, we have
$\widehat{\textnormal{FDR}}^{*}_{j}=\frac{\mu_{j}(j\in\mathcal{R}_{k})}{1\vee(R+\gamma_{j})}.$
Therefore
$\left|\widehat{\textnormal{FDR}}^{*}_{j}-\frac{\mu_{j}(j\in\mathcal{R}_{k})}{1\vee R}\right|=\left|\frac{\mu_{j}(j\in\mathcal{R}_{k})}{1\vee(R+\gamma_{j})}-\frac{\mu_{j}(j\in\mathcal{R}_{k})}{1\vee R}\right|\leq\frac{b}{1\vee R^{2}}\cdot\frac{1\vee R}{1\vee(R-b)}\leq\frac{4b^{2}}{1\vee R^{2}}.$
For notational convenience, we suppress taking maximum with $1$ for $R$. Through out the section, we have $R$ represents $1\vee R$ and $\sum_{k}R_{k}$ represents $1\vee\sum_{k}R_{k}$. Then we can write
$\widehat{\textnormal{FDR}}^{*}_{j}=\frac{\mu_{j}(j\in\mathcal{R}_{k})}{\sum_{k}R_{k}}+O\left(\frac{b^{2}}{R^{2}}\right).$
Define
$T_{k}=\sum_{j\in B_{k}}\mu_{j}(j\in\mathcal{R}_{k})\cdot\frac{\psi_{j}(p_{j})}{\mathbb{E}[\psi_{j}(p_{j})\mid\boldsymbol{S}_{j}]}.$
We have
$\sum_{j\in B_{k}}\widehat{\textnormal{FDR}}_{j}=\frac{T_{k}}{\sum_{k}R_{k}}+O\left(\frac{c_{\phi}b^{3}}{R^{2}}\right)$
and
$\widehat{\textnormal{FDR}}=\sum_{k=1}^{K}\sum_{j\in B_{k}}\widehat{\textnormal{FDR}}_{j}=\frac{\sum_{k}T_{k}}{\sum_{k}R_{k}}+O\left(\frac{Kc_{\phi}b^{3}}{R^{2}}\right)$
Note we can compute $(\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})$ from $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$ for any $j\in B_{k}$. Moreover, by Lemma D.1, for any $j\in B_{k}$, $p_{j}$ depends on $Y$ only through $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$. So $T_{k}$ depends on $Y$ only through $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$. Therefore, condition on RSS, both $\left\{T_{k}:\;k\in[K]\right\}$ and $\left\{R_{k}:\;k\in[K]\right\}$ are sets of conditional independent random variables respectively. We denote $T=\sum_{k}T_{k}$ for simplicity.
Next we prove the theorem in three steps.
Step 1. $\textnormal{Var}\left({\sum_{k}T_{k}}/{\sum_{k}R_{k}}\mid\textnormal{RSS}\right)\leq C(c_{\phi}^{2}b^{4}/Kq^{4}+c_{\phi}^{2}b^{2}K^{2}\rho)$ almost surely for some $C>0$ that does not depend on RSS.
[…]
$\textnormal{Var}_{\textnormal{RSS}}\left[\frac{T}{R}\right]=$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}$
$=$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R,R^{\prime}>Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Kq)$
$+\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R\wedge R^{\prime}\leq Kq),$
where $(T^{\prime},R^{\prime})$ is an IID copy of $(T,R)$ conditional on RSS. We will control each term separately. First,
$\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R,R^{\prime}>Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Kq)$
$=$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{TR^{\prime}-T^{\prime}R}{RR^{\prime}}\right)^{2}\mid R,R^{\prime}>Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Kq)$
$\leq$ $\frac{1}{2K^{4}q^{4}}\cdot\mathbb{E}_{\textnormal{RSS}}\left[\left(TR^{\prime}-T^{\prime}R\right)^{2}\mid R,R^{\prime}>Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Kq)$
$\leq$ $\frac{1}{2K^{4}q^{4}}\cdot\mathbb{E}_{\textnormal{RSS}}\left[\left(TR^{\prime}-T^{\prime}R\right)^{2}\right]$
$=$ $\frac{1}{2K^{4}q^{4}}\cdot\left(2\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]-2\left(\mathbb{E}_{\textnormal{RSS}}^{2}\left[TR\right]-\mathbb{E}_{\textnormal{RSS}}^{2}\left[T\right]\mathbb{E}_{\textnormal{RSS}}^{2}\left[R\right]\right)\right)$
$\leq$ $\frac{1}{K^{4}q^{4}}\cdot\left(\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]+2\mathbb{E}_{\textnormal{RSS}}\left[T\right]\mathbb{E}_{\textnormal{RSS}}\left[R\right]\cdot|\text{Cov}_{\textnormal{RSS}}(T,R)|\right)$
The last inequality is obtained by considering cases where $\text{Cov}_{\textnormal{RSS}}(T,R)$ is positive or negative separately. Note $T$ and $R^{\prime}$ are conditionally independent given RSS. We have
$\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]=\textnormal{Var}_{\textnormal{RSS}}(T)\textnormal{Var}_{\textnormal{RSS}}(R^{\prime})+\textnormal{Var}_{\textnormal{RSS}}(T)(\mathbb{E}_{\textnormal{RSS}}R^{\prime})^{2}+\textnormal{Var}_{\textnormal{RSS}}(R^{\prime})(\mathbb{E}_{\textnormal{RSS}}T)^{2}$
Note $T_{k}\in[0,c_{\phi}b]$, we have $\mathbb{E}_{\textnormal{RSS}}[T]\leq Kc_{\phi}b$ and $\textnormal{Var}_{\textnormal{RSS}}[T]=\sum_{k}\textnormal{Var}_{\textnormal{RSS}}[T_{k}]\leq Kc_{\phi}^{2}b^{2}$. Similarly, $\mathbb{E}_{\textnormal{RSS}}[R^{\prime}]\leq Kb$ and $\textnormal{Var}_{\textnormal{RSS}}[R^{\prime}]\leq Kb^{2}$. Thus
$\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]\leq K^{2}c_{\phi}^{2}b^{4}+K^{3}c_{\phi}^{2}b^{4}=O(b^{4}K^{3})$
and
$|\text{Cov}_{\textnormal{RSS}}(T,R)|\leq\sqrt{\textnormal{Var}_{\textnormal{RSS}}[T]\textnormal{Var}_{\textnormal{RSS}}[R]}\leq Kb^{2}c_{\phi}.$
Therefore the first term is bounded by $O(b^{4}/K)$. For the second term,
$\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Kq\right]$
$\leq$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}+\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Kq\right]$
$\leq$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(T+T^{\prime}\right)^{2}\mid R\wedge R^{\prime}\leq Kq\right]$
$\leq$ $2K^{2}c_{\phi}^{2}b^{2}.$
The last inequality is because $T\leq Kc_{\phi}b$. By Proposition 3.1, we have
$\mathbb{P}_{\textnormal{RSS}}(R\wedge R^{\prime}\leq Kq)\leq 2\mathbb{P}_{\textnormal{RSS}}(R\leq Kq)\leq 2\rho.$
Combine the results above, we have
$\textnormal{Var}_{\textnormal{RSS}}\left[\frac{T}{R}\right]\leq c_{\phi}^{2}\cdot\left[\frac{K^{2}b^{4}+5K^{3}b^{4}}{K^{4}q^{4}}+2K^{2}b^{2}\cdot 2\rho\right].$
The bound is of $O\left({b^{4}}/{K}\right)$ if $\rho=O(K^{-3})$.
Step 2. $\textnormal{Var}\left[\mathbb{E}\left({\sum_{k}T_{k}}/{\sum_{k}R_{k}}\mid\textnormal{RSS}\right)\right]=O(1/(n-d))$.
We start by bounding the perturbation of $\mu_{j}(j\in\mathcal{R}_{k})$ when we perturb RSS while holding $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ fixed. We write
[…]
where $r_{j}=\sqrt{\textnormal{RSS}+u_{j}^{2}}$, $I_{j}(v_{j};X_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})=\mathbf{1}\left\{j\in\mathcal{R}_{k}\right\}$, and $q_{j}=d\mu_{j}/dv_{j}$ is the probability density. Direct calculation shows
$q_{j}(v_{j})=c_{n-d}\cdot\sqrt{n-d}\cdot\left(1-\frac{v_{j}^{2}}{r_{j}^{2}}\right)^{\frac{n-d-2}{2}}\cdot\frac{1}{r_{j}},\quad\text{where}\quad c_{n-d}=\frac{\Gamma\left(\frac{n-d+1}{2}\right)}{\sqrt{(n-d)\pi}\,\Gamma\left(\frac{n-d}{2}\right)}.$
Note that $q_{j}(r_{j})=q_{j}(-r_{j})=0$. By the bounded convergence theorem,
$\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{k})=\int_{-r_{j}}^{r_{j}}I_{j}(v_{j};X_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})\cdot\partial_{\textnormal{RSS}}\,q_{j}(v_{j})\,dv_{j},$
where
$\partial_{\textnormal{RSS}}\,q_{j}=((n-d-1)v_{j}^{2}-r_{j}^{2})\cdot\frac{c_{n-d}\sqrt{n-d}}{2}\cdot(r_{j}^{2}-v_{j}^{2})^{\frac{n-d-4}{2}}\cdot(r_{j}^{2})^{-\frac{n-d+1}{2}}.$
Note that $\partial_{\textnormal{RSS}}\,q_{j}\geq 0$ if and only if $v_{j}^{2}\geq\frac{r_{j}^{2}}{n-d-1}$ and
[…]
Since $I_{j}(v_{j};X_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})\in\left\{0,1\right\}$, we have
$|\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{k})|$ $\leq\max\left\{\int_{r_{j}^{2}\geq v_{j}^{2}\geq\frac{r_{j}^{2}}{n-d-1}}\partial_{\textnormal{RSS}}\,q_{j}\,dv_{j},\;-\int_{v_{j}^{2}\leq\frac{r_{j}^{2}}{n-d-1}}\partial_{\textnormal{RSS}}\,q_{j}\,dv_{j}\right\}$
$=\int_{r_{j}^{2}\geq v_{j}^{2}\geq\frac{r_{j}^{2}}{n-d-1}}\partial_{\textnormal{RSS}}\,q_{j}\,dv_{j}$
$=c_{n-d}\cdot\frac{\sqrt{n-d}}{\sqrt{n-d-1}}\left(1-\frac{1}{n-d-1}\right)^{\frac{n-d-2}{2}}\cdot\frac{1}{r_{j}^{2}}$
$\leq\frac{1}{\textnormal{RSS}}.$
The last equation is calculated by Mathematica (2019) and the last inequality is by $c_{n-d}\leq 1/2$ and $r_{j}^{2}\geq\textnormal{RSS}$.
Then let us look at the effect of perturbing RSS on the two-sided t-test $p$-values. We have
$p_{j}=2\cdot F_{t_{n-d}}\left(-\frac{|v_{j}|}{\sqrt{\textnormal{RSS}/(n-d)}}\right),$
where $F_{t_{n-d}}$ is the cumulative distribution function of the t-distribution of degree of freedom $n-d$. Then
$\partial_{\textnormal{RSS}}\,p_{j}=c_{n-d}\cdot\left(1+\frac{v_{j}^{2}}{\textnormal{RSS}}\right)^{-\frac{n-d+1}{2}}\cdot\frac{|v_{j}|}{\sqrt{\textnormal{RSS}/(n-d)}}\cdot\frac{1}{\textnormal{RSS}}.$
Now consider the event $A_{\textnormal{RSS}}=\left\{0.5\,(n-d)\leq\textnormal{RSS}\leq 1.5\,(n-d)\right\}$. Recall we assume $\sigma^{2}=1$. Since $\textnormal{RSS}\sim\chi^{2}_{n-d}$ is sub-exponential, we have $\mathbb{P}(A^{\mathsf{c}})\leq 2\exp(-(n-d)/32)$ by the Bernstein concentration equalities.
On $A_{\textnormal{RSS}}$, we have
$|\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{k})|\leq\frac{2}{n-d}$
and
$\partial_{\textnormal{RSS}}\,p_{j}$ $\leq c_{n-d}\cdot\left(1+\frac{v_{j}^{2}/3}{(n-d)/2}\right)^{-\frac{n-d}{2}}\cdot|v_{j}|\cdot\sqrt{2}\cdot\frac{1}{n-d}$
$\leq c_{n-d}\cdot\frac{|v_{j}|}{1+v_{j}^{2}/3}\cdot\sqrt{2}\cdot\frac{1}{n-d}$
$\leq c_{n-d}\cdot\frac{\sqrt{3}}{2}\cdot\sqrt{2}\cdot\frac{1}{n-d}$
$\leq\frac{1}{n-d}.$
Note both $\mu_{j}(j\in\mathcal{R}_{k})$ and $\phi_{j}(p_{j})$ are bounded. Therefore, $\mu_{j}(j\in\mathcal{R}_{k})\cdot\phi_{j}(p_{j})$ is Lipschitz in RSS with a Lipschitz constant
$\left(2c_{\phi}+L_{\psi}\right)\cdot\frac{1}{n-d}.$
Moreover, $\sum_{k}T_{k}=\sum_{j=1}^{Kb}\mu_{j}(j\in\mathcal{R}_{k})\cdot\phi_{j}(p_{j})$ is then Lipschitz in RSS with a Lipschitz constant $Kb\left(2c_{\phi}+L_{\psi}\right)/(n-d)$. A key observation is that ${\sum_{k}T_{k}}/{R}$ is Lipschitz in RSS since $R$ is not a function of RSS.
Now consider the event $A_{R}=\left\{R>Kq\right\}$. Since RSS is independent of $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$, $A_{\textnormal{RSS}}$ and $A_{R}$ are independent. In addition, we have $\mathbb{P}(A_{R}^{\mathsf{c}})\leq\rho$.
On $A_{\textnormal{RSS}}\cap A_{R}$, we have ${\sum_{k}T_{k}}/{R}$ is $L$-Lipschitz in RSS with
$L=\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)\cdot\frac{1}{n-d}.$
Therefore, for any $\textnormal{RSS}\in A_{\textnormal{RSS}}$,
$\mathbb{E}\left[\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS},A_{R}\right]$
is $L$-Lipschitz in RSS.
For notational convenience, we denote
$E_{R}(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS},A_{R}\right],\quad E_{R^{\mathsf{c}}}(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS},A_{R}^{\mathsf{c}}\right],$
and
$E(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS}\right]=E_{R}(\textnormal{RSS})\cdot\mathbb{P}(A_{R})+E_{R^{\mathsf{c}}}(\textnormal{RSS})\cdot\mathbb{P}(A_{R}^{\mathsf{c}}).$
Note all $E_{R}(\textnormal{RSS})$, $E_{R^{\mathsf{c}}}(\textnormal{RSS})$, and $E(\textnormal{RSS})$ are bounded by $Kc_{\phi}b$.
Let $\textnormal{RSS}^{\prime}$ be an independent copy of RSS. Then
$(E(\textnormal{RSS})-E(\textnormal{RSS}^{\prime}))^{2}$
$=$ $(E_{R}(\textnormal{RSS})-E_{R}(\textnormal{RSS}^{\prime}))^{2}\cdot\mathbb{P}(A_{R})^{2}+(E_{R^{\mathsf{c}}}(\textnormal{RSS})-E_{R^{\mathsf{c}}}(\textnormal{RSS}^{\prime}))^{2}\cdot\mathbb{P}(A_{R}^{\mathsf{c}})^{2}+$
$\quad 2(E_{R}(\textnormal{RSS})-E_{R}(\textnormal{RSS}^{\prime}))(E_{R^{\mathsf{c}}}(\textnormal{RSS})-E_{R^{\mathsf{c}}}(\textnormal{RSS}^{\prime}))\cdot\mathbb{P}(A_{R})\mathbb{P}(A_{R}^{\mathsf{c}})$
$=$ $(E_{R}(\textnormal{RSS})-E_{R}(\textnormal{RSS}^{\prime}))^{2}+O(c_{\phi}^{2}b^{2}\cdot K^{2}\rho)$
Therefore,
$\textnormal{Var}[E(\textnormal{RSS})]=$ $\frac{1}{2}\mathbb{E}\left[(E(\textnormal{RSS})-E(\textnormal{RSS}^{\prime}))^{2}\right]$
$=$ $\frac{1}{2}\mathbb{E}\left[(E(\textnormal{RSS})-E(\textnormal{RSS}^{\prime}))^{2}\mid A_{\textnormal{RSS}}\right]\cdot\mathbb{P}(A_{\textnormal{RSS}})+$
$\quad\frac{1}{2}\mathbb{E}\left[(E(\textnormal{RSS})-E(\textnormal{RSS}^{\prime}))^{2}\mid A_{\textnormal{RSS}}^{\mathsf{c}}\right]\cdot\mathbb{P}(A_{\textnormal{RSS}}^{\mathsf{c}})$
$=$ $\frac{1}{2}\mathbb{E}\left[(E_{R}(\textnormal{RSS})-E_{R}(\textnormal{RSS}^{\prime}))^{2}\mid A_{\textnormal{RSS}}\right]\cdot\mathbb{P}(A_{\textnormal{RSS}})$
$\quad+O(c_{\phi}^{2}b^{2}\cdot K^{2}\rho)+O\left(c_{\phi}^{2}b^{2}\cdot K^{2}e^{-{(n-d)}/{32}}\right)$
Since $E_{R}(\textnormal{RSS})$ is $L$-Lipschitz in RSS,
$\frac{1}{2}\mathbb{E}\left[(E_{R}(\textnormal{RSS})-E_{R}(\textnormal{RSS}^{\prime}))^{2}\mid A_{\textnormal{RSS}}\right]\cdot\mathbb{P}(A_{\textnormal{RSS}})$
$\leq$ $\frac{1}{2}\mathbb{E}\left[L^{2}(\textnormal{RSS}-\textnormal{RSS}^{\prime})^{2}\mid A_{\textnormal{RSS}}\right]\cdot\mathbb{P}(A_{\textnormal{RSS}})$
$\leq$ $\frac{1}{2}\mathbb{E}\left[L^{2}(\textnormal{RSS}-\textnormal{RSS}^{\prime})^{2}\right]$
$\leq$ $L^{2}\textnormal{Var}(\textnormal{RSS})$
$=$ $\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{2}{n-d},$
where the last line uses the fact that $(n-d)\textnormal{RSS}\sim\chi^{2}(n-d)$ and hence $\textnormal{Var}(\textnormal{RSS})=2/(n-d)$. Then we have
$\textnormal{Var}\left[E(\textnormal{RSS})\right]\leq\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}b^{2}\cdot K^{2}\rho)+O\left(c_{\phi}^{2}b^{2}\cdot K^{2}e^{-(n-d)/32}\right).$
That is
$\textnormal{Var}\left[\mathbb{E}\left(\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS}\right)\right]\leq\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}b^{2}\cdot K^{2}(\rho\vee e^{-(n-d)/32})).$
Combining Step 1 and Step 2, we have
$\textnormal{Var}\left[\frac{\sum_{k}T_{k}}{R}\right]$ $=\mathbb{E}\left[\textnormal{Var}_{\textnormal{RSS}}\left(\frac{T}{R}\right)\right]+\textnormal{Var}\left[\mathbb{E}\left(\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS}\right)\right]$
$\leq O\left(\frac{c_{\phi}^{2}b^{4}}{Kq^{4}}\right)+\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}b^{2}\cdot K^{2}(\rho\vee e^{-(n-d)/32}))$
Finally,
$\textnormal{Var}\left[\widehat{\textnormal{FDR}}\right]=$ $\textnormal{Var}\left[\frac{\sum_{k}T_{k}}{\sum_{k}R_{k}}+O\left(\frac{Kc_{\phi}b^{3}}{R^{2}}\right)\right]$
$\leq$ $2\textnormal{Var}\left[\frac{\sum_{k}T_{k}}{\sum_{k}R_{k}}\right]+O\left(2\mathbb{E}\left[\frac{K^{2}c_{\phi}^{2}b^{6}}{R^{4}}\right]\right)$
$\leq$ $2\textnormal{Var}\left[\frac{\sum_{k}T_{k}}{\sum_{k}R_{k}}\right]+O\left(\frac{c_{\phi}^{2}b^{6}}{K^{2}q^{4}}\right)+O(K^{2}c_{\phi}^{2}b^{4}\rho)$
$\leq$ $O\left(c_{\phi}^{2}\left(\frac{b^{4}}{q^{4}}+\frac{b^{6}}{Kq^{4}}\right)\cdot\frac{1}{K}\right)+\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{4}{n-d}$
$+O(c_{\phi}^{2}b^{4}\cdot K^{2}(\rho\vee e^{-(n-d)/32})).$
The big-O notation only hides universal constant.
Proof of Corollary 3.1.
Result (1): The setting satisfies the conditions in Part 2 of Proposition 3.1 with $\xi=K^{-\varepsilon}$ and $q=\xi/3=O(K^{-\varepsilon})$. Then $\rho=\exp\left(-K^{1-\varepsilon}/36\right)$. Plugging them into the bound in Theorem 3.1, we have
$\textnormal{Var}(\widehat{\textnormal{FDR}})=O\left(K^{-1+4\varepsilon}+\frac{K^{2\varepsilon}}{n-d}+K^{2}\exp\left(-K^{1-\varepsilon}/36\right)\right)=O(K^{-1+4\varepsilon}).$
Result (2): The setting satisfies the conditions in Part 2 of Proposition 3.1 with $q=\xi/3$, and $\rho=\exp\left(-K\xi/36\right)$. Plugging them into the bound in Theorem 3.1, we have
$\textnormal{Var}(\widehat{\textnormal{FDR}})=O\left(K^{-1}+\frac{1}{n-d}+K^{2}\exp\left(-(n-d)/32\right)\right)=O(K^{-1}).$
Appendix E Extended numerical results
Figure 13 shows the extended results of $\widehat{\textnormal{FDR}}$ along with estimated one-standard-error bars in Section 4.1 and Figure 14 shows the extended results for Section 4.2.
(a) APV, $201$, $767$
(b) ATV, $147$, $328$
(c) IDV, $206$, $825$
(d) LPV, $148$, $515$
(e) NFV, $207$, $842$
(f) RTV, $205$, $793$
(g) SQV, $206$, $824$
(h) X3TC, $283$, $629$
(i) ABC, $283$, $623$
(j) AZT, $283$, $626$
(k) D4T, $281$, $625$
(l) DDI, $283$, $628$
(m) TDF, $215$, $351$
(n) DLV, $305$, $730$
(o) EFV, $312$, $732$
(p) NVP, $313$, $744$
Figure 13: Extended results of $\widehat{\textnormal{FDR}}$ along with estimated one-standard-error bars in Section 4.1. The string and number in the caption of the subfigures represent the name of the drug, the number of variables $d$, and the number of observations $n$, respectively.
Figure 14: Extended results of $\widehat{\textnormal{FDR}}$ along with estimated one-standard-error bars in Section 4.2. From top-left to bottom-right are stimulation settings 1 to 14.
Appendix F Running time of $\widehat{\textnormal{FDR}}$ in simulation problems
Figure 15 shows a box plot of the running time of evaluating $\widehat{\textnormal{FDR}}$ (without estimating its s.e. by bootstrap) at all $10$ different $\lambda$ in the the simulation problems in Section 5 on a single-cored computer. The runtime is typically several minutes.
Figure 15: Box plot of the running time of evaluating $\widehat{\textnormal{FDR}}$ (without estimating its s.e. by bootstrap) at all $10$ different $\lambda$ in the simulation problems in Section 5 on a single-cored computer.

### v4

@@
-Estimating the FDR of variable selection
+Estimating the False Discovery Rate of Variable Selection
 Abstract
@@
 1 Introduction
-When deciding which variables to include in a statistical model, a data analyst commonly wants to balance multiple goals including predictive accuracy, accurate detection of signal variables, parsimony, and interpretability.
-Perhaps the best-known paradigm is to use a variable-selecting method such as the Lasso (Tibshirani, 1996), forward stepwise regression (Hocking, 1976), or the graphical Lasso (Friedman et al., 2008), with a tuning parameter that controls the model complexity, and select the value of the tuning parameter by minimizing the cross-validation (CV) estimate of out-of-sample prediction error.
-This paradigm is versatile but prioritizes prediction error above all else, while ignoring the other goals mentioned above.
-In particular, CV can be quite liberal in including irrelevant explanatory variables (see e.g.
-Baumann, 2003), an undesirable feature in applications where accurate variable selection is a priority.
-In such applications, we could abandon estimation methods like the Lasso and instead apply multiple testing methods such as knockoffs (Barber and Candès, 2016; Candès et al., 2018), or the Benjamini–Hochberg procedure or its dependence-corrected variants (Benjamini and Hochberg, 1995; Benjamini and Yekutieli, 2001; Fithian and Lei, 2022), which guarantee control of the false discovery rate (FDR) at a user-specified level.
-However, multiple testing methods do not yield fitted models, and the model defined by including only the selected variables can have very poor fit to the data.
-In this work we introduce a new approach for assessing variable selection accuracy of estimation methods like the Lasso: an estimator for the FDR of any variable selection procedure, which can be evaluated alongside the cross-validation curve to inform the analyst about the tradeoff between variable selection accuracy and prediction accuracy over the entire regularization path.
-Our estimator can be applied in common statistical settings including the Gaussian linear model, the Gaussian graphical model, and the nonparametric model-X framework of Candès et al.
-(2018), requiring no additional statistical assumptions in these settings beyond those required for standard hypothesis tests.
-Its bias is non-negative in finite samples and typically small, and its standard error can be assessed using the bootstrap.
-In the special case of thresholding independent $p$-values for multiple hypothesis tests, our method is almost identical to the FDR estimation method advocated by Storey (2002), but slightly less conservative.
+When selecting variables for inclusion in a statistical model, data analysts typically seek to balance multiple goals, including predictive accuracy, reliable identification of signal variables, model parsimony, and interpretability.
+Perhaps the best-known paradigm is to use a variable selection method such as the Lasso (Tibshirani, 1996), forward stepwise regression (Hocking, 1976), or the graphical Lasso (Friedman et al., 2008),in conjunction with a tuning parameter that governs model complexity.
+The optimal value of this tuning parameter is commonly chosen by minimizing the cross-validation (CV) estimate of out-of-sample prediction error.
+Although this framework is flexible and effective, it inherently prioritizes prediction accuracy over other desirable properties.
+In particular, CV-based selection can be quite liberal in including irrelevant explanatory variables (see, e.g., Baumann, 2003), which is undesirable in applications where accurate variable selection is of primary importance.
+In such applications, one may forgo estimation-based approaches like the Lasso in favor of multiple testing procedures, such as knockoffs (Barber and Candès, 2016; Candès et al., 2018), or the Benjamini–Hochberg procedure and its dependence-corrected variants (Benjamini and Hochberg, 1995; Benjamini and Yekutieli, 2001; Fithian and Lei, 2022), which guarantee control of the false discovery rate (FDR) at a user-specified level.
+However, multiple testing methods do not yield fitted models, and the model formed by including only the selected variables may exhibit very poor fit to the data.
+[^4] For example, in the knockoff procedure, certain correlation structures can cause a strong signal variable to have a negative knockoff statistic with approximately 50% probability (Li and Fithian, 2021).
+If this occurs, the strong sigal variable will not be selected regardless of how large the FDR threshold is.
+In this work, we introduce a novel approach for assessing variable selection accuracy of estimation methods such as the Lasso: an estimator of the FDR applicable to any variable selection procedure under common statistical settings.
+This estimator can be evaluated in conjunction with the cross-validation curve, enabling analysts to examine the trade-off between variable selection accuracy and predictive performance across the entire regularization path.
+The proposed method is applicable in common statistical settings, including the Gaussian linear model, the Gaussian graphical model, and the nonparametric model-X framework of Candès et al.
+(2018), without imposing any additional assumptions beyond those required for standard hypothesis testing.
+The estimator exhibits non-negative finite-sample bias that is typically small, and its standard error can be assessed via the bootstrap.
+The idea of estimating the FDR of a selection procedure dates back to Storey (2002), who proposed a conservative estimator for variable selection based on thresholding independent $p$-values.
+Goeman and Solari (2011) further developed an upper confidence bound for the number of null variables in any given selection set, yielding a conservative median estimator.
+Their method relies on a closed testing framework, and a feasible implementation typically requires independent or specific positively correlated $p$-values.
+In addition, certain multiple testing procedures, such as Benjamini–Hochberg (Benjamini and Hochberg, 1995) and knockoffs (Barber and Candès, 2016), incorporate intrinsic False Discovery Proportion (FDP) estimators to achieve FDR control based on their particular variable ordering and testing strategy.
+In contrast, our proposed FDR estimator is broadly applicable across common statistical settings and can be used with any variable selection procedure deemed appropriate by the analyst.
 (a) Easy variable selection scenario
@@
 Figure 1: Cross-validation MSE (blue), true FDR (black), and our FDR estimator (red), for Lasso regression in two scenarios with explanatory variables that are (a) independent, and (b) highly correlated.
-In scenario (a) the minimum-CV model (solid blue vertical line) has high FDR, while the one-standard-error rule (dashed vertical line) achieves a reasonably low FDR.
-In scenario (b) there is no model that simultaneously achieves good predictive performance and low FDR.
+In scenario (a), the minimum-CV model (solid blue vertical line) has high FDR, while the one-standard-error rule (dashed vertical line) achieves a reasonably low FDR.
+In scenario (b), there is no model that simultaneously achieves good predictive performance and low FDR.
 In both scenarios, our FDR estimator successfully captures important information about variable selection performance that is not evident from the CV curve.
-Figure 1 illustrates why our method is useful by contrasting two Lasso regression scenarios with very different tradeoffs between prediction error and FDR.
-In both scenarios, we generate $n=600$ data points from a Gaussian linear model with $d=200$ standard Gaussian explanatory variables, of which only $d_{1}=20$ are true signal variables with the same moderate signal strength.
-The only difference is that the explanatory variables are independent in the first scenario, but correlated in the second (specifically, they follow $\text{AR}(0.8)$ correlation).
-This difference dramatically changes the variable selection performance: When the explanatory variables are independent, it is possible to select a good predictive model if we use the one-standard-error rule (but not if we minimize CV error).
-When the explanatory variables are correlated, however, all models with reasonably good predictive error have FDR exceeding $40\%$.
-While the CV curve gives no indication of the variable selection performance in either scenario, our estimator is accurate in both.
-The next section explains how we calculate it.
+Figure 1 illustrates the utility of our proposed method by comparing two Lasso regression scenarios that exhibit markedly different trade-offs between prediction error and FDR.
+In both scenarios, we generate $n=600$ observations from a Gaussian linear model with $d=200$ standard Gaussian explanatory variables, among which only $d_{1}=20$ are signal variables of equal moderate strength.
+The sole difference between the two settings lies in the dependence structure of the explanatory variables: in the first scenario, the variables are independent, whereas in the second, they exhibit an autoregressive $\text{AR}(1)$ correlation structure with a correlation coefficient $0.8$.
+This distinction substantially alters the variable selection performance.
+When the explanatory variables are independent, a model exhibiting both high predictive accuracy and reliable variable selection can be obtained using the one-standard-error rule, though not by minimizing the CV error.
+In contrast, when the variables are correlated, all models achieving competitive predictive accuracy exhibit an FDR exceeding $40\%$.
+While the CV curve alone offers no insight into variable selection performance in either case, our proposed estimator remains accurate across both scenarios.
+The following section introduces its construction.
 1.1 FDR estimation in the Gaussian linear model
@@
 Benjamini and Hochberg (1995) define the false discovery rate (FDR) as the expectation of the false discovery proportion (FDP), defined here as the fraction of selected variables that are actually irrelevant.
-If $\mathcal{H}_{0}$ is the index set of all noise variables and $R=|\mathcal{R}|$ is the total number of selections, we have
+Let $\mathcal{H}_{0}$ be the index set of all noise variables and $R=|\mathcal{R}|$ be the total number of selections, we have
 $\textnormal{FDR}=\mathbb{E}[\textnormal{FDP}]\quad\textnormal{and}\quad\textnormal{FDP}=\frac{|\mathcal{R}\cap\mathcal{H}_{0}|}{R},$ (3)
-using the convention that $/0=0$ 0; i.e., $\textnormal{FDP}=0$ when no variables are selected.
-Note that we can linearly decompose the FDR as
+using the convention that $0/0=0$; i.e., $\textnormal{FDP}=0$ when no variables are selected.
+Inspired by Fithian and Lei (2022), we linearly decompose the FDR as
 $\textnormal{FDR}=\sum_{j=1}^{d}\textnormal{FDR}_{j},\quad\textnormal{FDR}_{j}=\mathbb{E}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{|\mathcal{R}|}\right]\cdot\mathbf{1}\left\{j\in\mathcal{H}_{0}\right\}.$ (4)
-Using this decomposition, a promising estimator of the FDR is
+This technique allows us to treat each term separately and decouple the complicated interactions by conditioning. A promising estimator of the FDR is
 $\widehat{\textnormal{FDR}}=\sum_{j=1}^{d}\widehat{\textnormal{FDR}}_{j},\quad\widehat{\textnormal{FDR}}_{j}=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{|\mathcal{R}|}\mid\boldsymbol{S}_{j}\right]\cdot\frac{\mathbf{1}\left\{p_{j}>0.1\right\}}{0.9},$ (5)
@@
 The bias of $\mathbb{E}[\widehat{\textnormal{FDR}}$] is small if most signal variables have $p_{j}\leq 0.1$ with high probability.
-The definition of our estimator $\widehat{\textnormal{FDR}}$ in (5) is highly general and makes no assumptions about the form of the variable selection procedure $\mathcal{R}$.
-As a result, it applies to any selection procedure including the Lasso, forward stepwise regression, LARS (Efron et al., 2004), the elastic net (Zou and Hastie, 2005), and many others.
-Our method requires no further assumptions beyond the standard linear model assumptions in (1) required to justify the marginal $t$-tests for each variable.
-Indeed, in the common case where the method $\mathcal{R}$ is a function of $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}$ and $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ only (which is true of all methods referenced above), the only statistical assumption required to justify the inequality (6) is that the $t$-statistic follows a $t_{n-d}$ distribution under $H_{j}$, conditional on $\boldsymbol{S}_{j}$.
-This assumption may hold approximately even when the parametric assumptions of the linear model fail.
-As we see next, our method generalizes far beyond the Gaussian linear model to other parametric and nonparametric contexts, provided that our modeling assumptions supply an appropriate sufficient statistic $\boldsymbol{S}_{j}$ for each null hypothesis $H_{j}$.
-Section 2 defines our method generically and discusses additional applications including the graphical Lasso (Friedman et al., 2008) and nonparametric model-X assumptions introduced by (Candès et al., 2018).
-Section 3 discusses theoretical analysis of the standard error in stylized settings, as well as our bootstrap scheme for estimating the standard error, which is used to generate the error bars in all figures, and Sections 4 and 5 illustrate our method’s behavior in real data and simulation examples, respectively.
-Section 6 discusses computational issues, and Section 7 concludes.
+The definition of our estimator $\widehat{\textnormal{FDR}}$ in (5) is highly general and imposes no assumptions on the form of the variable selection procedure $\mathcal{R}$.
+Consequently, it is applicable to any selection method, including the Lasso, forward stepwise regression, LARS (Efron et al., 2004), the elastic net (Zou and Hastie, 2005), among many others.
+Our approach requires no assumptions beyond those of the standard linear model in (1), which are required to justify the use of marginal $t$-tests for each variable.
+In the common setting where the procedure $\mathcal{R}$ depends only on $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}$ and $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$–a property shared by all methods listed above–the sole statistical assumption needed to justify the inequality (6) is that the $t$-statistic follows a $t_{n-d}$ distribution under $H_{j}$, conditional on $\boldsymbol{S}_{j}$.
+Notably, this assumption may still hold approximately even when the parametric assumptions of the linear model are violated.
+As we show next, our method extends far beyond the Gaussian linear model to other parametric and nonparametric settings, provided that the modeling assumptions yield an appropriate sufficient statistic $\boldsymbol{S}_{j}$ for each null hypothesis $H_{j}$.
+Section 2 describes our method in full generality and discusses additional applications, including the graphical Lasso (Friedman et al., 2008) and the nonparametric model-X framework of Candès et al.
+(2018).
+Section 3 presents a theoretical analysis of the standard error in stylized settings and describes our bootstrap procedure for estimating the standard error, which is used to construct the error bars in all figures.
+Sections 4 and 5 illustrate the behavior of our estimator in real-data and simulation studies, respectively.
+Section 6 addresses computational considerations, and Section 7 concludes.
 2 General formulation of our estimator
 2.1 Notation and statistical assumptions
-Throughout the paper, we will use upper-case letters to represent variables and lower-case letters to denote individual observations.
-For example, $y_{i}$ is the $i$th observation of variable $Y$ and $x_{i,j}$ is the $i$th observation of variable $X_{j}$.
-We use the upper-case bold-face letter to denote a vector of all $n$ observations, e.g.
-$\boldsymbol{X}_{j}=(x_{1,j},\ldots,x_{n,j})^{\mkern-1.5mu\mathsf{T}}$.
-We use $[d]$ to denote the set $\{1,\ldots,d\}$.
-For $B\subseteq[d]$, we denote as $\boldsymbol{X}_{B}$ the block of matrix $\boldsymbol{X}$ including only the variables in $B$, and we use $\boldsymbol{X}_{-j}$ as a shorthand for $\boldsymbol{X}_{[d]\setminus\{j\}}$ and $\boldsymbol{X}_{-B}$ for $\boldsymbol{X}_{[d]\setminus B}$.
-We denote by $\boldsymbol{D}$ a generic random object that encodes the entire data set, whose distribution $P$ we will assume lies in some known family $\mathcal{P}$.
-If $\mathcal{P}$ is a parametric family, we generically use $\boldsymbol{\theta}$ to denote its parameters, and $P_{\boldsymbol{\theta}}$ to denote the corresponding distribution.
-In the Gaussian linear model in Section 1.1, irrelevance of explanatory variable $X_{j}$ is represented by the null hypothesis $H_{j}:\boldsymbol{\theta}_{j}=0$, but other problem settings will involve other null hypotheses.
+Throughout the paper, we use uppercase letters to represent random variables and lowercase letters to denote individual observations.
+For example, $y_{i}$ denotes the $i$th observation of the variable $Y$, and $x_{i,j}$ denotes the $i$th observation of the variable $X_{j}$.
+We use bold uppercase letters to denote the collection of all $n$ observations.
+For instance, $\boldsymbol{X}_{j}=(x_{1,j},\ldots,x_{n,j})^{\mkern-1.5mu\mathsf{T}}$ represents the vector of observations for the random variable $X_{j}$, and $\boldsymbol{X}\in\mathbb{R}^{n\times d}$, whose $(i,j)$th element is $x_{i,j}$, denotes the matrix of observations for all explanatory variables.
+Moreover, let $d]={1,\ldots,d}$.
+For any subset $A\subseteq[d]$, we denote by $\boldsymbol{X}_{A}$ the submatrix of $\boldsymbol{X}$ containing only the columns corresponding to the variables in $A$.
+Similarly, we use $\boldsymbol{X}_{-j}$ as shorthand for $\boldsymbol{X}_{[d]\setminus{j}}$ and $\boldsymbol{X}_{-A}$ for $\boldsymbol{X}_{[d]\setminus A}$.
+We denote by $\boldsymbol{D}$ a generic random object representing the entire dataset, whose distribution $P$ is assumed to belong to a known family $\mathcal{P}$.
+If $\mathcal{P}$ is a parametric family, we use $\boldsymbol{\theta}$ to denote its parameters and write $P_{\boldsymbol{\theta}}$ for the corresponding distribution.
+In the Gaussian linear model discussed in Section 1.1, the irrelevance of explanatory variable $X_{j}$ is expressed through the null hypothesis $H_{j}:\boldsymbol{\theta}_{j}=0$, while other problem settings may involve different forms of null hypotheses.
+Moreover, we denote by $\mathcal{R}$ the index set of variables selected by a generic variable selection procedure based on the random dataset $\boldsymbol{D}$, and define $R=|\mathcal{R}|$ as the total number of selected variables.
 To define our estimator in a generic modeling context and prove that its bias is non-negative, the sole statistical assumption we will rely on is the availability of a sufficient statistic $\boldsymbol{S}_{j}$ for the submodel corresponding to each null hypothesis $H_{j}$.
@@
 To estimate each variable’s contribution $\textnormal{FDR}_{j}$, we proceed by estimating each of the two factors in turn.
-Step 1: Estimating the first factor:
+Step 1: Estimating the first factor.
 Crucially, note that the factor $\mathbb{E}[\mathbf{1}\{j\in\mathcal{R}\}/R]$ is only relevant when $j\in\mathcal{H}_{0}$.
-As a result, we can assume when estimating it that $H_{j}$ holds.
-Since $\boldsymbol{S}_{j}$ is sufficient for the submodel $H_{j}$, we can do so by Rao-Blackwellizing the integrand (Rao, 1945; Blackwell, 1947), obtaining
+Therefore, when estimating it, we may assume that $H_{j}$ holds.
+Since $\boldsymbol{S}_{j}$ is sufficient for the submodel under $H_{j}$, we can estimate the first factor by Rao-Blackwellizing the integrand (Rao, 1945; Blackwell, 1947), obtaining
 $\widehat{\textnormal{FDR}}^{*}_{j}(\boldsymbol{S}_{j}):=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right].$ (7)
 For $j\in\mathcal{H}_{0}$, $\widehat{\textnormal{FDR}}_{j}^{*}$ is unbiased by construction, and it is the UMVU estimator for the first factor of $\textnormal{FDR}_{j}$ whenever $\boldsymbol{S}_{j}$ is complete (Lehmann and Scheffé, 1950).
-Step 2: Estimating the second factor:
+Since the conditional null distribution of the data $\boldsymbol{D}$ given $\boldsymbol{S}_{j}$ is known by the sufficiency, $\widehat{\textnormal{FDR}}^{*}_{j}$ is computable and can generally be approximated using a Monte Carlo algorithm.
+For certain variable selection procedures, we have developed fast and exact computational methods.
+Further details are provided in Section 6.
+Step 2: Estimating the second factor.
 We could obtain a conservative estimator by simply upper-bounding the indicator $\mathbf{1}\{j\in\mathcal{H}_{0}\}\leq 1$, resulting in the estimator $\widehat{\textnormal{FDR}}^{*}=\sum_{j}\widehat{\textnormal{FDR}}_{j}^{*}$.
@@
 $\phi_{j}(\boldsymbol{D})=\frac{\psi_{j}(\boldsymbol{D})}{\mathbb{E}_{H_{j}}[\psi_{j}(\boldsymbol{D})\mid\boldsymbol{S}_{j}]},$ (8)
-or $\phi_{j}(\boldsymbol{D})=1$ if $\mathbb{E}_{H_{j}}[\psi_{j}|\boldsymbol{S}_{j}]=0$. Note we have $\mathbb{E}_{H_{j}}[\phi_{j}\mid\boldsymbol{S}_{j}]=1$ almost surely.
-Commonly there is a standard test for $H_{j}$, whose $p$-value $p_{j}$ is uniform and independent of $\boldsymbol{S}_{j}$ under $H_{j}$.
+or $\phi_{j}(\boldsymbol{D})=1$ if $\mathbb{E}_{H_{j}}[\psi_{j}|\boldsymbol{S}_{j}]=0$.
+Note we have $\mathbb{E}_{H_{j}}[\phi_{j}\mid\boldsymbol{S}_{j}]=1$ almost surely.
+For the same reason as in Step 1, the normalizer $\mathbb{E}_{H_{j}}[\psi_{j}(\boldsymbol{D})\mid\boldsymbol{S}_{j}]$ is always computable.
+As an example, commonly, there is a standard test for $H_{j}$, whose $p$-value $p_{j}$ is uniform and independent of $\boldsymbol{S}_{j}$ under $H_{j}$.
 Then, we can take $\psi_{j}=\mathbf{1}\{p_{j}>\zeta\}$ for $\zeta\in(0,1)$, and
 $\phi_{j}(p_{j})=\frac{\mathbf{1}\{p_{j}>\zeta\}}{1-\zeta}.$ (9)
-If the signal is strong enough for the test to reject $H_{j}$ with high probability at level $\zeta$, then variable $j$’s contribution will almost always be zero.
-We adopt (9) as the canonical form of $\phi_{j}$, using $\zeta=0.1$ for all simulations.
+The parameter $\zeta$ governs the bias–variance trade-off: a larger $\zeta$ leads to a smaller bias but a larger variance.
+Our numerical experiments indicate that $\zeta=0.1$ provides a good balance between bias and variance.
+When the signal is strong enough for the test to reject $H_{j}$ with high probability at level $\zeta$, the contribution of variable $j$ will almost always be zero.
+In Appendix A.3, we establish an upper bound on the bias contributed by a strong-signal variable in the Gaussian linear model, showing that it decays exponentially with the signal strength.
+We adopt (9) as the canonical form of $\phi_{j}$.
 Combining (7) and (8), we obtain the estimator
 $\widehat{\textnormal{FDR}}=\sum_{j=1}^{d}\widehat{\textnormal{FDR}}_{j},\quad\text{where }\widehat{\textnormal{FDR}}_{j}(\boldsymbol{D})=\widehat{\textnormal{FDR}}_{j}^{*}(\boldsymbol{S}_{j})\cdot\phi_{j}(\boldsymbol{D}).$ (10)
-We can write the canonical version of $\widehat{\textnormal{FDR}}_{j}$ more concretely as
+When we adopt the canonical $\phi_{j}$, we can write the canonical version of $\widehat{\textnormal{FDR}}_{j}$ more concretely as
 $\widehat{\textnormal{FDR}}_{j}(\boldsymbol{D})=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]\cdot\frac{\mathbf{1}\{p_{j}>\zeta\}}{1-\zeta}.$ (11)
-These two factors are independent if $p_{j}\perp\!\!\!\perp\boldsymbol{S}_{j}$, as is commonly the case.
+These two factors are independent when $p_{j}\perp\!\!\!\perp\boldsymbol{S}_{j}$.
+[^5] When a suitable $p_{j}$ that is uniform and independent of $\boldsymbol{S}_{j}$ under $H_{j}$ is not available to adopt the canonical form, researchers may instead choose an appropriate $\psi_{j}$ function and employ the general formulation (10) of $\widehat{\textnormal{FDR}}$.
 The estimator derived above, in its generic form (10), is always conservatively biased:
 Theorem 2.1.
-For any selection procedure $\mathcal{R}$, and for any estimator $\psi_{j}:\;\boldsymbol{D}\to[0,1]$ of $\mathbf{1}\{j\in\mathcal{H}_{0}\}$, the estimator $\widehat{\textnormal{FDR}}$ defined in (710) has non-negative bias:
+Suppose for each $j=1,\ldots,d$, $\boldsymbol{S}_{j}$ is a sufficient statistic under the null submodel $H_{j}$.
+Then for any selection procedure $\mathcal{R}$, and for any estimator $\psi_{j}:\;\boldsymbol{D}\to[0,1]$ of $\mathbf{1}\{j\in\mathcal{H}_{0}\}$, the estimator $\widehat{\textnormal{FDR}}$ defined in (7, 8, 10) has non-negative bias:
 $\mathbb{E}[\widehat{\textnormal{FDR}}]\geq\textnormal{FDR}.$
@@
 In other words, the sparsity pattern of the undirected dependence graph $\mathcal{G}$ for $X$ is identical to the sparsity pattern of $\boldsymbol{\Theta}$.
-In lieu of estimating all $d(d-1)/2$ off-diagonal entries of $\boldsymbol{\Theta}$, a common modeling assumption is that the dependence graph is sparse, and by extension that $\boldsymbol{\Theta}$ is sparse as well (Meinshausen and Bühlmann, 2006; Yuan and Lin, 2007; Friedman et al., 2008).
+In lieu of estimating all d(d1)/2d(d-1)/2 off-diagonal entries of $\boldsymbol{\Theta}$, a common modeling assumption is that the dependence graph is sparse, and by extension that $\boldsymbol{\Theta}$ is sparse as well (Meinshausen and Bühlmann, 2006; Yuan and Lin, 2007; Friedman et al., 2008).
 We associate to each pair $j\neq k$ the null hypothesis that $(j,k)$ is absent from the true dependence graph $\mathcal{G}$, or equivalently $H_{jk}:\;\boldsymbol{\Theta}_{jk}=0$.
@@
 Since our general method can be adapted for many different selection algorithms in many different statistical modeling contexts, it is very difficult to provide a general analysis that covers all contexts in which our method might be applied, but we can offer theoretical results in a narrower context, as well as a practical estimation strategy that we have found to be reasonably reliable in our simulation studies.
-Section 3.1 develops a theoretical finite-sample bound for our estimator under stylized assumptions, leading to conditions under which the variance of $\widehat{\textnormal{FDR}}$ is vanishing, while Section 3.2 discusses a specialized bootstrap method for standard error estimation, which we use to produce all figures in this work.
+Section 3.1 develops a theoretical finite-sample bound for our estimator under stylized assumptions, leading to conditions under which the variance of $\widehat{\textnormal{FDR}}$ is vanishing.
+Section 3.2 discusses a specialized bootstrap method for standard error estimation, which we use to produce all figures in this work.
+Section 3.3 introduces the theoretical property of our bootstrap scheme.
 3.1 Theoretical bound for standard error
@@
 Assumption 3.1 (block-orthogonal variables).
-Let $\boldsymbol{X}$ be an $n\times d$ matrix, and let $B_{1},\ldots,B_{K}\subseteq[d]$ represent a partition of $[d]$ into $K$ disjoint subsets (that is, $\bigcup_{k}B_{k}=[d]$).
-We say the matrix $\boldsymbol{X}$ has block orthogonal variables with respect to the partition $B$ if $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{B_{\ell}}=\boldsymbol{0}_{|B_{k}|\times|B_{\ell}|}$, for all $k\neq\ell$.
-Assumption 3.1 allows us to treat the Lasso on the entire matrix $\boldsymbol{X}$ as a separable optimization problem on the $K$ different blocks of explanatory variables.
-When this assumption is in effect, we will denote the size of the largest block as $b=\max_{k}|B_{k}|$.
+Let $\boldsymbol{X}$ be an $n\times d$ matrix, and let index set $(b)\subseteq[d]$ for $b=1,\ldots,B$ represent a partition of $[d]$ into $B$ disjoint subsets (that is, $\bigcup_{b}(b)=[d]$).
+For each $b$, let $\boldsymbol{X}_{(b)}\in\mathbb{R}^{n\times m_{b}}$ be the submatrix of $\boldsymbol{X}$ containing only the columns corresponding to the variables in $(b)$, where $m_{b}=|(b)|$ is size of block $(b)$.
+We say that $\boldsymbol{X}$ has block orthogonal variables with respect to the partition $\{(b)\}_{b=1}^{B}$ if
+$\boldsymbol{X}_{(b_{1})}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{(b_{2})}=\boldsymbol{0}_{m_{b_{1}}\times m_{b_{2}}},\quad\text{for all }b_{1}\neq b_{2}.$
+Assumption 3.1 allows us to treat the Lasso on the entire matrix $\boldsymbol{X}$ as a separable optimization problem on the $B$ different blocks of explanatory variables.
+When this assumption is in effect, we will denote the size of the largest block as $m=\max_{b}m_{b}$.
 We will assume throughout this section that every explanatory variable has mean zero and unit norm: $\sum_{i=1}^{n}x_{i,j}=0$ and $\|\boldsymbol{X}_{j}\|^{2}=1$, for all $j=1,\ldots,d$.
 This last assumption comes with little practical loss of generality since conventional implementations of the Lasso begin by normalizing all variables in this way.
-For our results below, it will be important to be able to provide a probabilistic lower bound on the average number $q$ of selections per block:
+For our results below, it will be important to be able to provide a probabilistic lower bound $q$ on the average number of selections per block:
 Condition 3.1 (enough selections).
-We have $\mathbb{P}(R\leq Kq)\leq\rho$ for some $q,\rho\geq 0$.
+We have $\mathbb{P}(R\leq Bq)\leq\rho$ for some $q,\rho\geq 0$.
 Because Condition 3.1 always holds for some $q$ and $\rho$, it has force only after we show that it holds for specific values of $q$ and $\rho$, defined in terms of other problem quantities.
@@
   1. 1.
-    Let $\delta(z):=2(1-\Phi(z))$ denote the probability that a Gaussian random variable with mean 00 and variance 1 exceeds $z$ in absolute value.
-Then $R$ is stochastically larger than a $\textnormal{Binom}(K,\delta(\lambda/\sigma))$ random variable, and for any $q<\delta(\lambda/\sigma)$, the Lasso procedure satisfies Condition 3.1 with
-    $\rho=\exp\left(-\;\frac{(\delta(\lambda/\sigma)-q)^{2}}{2\delta}\cdot K\right).$
+    Let $\delta(z):=2(1-\Phi(z))$ denote the probability that a Gaussian random variable with mean 0 and variance 1 exceeds $z$ in absolute value.
+Then $R$ is stochastically larger than a $\textnormal{Binom}(B,\delta(\lambda/\sigma))$ random variable, and for any $q<\delta(\lambda/\sigma)$, the Lasso procedure satisfies Condition 3.1 with
+    $\rho=\exp\left(-\;\frac{(\delta(\lambda/\sigma)-q)^{2}}{2\delta}\cdot B\right).$
   2. 2.
-    Assume no fewer than $K\xi$ blocks contain at least one variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$ for some $\xi\in(0,1]$, where $\boldsymbol{\mu}=\boldsymbol{X}\boldsymbol{\theta}$ is the mean of $\boldsymbol{Y}$.
-Then $R$ is stochastically larger than a $\textnormal{Binom}(K,\xi/2)$ random variable, and for any $q<\xi/2$, the Lasso procedure satisfies Condition 3.1 with
-    $\rho=\exp\left(-\;\frac{(\xi-2q)^{2}}{4\xi}\cdot K\right).$
-A finite-sample upper bound for $\text{Var}(\widehat{\textnormal{FDR}})$ for Lasso with block-orthogonal variables then follows.
+    Assume no fewer than $B\xi$ blocks contain at least one variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$ for some $\xi\in(0,1]$, where $\boldsymbol{\mu}=\boldsymbol{X}\boldsymbol{\theta}$ is the mean of $\boldsymbol{Y}$.
+Then $R$ is stochastically larger than a $\textnormal{Binom}(B\xi,1/2)$ random variable, and for any $q<\xi/2$, the Lasso procedure satisfies Condition 3.1 with
+    $\rho=\exp\left(-\;\frac{(\xi-2q)^{2}}{4\xi}\cdot B\right).$
+A finite-sample upper bound $\text{Var}(\widehat{\textnormal{FDR}}) for$ Lasso with block-orthogonal variables then follows.
 Theorem 3.1 (finite sample variance bound).
@@
 If Condition 3.1 holds for given values of $q$ and $\rho$, we have
-$\textnormal{Var}(\widehat{\textnormal{FDR}})\leq O\left(\left(\frac{b^{4}}{q^{4}}+\frac{b^{6}}{Kq^{4}}\right)\cdot\frac{1}{K}+\frac{b^{2}}{q^{2}}\cdot\frac{1}{n-d}+b^{4}K^{2}(\rho\vee e^{-(n-d)/32})\right),$
+$\textnormal{Var}(\widehat{\textnormal{FDR}})=O\left(\left(\frac{m^{4}}{q^{4}}+\frac{m^{6}}{Bq^{4}}\right)\cdot\frac{1}{B}+\frac{m^{2}}{q^{2}}\cdot\frac{1}{n-d}+m^{6}B^{2}(\rho\vee e^{-(n-d)/32})\right),$
 where the big-O notation only hides constants that depend on $c_{\phi}$, $L_{\phi}$, and $\sigma^{2}$.
@@
 As a result, resampling the data conditional on $\boldsymbol{S}_{j}$ would not change the selection in all other blocks.
-Hence, the number of selections is conditionally stable: with resampling conditional on $\boldsymbol{S}_{j}$ it is $R(\boldsymbol{D})+O(b)$.
+Hence, the number of selections is conditionally stable: with resampling conditional on $\boldsymbol{S}_{j}$ it is $R(\boldsymbol{D})+O(m)$.
   2. 2.
-    For any $j\in B_{k}$, let $X_{B_{k}-j}$ denotes the variables in block $B_{k}$ except $X_{j}$.
-The selection of variable $X_{j}$ depends on $\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
+    For any $j\in(b)$, let $\boldsymbol{X}_{(b);-j}$ denote the variables in block $(b)$ except $\boldsymbol{X}_{j}$.
+The selection of variable $\boldsymbol{X}_{j}$ depends on $\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ only through $\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
 Formally, we have
-    $E_{j}:=\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}]=\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}]\approx\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{S}_{j}].$
+    $E_{j}:=\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}]=\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}]\approx\mathbb{E}_{H_{j}}[\mathbf{1}\left\{j\in\mathcal{R}\right\}\mid\boldsymbol{S}_{j}].$
     Note these $E_{j}$ are block-independent.
 [^7] Alert readers might have noticed that conditioning on $\boldsymbol{S}_{j}$ is not the same as conditioning on $\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$: there is an additional $\left\|\boldsymbol{Y}\right\|$ term.
-The analysis of the effects of $\left\|\boldsymbol{Y}\right\|$ is rather technical and it is the source of the $O(1/(n-d))$ term in the variance bound.
+The analysis of the effects of $\left\|\boldsymbol{Y}\right\|$ is rather technical and it is the source of the O(1/(nd))O(1/(n-d)) term in the variance bound.
   3. 3.
-    Likewise, the p-value $p_{j}$ depends on $\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
+    Likewise, the p-value $p_{j}$ depends on $\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ only through $\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
 So $\phi_{j}(p_{j})$ are block-independent.
 As a result, we have
-$\widehat{\textnormal{FDR}}_{j}\;=\;\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]\cdot\phi_{j}\;\approx\;\frac{E_{j}\cdot\phi_{j}}{R+O(b)}\;\approx\;\frac{E_{j}\cdot\phi_{j}}{R}.$
-The last approximation holds when $R>Kq$ with high probability and $K$ is substantially larger than $b$. Then
-$\widehat{\textnormal{FDR}}\;=\;\sum_{j}\widehat{\textnormal{FDR}}_{j}\;\approx\;\frac{\sum_{k}\left(\sum_{j\in B_{k}}E_{j}\cdot\phi_{j}\right)}{R}.$
-This is an average of $K$ independent random variables. Therefore, the variance is of order $O(1/K)$ by the concentration laws. A formal asymptotic result is shown in Corollary 3.1.
+$\widehat{\textnormal{FDR}}_{j}\;=\;\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]\cdot\phi_{j}\;\approx\;\frac{E_{j}\cdot\phi_{j}}{R+O(m)}\;\approx\;\frac{E_{j}\cdot\phi_{j}}{R}.$
+The last approximation holds when $R>Bq$ with high probability and $B$ is substantially larger than $m$. Then
+$\widehat{\textnormal{FDR}}\;=\;\sum_{j}\widehat{\textnormal{FDR}}_{j}\;\approx\;\frac{\sum_{k}\left(\sum_{j\in(b)}E_{j}\cdot\phi_{j}\right)}{R}.$
+This is an average of $B$ independent random variables. Therefore, the variance is of order $O(1/B)$ by the concentration laws. A formal asymptotic result is shown in Corollary 3.1.
 Corollary 3.1 (Asymptotically vanishing variance).
-Under the same settings as in Theorem 3.1, consider a series of problems where we hold $c_{\phi}$, $b$, $L_{\psi}$ and $\sigma$ fixed but allow $\boldsymbol{\theta}$, $\lambda$, and $n$ to vary as $K$ increases.
-Suppose that at least $s(K)$ blocks contain a variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$, where $\boldsymbol{\mu}=\boldsymbol{X}\boldsymbol{\theta}$ is the mean of $\boldsymbol{Y}$.
+Under the same settings as in Theorem 3.1, consider a series of problems where we hold $c_{\phi}$, $m$, $L_{\phi}$ and $\sigma$ fixed but allow $\boldsymbol{\theta}$, $\lambda$, and $n$ to vary as $B$ increases.
+Suppose that at least $s(B)$ blocks contain a variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$, where $\boldsymbol{\mu}=\boldsymbol{X}\boldsymbol{\theta}$ is the mean of $\boldsymbol{Y}$.
   1. 1.
-    For $40<\varepsilon<1/$ 4, suppose $s(K)\geq K^{1-4\varepsilon}$, and let $K\to\infty$ with $n-d\geq K^{1-2\varepsilon}$. Then
-    $\textnormal{Var}(\widehat{\textnormal{FDR}})=O(K^{-1+4\varepsilon})\to 0.$
+    For $0<\varepsilon<1/4$, suppose $s(B)\geq B^{1-4\varepsilon}$, and let $B\to\infty$ with $n-d\geq B^{1-2\varepsilon}$. Then
+    $\textnormal{Var}(\widehat{\textnormal{FDR}})=O(B^{-1+4\varepsilon})\to 0.$
   2. 2.
-    For $10<\xi<$ 1, suppose $s(K)\geq K\xi$, and let $K\to\infty$ with $n-d\geq 64\log(K)$. Then
-    $\textnormal{Var}(\widehat{\textnormal{FDR}})=O(K^{-1})\to 0.$
+    For $0<\xi<1$, suppose $s(B)\geq B\xi$, and let $B\to\infty$ with $n-d\geq 64\log(B)$. Then
+    $\textnormal{Var}(\widehat{\textnormal{FDR}})=O(B^{-1})\to 0.$
 […]
@@
 […]
-To ameliorate this issue, we try to construct $\widehat{P}$ in such a way that most of the null variables in $P$ remain null in $\widehat{P}$, but the signal variables are estimated with little bias.
-Formally, let $\widehat{\mathcal{H}_{0}}$ denote an estimate of the index set of noise variables.
-In the parametric setting $\boldsymbol{D}\sim P_{\boldsymbol{\theta}}$, we find a sparse MLE of $\boldsymbol{\theta}$, denoted as $\hat{\boldsymbol{\theta}}$, under the constraints that $\theta_{j}=0$ for all $j\in\widehat{\mathcal{H}_{0}}$.
-Then we set $\widehat{P}=P_{\hat{\boldsymbol{\theta}}}$ and draw bootstrap samples to calculate $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$.
-We formally describe our bootstrap scheme in the parametric setting in Algorithm 1.
+[…]
 Algorithm 1 $\widehat{\textnormal{FDR}}$ standard error estimation (parametric)
@@
 Compute: the MLE under the intersection null model with $\theta_{j}=0$ for all $j\in\widehat{\mathcal{H}_{0}}$
-$\hat{\boldsymbol{\theta}}(\widehat{\mathcal{H}_{0}};\boldsymbol{D})=\underset{\theta_{j}=0,\,j\in\widehat{\mathcal{H}_{0}}}{\operatorname*{argmax}}\;P_{\boldsymbol{\theta}}(\boldsymbol{D})$
+$\hat{\boldsymbol{\theta}}^{(*)}(\widehat{\mathcal{H}_{0}};\boldsymbol{D})=\underset{\theta_{j}=0,\,j\in\widehat{\mathcal{H}_{0}}}{\operatorname*{argmax}}\;P_{\boldsymbol{\theta}}(\boldsymbol{D})$
 Loop: for $m=1,2,\ldots,M$ do
-  generate sample $\hat{\boldsymbol{D}}^{(m)}\sim P_{\hat{\boldsymbol{\theta}}}$;  compute $\widehat{\textnormal{FDR}}^{(m)}=\widehat{\textnormal{FDR}}(\hat{\boldsymbol{D}}^{(m)})$
+  generate sample $\hat{\boldsymbol{D}}^{(m)}\sim P_{\hat{\boldsymbol{\theta}}^{(*)}}$;  compute $\widehat{\textnormal{FDR}}^{(m)}=\widehat{\textnormal{FDR}}(\hat{\boldsymbol{D}}^{(m)})$
 Output: Sample standard error
@@
 Example 3.1 (Gaussian linear model).
-In the Gaussian linear model, the negative log-likelihood $-\log P_{\theta}(\boldsymbol{D})$ is proportional to the mean-squared error.
-Therefore, $\boldsymbol{\hat{\theta}}(\widehat{\mathcal{H}_{0}})$ is the least square coefficients regressing $Y$ on variables $X_{j}$ for $j\in\widehat{\mathcal{H}_{0}}^{\mathsf{c}}$.
+[…]
 Example 3.2 (Gaussian graphical model).
@@
 Yuan and Lin, 2007), we apply a simple heuristic algorithm described in Appendix A.4 to save computation.
-For the model-X setting, we can employ an analogous method by using a variant of the i.i.d.
-bootstrap that constrains $H_{j}:\;Y\perp\!\!\!\perp X_{j}\mid X_{-j}$ to be true for all $j\in\widehat{\mathcal{H}_{0}}$.
-Specifically, we first draw bootstrap samples of $(Y,X_{-\widehat{\mathcal{H}_{0}}})$ from the empirical distribution of the data.
-Then, we sample $X_{\widehat{\mathcal{H}_{0}}}$ from the conditional distribution $X_{\widehat{\mathcal{H}_{0}}}\mid(Y,X_{-\widehat{\mathcal{H}_{0}}})$, which is equal to $X_{\widehat{\mathcal{H}_{0}}}\mid X_{-\widehat{\mathcal{H}_{0}}}$ under the intersection null $\bigcap_{j\in\widehat{\mathcal{H}_{0}}}H_{j}$.
-We formally describe our bootstrap scheme in the model-X nonparametric setting in Algorithm 2.
-If the researchers make model assumptions, the set $\widehat{\mathcal{H}_{0}}$ can be constructed in the same way as in the parametric settings.
-If no model assumption is made and the likelihood is unavailable, researchers can choose $\widehat{\mathcal{H}_{0}}$ via p-value-based approaches.
-For example, we can calculate the conditional randomization test p-values (Candès et al., 2018) for each variable and define $\widehat{\mathcal{H}_{0}}$ as all variables whose p-values are larger than $0.1$.
-This choice lets the majority of null variables remain null in $\widehat{P}$ while retaining the signal for strong signal variables.
 Algorithm 2 $\widehat{\textnormal{FDR}}$ standard error estimation (non-parametric, model-X)
 Input: The observed data matrix $\boldsymbol{D}$ whose $n$ rows are independent observations $(y_{i},x_{i,1},\ldots,x_{i,d})$; an index set $\widehat{\mathcal{H}_{0}}$ that we believe $H_{j}:\,\theta_{j}=0$ is true for $j\in\widehat{\mathcal{H}_{0}}$.
-Preparation: For simplicity, dentoe $x_{i,\widehat{\mathcal{H}_{0}}}=(x_{i,j},\;j\in\widehat{\mathcal{H}_{0}})$ as the $i$th observation of $X_{\widehat{\mathcal{H}_{0}}}$.
-Let $\widehat{P}_{-}$ denote the empirical distribution of $(Y,X_{-\widehat{\mathcal{H}_{0}}})$
-$\widehat{P}_{-}(\,\cdot\,)=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{(y_{i},x_{i,-\widehat{\mathcal{H}_{0}}})\in\,\cdot\,\right\}.$
+[…]
+[…]
 Loop: for $m=1,2,\ldots,M$ do
@@
 $\widehat{\textnormal{s.e.}}=\sqrt{\frac{1}{M-1}\sum_{m=1}^{M}\left(\widehat{\textnormal{FDR}}^{(m)}-\frac{1}{M}\sum_{l=1}^{M}\widehat{\textnormal{FDR}}^{(l)}\right)^{2}}$
+[…]
+If the researchers make model assumptions, the set $\widehat{\mathcal{H}_{0}}$ can be constructed in the same way as in the parametric settings.
+If no model assumption is made and the likelihood is unavailable, researchers can choose $\widehat{\mathcal{H}_{0}}$ via p-value-based approaches.
+For example, we can calculate the conditional randomization test p-values (Candès et al., 2018) for each variable and define $\widehat{\mathcal{H}_{0}}$ as all variables whose p-values are larger than $0.1$.
+This choice lets the majority of null variables remain null in $\widehat{P}$ while retaining the signal for strong signal variables.
+3.3 Theory of parametric bootstrap variance estimation
+It is difficult to study the theoretical property of the bootstrap variance estimator in full generality.
+Nevertheless, we make substantial progress for Gaussian linear models with Lasso selection.
+In particular, we revisit the stylized block-orthogonal Gaussian linear model used in our theoretical variance analysis (Section 3.1) with known error variance $\sigma^{2}$ for simplicity and consider the Lasso selection with a appropriately chosen penalty level in an asymptotic regime where the true and estimated FDR are bounded away from 0 and $1$ in the limit.
+For variance estimation, we consider the parametric bootstrap in Algorithm 1 with $M=\infty$, which assumes away Monte Carlo error, and the estimated null set
+$\widehat{\mathcal{H}_{0}}=\{j:\ \widehat{\theta}^{\textnormal{Lasso}}_{j}(\bar{\lambda})=0\},$
+where $\bar{\lambda}>\lambda$ is a slightly inflated penalty level used only for estimating the bootstrap model.
+We then choose $\hat{\boldsymbol{\theta}}^{(*)}$ as the OLS estimator restricting to $\widehat{\mathcal{H}_{0}}$, as described in Example 3.1.
+We call the resulting model an OLS-after-Lasso model.
+Define $\widehat{\textnormal{FDR}}^{(*)}$ as a generic bootstrap sample of $\widehat{\textnormal{FDR}}$ computed under OLS-after-Lasso model.
+In this setting, we prove that the asymptotic variance of the parametric bootstrap FDR estimates conditional on data matches the asymptotic variance of our FDR estimator.
+We elaborate the setting and technical details in Appendix E.1.
+Theorem 3.2.
+Consider a Gaussian linear model with known $\sigma^{2}$ that satisfies the Assumption 3.1.
+In the setting of Theorem E.2 in Appendix E.2, there exist deterministic sequences $\nu_{d}=o(1)$ and $\tilde{\alpha}_{d}$ that are bounded away from 0 and $1$, and a data-dependent sequence $\alpha_{d}^{(*)}=\alpha_{d}^{(*)}(\boldsymbol{X},\boldsymbol{Y})$ such that
+$\sup_{t\in\mathbb{R}}\biggl|\mathbb{P}\!\biggl(\frac{1}{\nu_{d}}\bigl(\widehat{\textnormal{FDR}}-\tilde{\alpha}_{d}\bigr)\leq t\biggr)-\Phi(t)\biggr|=o(1),$
+and
+$\sup_{t\in\mathbb{R}}\biggl|\mathbb{P}\!\biggl(\frac{1}{\nu_{d}}\bigl(\widehat{\textnormal{FDR}}^{(*)}-\alpha_{d}^{(*)}\bigr)\leq t\,\Big|\,\boldsymbol{X},\boldsymbol{Y}\biggr)-\Phi(t)\biggr|=o_{\mathbb{P}}(1),$
+where $\Phi$ denotes the CDF of $\mathcal{N}(0,1)$.
 4 Real world examples
@@
 In the data set for a given drug, the response $y_{i}$ is a measure of sample $i$’s resistance to that drug, and the explanatory variables are binary indicators of whether each of several hundred mutations was present or absent in that sample.
+As in Barber and Candès (2016), we adopt the Gaussian linear model assumption in this study.
 As a measure of replicability, Barber and Candès (2016) assessed whether detected genes were also present in a list of treatment-selected mutations (TSM), which a previous study had found were enriched in patients who had been treated by drugs in each of the three classes (Rhee et al., 2005).
@@
 The vertical lines show the locations of the complexity parameters selected by CV and the one-standard-error rule.
-Figure 2 shows our FDR estimates alongside the CV estimate of MSE, for four selected experiments representing a variety of statistical scenarios.
-Results for all drugs are available in Appendix E.
-The error bars for both CV and $\widehat{\textnormal{FDR}}$ our estimator show $\pm$±one standard error of $\widehat{\textnormal{FDR}}$ estimated by bootstrap.
-The $\widehat{\textnormal{FDP}}$ curves in the figures are estimated using the TSM signals as an approximation to the ground truth.
+[…]
 The four panels of Figure 2 illustrate how different variable selection problems can have very different trade-offs between variable selection accuracy and prediction accuracy.
@@
 Panel (b), by contrast, shows a regime in which no lasso estimator can simultaneously achieve competitive model fit and low FDR.
-Panel (c) shows an experiment in which the estimated FDR is non-monotone in the complexity parameter $\lambda$: both our estimator and the TSM $\widehat{\textnormal{FDP}}$ judge the fifth variable to enter the model, gene P71.V, to be a likely noise variable, because its $t$-statistic is $-0.38$, corresponding to a $p$-value of $0.70$, and it is also absent from the TSM list (the $t$-statistics for the first four variables to enter all exceed 5).
+Panel (c) shows an experiment in which the estimated FDR is non-monotone in the complexity parameter $\lambda$: both our estimator and the TSM $\widehat{\textnormal{FDP}}$ judge the fifth variable to enter the model, gene P71.V, to be a likely noise variable, because its $t$-statistic is 0.38-0.38, corresponding to a $p$-value of $0.70$, and it is also absent from the TSM list (the $t$-statistics for the first four variables to enter all exceed 5).
 Panel (d) shows one of the three experiments in which our $\widehat{\textnormal{FDR}}$ departs substantially from the TSM $\widehat{\textnormal{FDP}}$.
@@
 Figure 4 shows two representative examples of the results from running the graphical Lasso on the $14$ experimental settings.
-Results for all experiments are available in Appendix E.
+Results for all experiments are available in Appendix F.
 As in the HIV study, we show our $\widehat{\textnormal{FDR}}$ and its estimated standard error, alongside an estimate of FDP calculated using the causal DAG supplied by Sachs et al.
@@
   1. 1.
-[…]
+    IID normal: independent Gaussian random variables $X_{j}\overset{\textnormal{i.i.d.}}{\sim}\mathcal{N}(0,1)$.
   2. 2.
@@
   4. 4.
-[…]
+    Sparse: Independent sparse binary observations $X_{ij}\overset{\textnormal{i.i.d.}}{\sim}\text{Bernoulli}(0.05)$.
 The signal variable set $\mathcal{H}_{0}^{\mathsf{c}}$ is a random subset of $[d]$ that has cardinality $d_{1}$, uniformly distributed among all such subsets.
 To mimic realistic scenarios where there is a mix between weak and strong signal variables, we generate nonzero signals with random signal strength
-[…]
+$\theta_{j}\overset{\textnormal{i.i.d.}}{\sim}\theta^{*}\cdot(1+\textnormal{Exp}(1))/2,$
 for all $j\in\mathcal{H}_{0}^{\mathsf{c}}$, where $\textnormal{Exp}(1)$ represents a rate-$1$ exponentially distributed random variable and $\theta^{*}$ is a fixed value controlling the average signal strength.
@@
 […]
-Figure 5: Performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ in the Gaussian linear model.
-The error bars show the $5\%$ to $95\%$ quantiles of the empirical distributions and the solid lines show the sample average.
-$\widehat{\textnormal{FDR}}$ shows a small non-negative bias in estimating FDR and $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$ successively captures the magnitude of uncertainty of $\widehat{\textnormal{FDR}}$.
+[…]
 Next, we stress-test the performance of $\widehat{\textnormal{FDR}}$ in several variations on the IID normal scenario above.
@@
 where we employ $\ell_{1}$-penalized logistic regression for variable selection. The other settings are the same as in Section 5.1.
-Figure 7: Performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ in the model-X settings.
-The error bars show the $5\%$ to $95\%$ quantiles of the empirical distributions and the solid lines show the sample average.
-$\widehat{\textnormal{FDR}}$ shows a small non-negative bias in estimating FDR and $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$ successively captures the magnitude of uncertainty of $\widehat{\textnormal{FDR}}$.
+[…]
 Figure 7 shows the performance of our estimators where the regression problem (left panel) is high-dimensional with $d=500$ and $n=400<d$ and the classification problem (right panel) has $d=500$ and $n=1500$.
@@
 […]
-Figure 8: Performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ in the Gaussian graphical model.
-The error bars show the $5\%$ to $95\%$ quantiles of the empirical distributions and the solid lines show the sample average.
-$\widehat{\textnormal{FDR}}$ shows a small non-negative bias in estimating FDR and $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$ successively captures the magnitude of uncertainty of $\widehat{\textnormal{FDR}}$.
+[…]
 5.4 Unfavorable case: equicorrelated Gaussian
@@
 But like in previous simulations, the $\widehat{\textnormal{s.e.}}$ tells the magnitude of uncertainty of $\widehat{\textnormal{FDR}}$ and can warn the researchers that $\widehat{\textnormal{FDR}}$ is not very reliable.
-Figure 9: Performance of $\widehat{\textnormal{FDR}}$ and $\widehat{\textnormal{s.e.}}$ on the MCC problem.
-The error bars show the $5\%$ to $95\%$ quantiles of the empirical distributions and the solid lines show the sample average.
-$\widehat{\textnormal{FDR}}$ shows a large variance but $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$ successively captures the magnitude of uncertainty of $\widehat{\textnormal{FDR}}$.
+[…]
 5.5 Robustness to model misspecification
@@
 $Y=\sum_{j=1}^{d}X_{j}\theta_{j}+\exp\left(\sum_{j=1}^{d}|X_{j}\theta_{j}|\right)\cdot\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,1).$
-[…]
+In the “t-noise” setting, we have the noise $\varepsilon$ in the Gaussian linear model follow the t-distribution with a degree of freedom $5$ instead of the standard normal distribution.
+The “Exponential” setting has the variables follow the heavy-tailed exponential distribution $X_{j}\overset{\textnormal{i.i.d.}}{\sim}\textnormal{Exp(1)}$ but are mistaken as Gaussian in the model-X assumption.
+Finally, In the “X Perturbation” setting, we multiply each observed variable by an IID uniformly distributed perturbation $U([0,2])$.
+As the figure shows, our $\widehat{\textnormal{FDR}}$ performs well in these scenarios.
 6 Computational details
@@
 These calculations can be easily parallelized.
-In a generic setting, we can use Monte-Carlo integration to approximate $\widehat{\textnormal{FDR}}^{*}_{j}$.
-The worst-case time complexity for computing $\widehat{\textnormal{FDR}}$ is $O(d\cdot M\cdot T)$, where $M$ is the number of Monte-Carlo samples and $T$ is the running time of the selection procedure.
+In a general setting, $\widehat{\textnormal{FDR}}^{*}_{j}$ can be approximated via Monte Carlo integration: generate $M$ samples of the data $\boldsymbol{D}$ from the known conditional distribution, compute the integrand ${\mathbf{1}\left\{j\in\mathcal{R}\right\}}/{R}$ for each sample, and take the average over them.
+The worst-case time complexity for computing $\widehat{\textnormal{FDR}}$ is $O(d\cdot M\cdot T)$, where $T$ is the running time of the selection procedure.
 We note that calculating the integrand of $\widehat{\textnormal{FDR}}_{j}^{*}$ for each Monte-Carlo sample has the same computation complexity as calculating one placebo statistic in the conditional randomization test (CRT, Candès et al.
@@
 estimation) at all $10$ different tuning parameter values in the simulation problems in Section 5 is typically several minutes on a single-cored computer.
-We report the exact running time in Appendix F.
-7 Summary
+We report the exact running time in Appendix G.
+7 Summary and discussion
 In this work we present a new approach for estimating FDR of any variable selection procedure.
@@
 R package is available online at the Github repository https://github.com/yixiangLuo/hFDR.
+While this paper is focused on three settings, our method can be applied insofar as a sufficient statistic $\boldsymbol{S}_{j}$ exists under the null submodel $H_{j}$, i.e., the data $\boldsymbol{D}$ has a known distribution conditional on $\boldsymbol{S}_{j}$ under $H_{j}$.
+Consider a common setting in the multiple testing literature where a test statistic $T_{j}$ is available for $H_{j}$ and $T_{1},\ldots,T_{d}$ are independent.
+If $H_{j}$ are all simple hypotheses (e.g., $H_{j}:\theta_{j}=0$), we can simply set $\boldsymbol{S}_{j}=(T_{1},\ldots,T_{j-1},T_{j+1},\ldots,T_{d}\}$.
+When the $H_{j}$ are composite but simple conditional on some auxiliary statistic $V_{j}$, we can set $\boldsymbol{S}_{j}=(T_{1},\ldots,T_{j-1},V_{j},T_{j+1},\ldots,T_{d}\}$.
+For example, if the researcher has two samples $X_{j1},\ldots,X_{jn_{j_{1}}}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}P$ and $Z_{j1},\ldots,Z_{jn_{j_{2}}}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Q$ for the hypothesis $H_{j}:P=Q$, $\boldsymbol{S}_{j}$ can be always chosen as the unordered set of $\{X_{j1},\ldots,X_{jn_{j_{1}}},Z_{j1},\ldots,Z_{jn_{j_{2}}}\}$.
+If $P=F_{\theta_{1}}$ and $Q=F_{\theta_{2}}$ for some exponential family $F$ with natural parameter $\theta$ and sufficient statistic $T(\cdot)$, then we can choose $\boldsymbol{S}_{j}=\sum_{i=1}^{j_{1}}T(X_{ji})+\sum_{k=1}^{j_{2}}T(Z_{jk})$.
+Therefore, our method can also be applied to a large class of multiple testing problems with independent test statistics.
 Reproducibility
-Our FDR estimator is implemented in an R package available online at the Github repository
-github.com/yixiangLuo/hFDR. The R code to reproduce all simulations and figures is available at github.com/yixiangLuo/hFDR_expr.
+Our FDR estimator is implemented in an R package available online at the Github repository github.com/yixiangLuo/hFDR.
+The R code to reproduce all simulations and figures is available at github.com/yixiangLuo/hFDR_expr.
 Appendix A Additional details in experiments and methodologies
 A.1 Simulation settings for the illustrative examples
-[…]
+In Figure 1(a), the explanatory variables are independent Gaussian – $X_{j}\overset{\textnormal{i.i.d.}}{\sim}\mathcal{N}(0,1)$.
+We have $d=200$ variables, $n=600$ observations, and $d_{1}=20$ variables are signal variables with random signal strength $\theta_{j}=0.25\cdot(1+\textnormal{Exp}_{j}(1))/2$, where $\textnormal{Exp}_{j}(1)$ with $j\in\mathcal{H}_{0}^{\mathsf{c}}$ are independent exponentially distributed random variable with rate $1$.
+The response $Y$ is generated by the Gaussian linear model with $\sigma=1$.
 In Figure 1(b), the explanatory variables $X_{j}$ ($j=1,\ldots,d)$ form an AR(1) Gaussian random process with $\text{cov}(X_{j},X_{j+1})=0.8$. The other settings are the same as above.
 A.2 Cross-validation for constructing $\widehat{\mathcal{H}_{0}}$ in $\widehat{\textnormal{s.e.}}$
-In this section, we describe the cross-validation algorithm we use to construct the set $\widehat{\mathcal{H}_{0}}$ in computing $\widehat{\textnormal{s.e.}}(\widehat{\textnormal{FDR}})$.
-Formally, let $\boldsymbol{D}_{\text{trn},k}$ and $\boldsymbol{D}_{\text{val},k}$ be the $k$th training and validation data set in a $K$-folds cross-validation, respectively.
-Throughout the section, we write $\boldsymbol{\hat{\theta}}(\widehat{\mathcal{H}_{0}}(\lambda);\boldsymbol{D}_{\text{trn},k})$ with $\widehat{\mathcal{H}_{0}}(\lambda)=\mathcal{R}^{\mathsf{c}}(\lambda;\boldsymbol{D}_{\text{trn},k})$ as $\boldsymbol{\hat{\theta}}(\lambda;\boldsymbol{D}_{\text{trn},k})$ for simplicity of notation.
-We have
-$\boldsymbol{\hat{\theta}}(\lambda;\boldsymbol{D}_{\text{trn},k})=\underset{\theta_{j}=0,\,j\not\in\mathcal{R}(\lambda;\boldsymbol{D}_{\text{trn},k})}{\operatorname*{argmax}}\;P_{\boldsymbol{\theta}}(\boldsymbol{D}_{\text{trn},k}).$
+[…]
+$\boldsymbol{\hat{\theta}}^{(*)}(\lambda;\boldsymbol{D}_{\text{trn},k})=\underset{\theta_{j}=0,\,j\not\in\mathcal{R}(\lambda;\boldsymbol{D}_{\text{trn},k})}{\operatorname*{argmax}}\;P_{\boldsymbol{\theta}}(\boldsymbol{D}_{\text{trn},k}).$
 Then we find the best $\lambda$ that achieves the largest averaged likelihood on the validation sets
-[…]
+$\lambda_{\text{cv}}=\underset{\lambda}{\operatorname*{argmax}}\;\frac{1}{K}\sum_{k=1}^{K}P_{\boldsymbol{\hat{\theta}}^{(*)}(\lambda;\boldsymbol{D}_{\text{trn},k})}(\boldsymbol{D}_{\text{val},k})$
 and define
 $\widehat{\mathcal{H}_{0}}=\left\{j:\;j\not\in\mathcal{R}(\lambda_{\text{cv}};\boldsymbol{D})\right\}.$
-A.3 Sparse model estimation with graphical Lasso
+A.3 Bias contribution from signal variables
+Theorem A.1.
+In the Gaussian linear model (1), consider the canonical version (11) of $\widehat{\textnormal{FDR}}_{j}$:
+$\widehat{\textnormal{FDR}}_{j}=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]\cdot\frac{\mathbf{1}\{p_{j}>\zeta\}}{1-\zeta},$
+with $p_{j}$ being the two-sided t-test p-value for $H_{j}$. Define $T_{j}$ as the t-test statistic with noncentrality
+$\mu_{j}=\frac{\theta_{j}}{\sigma\sqrt{(\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1})_{jj}}}.$
+Let $t_{1-\zeta/2}$ denote the $1-\zeta/2$ upper quantile of the central t-distribution with degree of freedom ndn-d.
+Then the contribution to the bias of $\widehat{\textnormal{FDR}}$ from variable $j$, when the signal strength $|\mu_{j}|>t_{1-\zeta/2}$, has upper bound
+$\mathbb{E}[\widehat{\textnormal{FDR}}_{j}]\leq\frac{1}{(1-\zeta)}\,\left[2\exp\!\left(-\frac{\bigl(|\mu_{j}|-\sqrt{2}\,t_{1-\zeta/2}\bigr)_{+}^{2}}{2}\right)\;+\;\exp\!\left(-\frac{n-d}{16}\right)\right].$
+Proof.
+Note
+$T_{j}\;\stackrel{{\scriptstyle d}}{{=}}\;\frac{Z+\mu_{j}}{\sqrt{U/(n-d)}},\qquad\textnormal{where }Z\sim\mathcal{N}(0,1),\ U\sim\chi^{2}_{n-d},\ Z\perp U.$
+Let $A:=\{U\leq 2(n-d)\}$. Split
+$\mathbb{P}(|T|<t_{1-\zeta/2})\leq\mathbb{P}(|T|<t_{1-\zeta/2},\ A)+\mathbb{P}(A^{c}).$
+On $A$ we have $\sqrt{U/(n-d)}\leq\sqrt{2}$, hence
+$\{|T|<t_{1-\zeta/2}\}\cap A\subseteq\{|Z+\mu_{j}|<\sqrt{2}\,t_{1-\zeta/2}\}.$
+Therefore
+$\mathbb{P}(|T|<t_{1-\zeta/2})\leq\mathbb{P}(|Z+\mu_{j}|<\sqrt{2}\,t_{1-\zeta/2})+\mathbb{P}(U>2(n-d)).$ (15)
+For the Gaussian term, with $a:=\sqrt{2}\,t_{1-\zeta/2}$,
+$\mathbb{P}(|Z+\mu_{j}|<a)=\mathbb{P}(-a-\mu_{j}<Z<a-\mu_{j})\leq\mathbb{P}(Z<a-\mu_{j})\leq\exp\!\left(-\frac{(\mu_{j}-a)_{+}^{2}}{2}\right),$
+and applying the same bound to $\mu_{j}$ gives
+$\mathbb{P}(|Z+\mu_{j}|<a)\leq 2\exp\!\left(-\frac{(|\mu_{j}|-a)_{+}^{2}}{2}\right).$
+For the chi-square term, use a Chernoff bound,
+$\mathbb{P}(U>2(n-d))\leq\exp\!\left(-\frac{n-d}{16}\right).$
+Therefore
+$\mathbb{P}(p_{j}>\zeta)=\mathbb{P}(|T_{j}|<t_{1-\zeta/2})\;\leq\;2\exp\!\left(-\frac{\bigl(|\mu_{j}|-\sqrt{2}\,t_{1-\zeta/2}\bigr)_{+}^{2}}{2}\right)\;+\;\exp\!\left(-\frac{n-d}{16}\right).$
+Note
+$\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]\leq 1\qquad a.s.$
+Combining all the components together, the inequality is proved. ∎
+Theorem A.1 states that the contribution to the bias of $\widehat{\textnormal{FDR}}$ from strong signal variables decays to 0 exponentially with the signal strength. In the proof, the bound
+$\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{R}\mid\boldsymbol{S}_{j}\right]\leq 1\qquad a.s.$
+is rather loose.
+In practice, we expect $R$ to be not much smaller than the number of strong signals for a reasonable variable selection procedure.
+Hence the bound established in Theorem A.1 is also a practical estimate for the bias coming from all the strong signals.
+A.4 Sparse model estimation with graphical Lasso
 We start with the unconstrained MLE
@@
 To select the signal variables with $\theta_{j}\neq 0$, we do a simple one-sided thresholding,
-=$\mathcal{R}=\left\{j:\;Z_{j}\geq 0.5\cdot\sqrt{2\log d}\right\},$
+$\mathcal{R}=\left\{j:\;Z_{j}\geq 0.5\cdot\sqrt{2\log d}\right\},$
 and set $\psi_{j}=\mathbf{1}\left\{p_{j}>0.5\right\}$, where $p_{j}$ is the one-sided marginal z-test p-value.
 Let $\boldsymbol{S}_{j}=\boldsymbol{Z}_{-j}-\boldsymbol{\Sigma}_{-j,j}\cdot Z_{j}$ and $\boldsymbol{\theta}=\boldsymbol{0}$.
-Figure 12 shows the standard deviation of $\widehat{\textnormal{FDR}}$ does not converge to 00 as the number of variables $d$ approach infinity.
-Figure 12: For variable selection on strongly positively equicorrelated $Z$-variables, standard deviation of $\widehat{\textnormal{FDR}}$ does not converge to 00 as the number of variables approach infinity.
+Figure 12 shows the standard deviation of $\widehat{\textnormal{FDR}}$ does not converge to 0 as the number of variables $d$ approach infinity.
+Figure 12: For variable selection on strongly positively equicorrelated $Z$-variables, standard deviation of $\widehat{\textnormal{FDR}}$ does not converge to 0 as the number of variables approach infinity.
 The noisy behavior of $\widehat{\textnormal{FDR}}$ is a consequence of the strong equicorrelation between the $Z$ variables.
@@
 The first-order optimal condition is
-$-\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}(\boldsymbol{Y}-\boldsymbol{X}\cdot\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}})+\boldsymbol{\nu}=\boldsymbol{0},$
+$\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}(\boldsymbol{Y}-\boldsymbol{X}\cdot\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}})+\boldsymbol{\nu}=\boldsymbol{0},$
 where $\boldsymbol{\nu}\in\mathbb{R}^{d}$ is the dual variable satisfying $\nu_{i}={\textnormal{sign}}(\widehat{\theta}^{\textnormal{Lasso}}_{i})\cdot\lambda$ for $\widehat{\theta}^{\textnormal{Lasso}}_{i}\neq 0$ and $\nu_{i}\in[-\lambda,\lambda]$ otherwise.
@@
 Similar ideas apply to the model-X settings.
-Appendix D Deferred proofs
+Appendix D Variance of FDR estimation
 In this section, we assume Assumption 3.1 and prove the propositions in Section 3.1.
 Throughout the section, we use $\mathcal{R}(\lambda)$ to denote the selection set of Lasso.
-Define $\mathcal{R}_{k}=\mathcal{R}\cap B_{k}$ as the selected variables in block $k$ and $R_{k}=|\mathcal{R}_{k}|$ as its cardinality.
-Throughout the section, we consider any pair of $(j,k)$ such that $j\in B_{k}$.
+Define $\mathcal{R}_{b}=\mathcal{R}\cap(b)$ as the selected variables in block $(b)$ and $R_{b}=|\mathcal{R}_{b}|$ as its cardinality.
+Throughout the section, we consider any pair of $(j,b)$ such that $j\in(b)$.
 We start with showing two technical lemmas.
 Lemma D.1.
-$\mathcal{R}_{k}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
-The p-value $p_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$, where $\textnormal{RSS}=\left\|\boldsymbol{Y}\right\|^{2}-\left\|\boldsymbol{X}(\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}\right\|^{2}$.
+$\mathcal{R}_{b}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
+The p-value $p_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$, where $\textnormal{RSS}=\left\|\boldsymbol{Y}\right\|^{2}-\left\|\boldsymbol{X}(\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}\right\|^{2}$.
 Proof.
-Let $\boldsymbol{Y}^{(k)}$ be the projection of $\boldsymbol{Y}$ onto the subspace spanned by the columns of $\boldsymbol{X}_{B_{k}}$, and denote the projection of $\boldsymbol{Y}$ onto the columns of $\boldsymbol{X}$ as $\widehat{\boldsymbol{Y}}=\boldsymbol{X}(\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
+Let $\boldsymbol{Y}^{(b)}$ be the projection of $\boldsymbol{Y}$ onto the subspace spanned by the columns of $\boldsymbol{X}_{(b)}$, and denote the projection of $\boldsymbol{Y}$ onto the columns of $\boldsymbol{X}$ as $\widehat{\boldsymbol{Y}}=\boldsymbol{X}(\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
 Then
-$\widehat{\boldsymbol{Y}}=\sum_{k=1}^{K}\boldsymbol{Y}^{(k)}$
-and the $\boldsymbol{Y}^{(k)}$s are orthogonal to each other under Assumption 3.1. We rewrite the Lasso coefficients
+$\widehat{\boldsymbol{Y}}=\sum_{b=1}^{B}\boldsymbol{Y}^{(b)}$
+and the $\boldsymbol{Y}^{(b)}$s are orthogonal to each other under Assumption 3.1. We rewrite the Lasso coefficients
 $\boldsymbol{\widehat{\theta}^{\textnormal{Lasso}}}(\lambda)$ $=\operatorname*{argmin}_{\boldsymbol{t}}\;\frac{1}{2}\,\left\|\boldsymbol{Y}-\boldsymbol{X}\boldsymbol{t}\right\|_{2}^{2}+\lambda\left\|\boldsymbol{t}\right\|_{1}$
 $=\operatorname*{argmin}_{\boldsymbol{t}}\;\frac{1}{2}\,\left\|\widehat{\boldsymbol{Y}}-\boldsymbol{X}\boldsymbol{t}\right\|_{2}^{2}+\lambda\left\|\boldsymbol{t}\right\|_{1}$
-$=\operatorname*{argmin}_{\boldsymbol{t}}\;\frac{1}{2}\,\left\|\sum_{k=1}^{K}\boldsymbol{Y}^{(k)}-\boldsymbol{X}_{B_{k}}\boldsymbol{t}_{B_{k}}\right\|_{2}^{2}+\lambda\sum_{k=1}^{K}\left\|\boldsymbol{t}_{B_{k}}\right\|_{1}$
-$=\operatorname*{argmin}_{\boldsymbol{t}}\;\sum_{k=1}^{K}\frac{1}{2}\,\left\|\boldsymbol{Y}^{(k)}-\boldsymbol{X}_{B_{k}}\boldsymbol{t}_{B_{k}}\right\|_{2}^{2}+\lambda\left\|\boldsymbol{t}_{B_{k}}\right\|_{1}.$
-We see solving for $\widehat{\theta}^{\textnormal{Lasso}}$ is equivalent to solving for Lasso coefficients for $K$ subproblems of regressing $\boldsymbol{Y}$ on $\boldsymbol{X}_{B_{k}}$.
-So $\widehat{\theta}^{\textnormal{Lasso}}_{B_{k}}(\lambda)$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}^{(k)}$, which is equal to $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
-Therefore, $\mathcal{R}_{k}=\left\{j\in B_{k}:\;\widehat{\theta}^{\textnormal{Lasso}}_{j}(\lambda)\neq 0\right\}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
+$=\operatorname*{argmin}_{\boldsymbol{t}}\;\frac{1}{2}\,\left\|\sum_{b=1}^{B}\boldsymbol{Y}^{(b)}-\boldsymbol{X}_{(b)}\boldsymbol{t}_{(b)}\right\|_{2}^{2}+\lambda\sum_{b=1}^{B}\left\|\boldsymbol{t}_{(b)}\right\|_{1}$
+$=\operatorname*{argmin}_{\boldsymbol{t}}\;\sum_{b=1}^{B}\frac{1}{2}\,\left\|\boldsymbol{Y}^{(b)}-\boldsymbol{X}_{(b)}\boldsymbol{t}_{(b)}\right\|_{2}^{2}+\lambda\left\|\boldsymbol{t}_{(b)}\right\|_{1}.$
+We see solving for $\widehat{\theta}^{\textnormal{Lasso}}$ is equivalent to solving for Lasso coefficients for $B$ subproblems of regressing $\boldsymbol{Y}$ on $\boldsymbol{X}_{(b)}$.
+So $\widehat{\theta}^{\textnormal{Lasso}}_{(b)}(\lambda)$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}^{(b)}$, which is equal to $\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
+Therefore, $\mathcal{R}_{b}=\left\{j\in(b):\;\widehat{\theta}^{\textnormal{Lasso}}_{j}(\lambda)\neq 0\right\}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
 Recall the p-value $p_{j}$ is a function of the t-statistic
@@
 We have
-$\left((\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\right)_{j}\cdot\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}=\left((\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{B_{k}})^{-1}\right)_{i(j,k)}\cdot\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},$
-where $i(j,k)$ is the index of variable $X_{j}$ in block $B_{k}$.
-Therefore $p_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$.
+$\left((\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X})^{-1}\right)_{j}\cdot\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}=\left((\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{(b)})^{-1}\right)_{i(j,b)}\cdot\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},$
+where $i(j,b)$ is the index of variable $X_{j}$ in block $(b)$.
+Therefore $p_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$.
 ∎
@@
 Proof of Proposition 3.1.
-Recall in the Gaussian linear model, we conditional on the explanatory variables $X$ and treat them as fixed.
-By Lemma D.1 and the block-orthogonality, $R_{k}$ are independent for $k=1,\ldots,K$.
-Our goal is to show $R=\sum_{k}R_{k}$ is stochastically larger than the sum of certain independent Bernoulli random variables.
+Recall in the Gaussian linear model, we condition on the explanatory variables $\boldsymbol{X}$ and treat them as fixed.
+By Lemma D.1 and the block-orthogonality, $R_{b}$ are independent for $b=1,\ldots,B$.
+Our goal is to show $R=\sum_{b}R_{b}$ is stochastically larger than the sum of certain independent Bernoulli random variables.
 Then the proposition follows by the concentration laws.
-Part 1.
-By Lemma D.1 and the first order condition, $R_{k}\geq 1$ if and only if there exist some $j\in B_{k}$ such that $|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda$.
-We have
-$\mathbb{P}(R_{k}\geq 1)=\mathbb{P}(\bigcup_{j\in B_{k}}\left\{|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda\right\})\geq\mathbb{P}(|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda)\geq 2(1-\Phi(\lambda/\sigma)).$
+Part 1. By Lemma D.1 and the first order condition, $R_{b}\geq 1$ if and only if there exist some $j\in(b)$ such that $|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda$. We have
+$\mathbb{P}(R_{b}\geq 1)=\mathbb{P}(\bigcup_{j\in(b)}\left\{|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda\right\})\geq\mathbb{P}(|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda)\geq 2(1-\Phi(\lambda/\sigma)).$
 The last inequality is because that $\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ is a Gaussian random variable with variance $\sigma^{2}$ (note $\boldsymbol{X}$ is normalized).
 Direct calculation shows the probability achieves minimum when $\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ has mean zero.
-As a consequence, $R=\sum_{k=1}^{K}R_{k}$ is stochastically larger than $\sum_{k=1}^{K}A_{k}$, where $A_{k}$ are $K$ independent Bernoulli random variables with success probability $\delta$.
+As a consequence, $R=\sum_{b=1}^{B}R_{b}$ is stochastically larger than $\sum_{b=1}^{B}A_{b}$, where $A_{b}$ are $B$ independent Bernoulli random variables with success probability $\delta$.
 By the multiplicative Chernoff bound, we have
-$\mathbb{P}(R\leq Kq)\;\leq\;\mathbb{P}\left(\sum_{k=1}^{K}A_{k}\leq\left(1-\frac{\delta-q}{\delta}\right)K\delta\right)\;\leq\;\exp\left(-\frac{(\delta-q)^{2}}{2\delta}\cdot K\right).$
+$\mathbb{P}(R\leq Bq)\;\leq\;\mathbb{P}\left(\sum_{b=1}^{B}A_{b}\leq\left(1-\frac{\delta-q}{\delta}\right)B\delta\right)\;\leq\;\exp\left(-\frac{(\delta-q)^{2}}{2\delta}\cdot B\right).$
 Part 2.
-To avoid triviality, assume $K\xi$ is an integer.
-Let $k$ denote any one of the $K\xi$ blocks that contains a variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$.
+To avoid triviality, assume $B\xi$ is an integer.
+Let $b$ denote any one of the $B\xi$ blocks that contains a variable with $|\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}|\geq\lambda$.
 Note $\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ is Gaussian with mean $\boldsymbol{X}_{j}\cdot\boldsymbol{\mu}$.
 By a similar reasoning as in Part 1,
-$\mathbb{P}(R_{k}\geq 1)\;\geq\;\mathbb{P}(|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda)\;\geq\;\frac{1}{2}.$
-As a consequence, $R=\sum_{k=1}^{K}R_{k}$ is stochastically larger than $\sum_{k=1}^{K\xi}A_{k}$, where $A_{k}$ are $K\xi$ independent Bernoulli random variables with success probability $1/2$.
+$\mathbb{P}(R_{b}\geq 1)\;\geq\;\mathbb{P}(|\boldsymbol{X}_{j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}|\geq\lambda)\;\geq\;\frac{1}{2}.$
+As a consequence, $R=\sum_{b=1}^{B}R_{b}$ is stochastically larger than $\sum_{b=1}^{B\xi}A_{b}$, where $A_{b}$ are $B\xi$ independent Bernoulli random variables with success probability $1/2$.
 By the multiplicative Chernoff bound, we have
-$\mathbb{P}(R\leq Kq)\;\leq\;\mathbb{P}\left(\sum_{k=1}^{K\xi}A_{k}\leq\left(1-\frac{\xi-2q}{\xi}\right)\frac{K\xi}{2}\right)\;\leq\;\exp\left(-\frac{(\xi-2q)^{2}}{4\xi}\cdot K\right).$
+$\mathbb{P}(R\leq Bq)\;\leq\;\mathbb{P}\left(\sum_{b=1}^{B\xi}A_{b}\leq\left(1-\frac{\xi-2q}{\xi}\right)\frac{B\xi}{2}\right)\;\leq\;\exp\left(-\frac{(\xi-2q)^{2}}{4\xi}\cdot B\right).$
 Proof of Theorem 3.1.
@@
 Hence in the proof, we assume $\sigma^{2}=1$ without loss of generality.
-Let $\boldsymbol{X}_{B_{k}}$ be the columns of $\boldsymbol{X}$ whose indices are in $B_{k}$ and $\boldsymbol{X}_{-B_{k}}$ be its complement.
-Let $\boldsymbol{X}_{B_{k}-j}$ be the same as $\boldsymbol{X}_{B_{k}}$ but with $\boldsymbol{X}_{j}$ excluded.
+Dependence on $\sigma^{2}$ can be restored by rescaling $\boldsymbol{Y}$ and $\lambda$.
+Let $\boldsymbol{X}_{(b)}$ be the columns of $\boldsymbol{X}$ whose indices are in $(b)$ and $\boldsymbol{X}_{-(b)}$ be its complement.
+Let $\boldsymbol{X}_{(b);-j}$ be the same matrix as $\boldsymbol{X}_{(b)}$ but with $\boldsymbol{X}_{j}$ excluded.
 Under Assumption 3.1,
-$\widehat{\textnormal{FDR}}^{*}_{j}$ $=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{|\mathcal{R}\cup\left\{j\right\}|}\mid\boldsymbol{S}_{j}\right]=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}_{k}\right\}}{|\mathcal{R}_{k}\cup\left\{j\right\}|+\sum_{l\neq k}R_{l}}\mid\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2}\right],$
+$\widehat{\textnormal{FDR}}^{*}_{j}$ $=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}\right\}}{|\mathcal{R}\cup\left\{j\right\}|}\mid\boldsymbol{S}_{j}\right]=\mathbb{E}_{H_{j}}\left[\frac{\mathbf{1}\left\{j\in\mathcal{R}_{b}\right\}}{|\mathcal{R}_{b}\cup\left\{j\right\}|+\sum_{l\neq b}R_{l}}\mid\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2}\right],$
 where $r_{j}^{2}=\textnormal{RSS}+u_{j}^{2}$ and $u_{j}$ is the norm of the projection of $\boldsymbol{Y}$ on to the direction of $(\boldsymbol{I}-\boldsymbol{X}_{-j}(\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{-j})^{-1}\boldsymbol{X}_{-j}^{\mkern-1.5mu\mathsf{T}})\boldsymbol{X}_{j}$.
-By Lemma D.1, $\sum_{l\neq k}R_{l}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ and $\mathcal{R}_{k}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
-The conditional expectation $\widehat{\textnormal{FDR}}^{*}_{j}$ is an one-dimensional integration over $u_{j}$.
+By Lemma D.1, $\sum_{l\neq b}R_{l}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{-(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ and $\mathcal{R}_{b}$ depends on $\boldsymbol{Y}$ only through $\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$.
+The conditional expectation $\widehat{\textnormal{FDR}}^{*}_{j}$ is a one-dimensional integration over $u_{j}$.
 To avoid confusion, we use $u_{j}$ to denote the value corresponding to the observed $\boldsymbol{Y}$ and $v_{j}$ to denote a placeholder for $u_{j}$ in the integral.
 We write
-$\widehat{\textnormal{FDR}}^{*}_{j}=\int_{\left\{v_{j}:\;j\in\mathcal{R}_{k}\right\}}\frac{1}{f(v_{j};\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})}\,\mu_{j}(dv_{j}\mid\boldsymbol{S}_{j}),$
+$\widehat{\textnormal{FDR}}^{*}_{j}=\int_{\left\{v_{j}:\;j\in\mathcal{R}_{b}\right\}}\frac{1}{f(v_{j};\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})}\,\mu_{j}(dv_{j}\mid\boldsymbol{S}_{j}),$
 where
-$f(v_{j};\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})=|\mathcal{R}_{k}\cup\left\{j\right\}|+\sum_{l\neq k}R_{l}$
+$f(v_{j};\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})=|\mathcal{R}_{b}\cup\left\{j\right\}|+\sum_{l\neq b}R_{l}$
 and
-$\mu_{j}(A\mid\boldsymbol{S}_{j})=\mathbb{P}_{H_{j}}(v_{j}\in A\mid\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})=\mathbb{P}_{H_{j}}(v_{j}\in A\mid\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2}).$
+$\mu_{j}(A\mid\boldsymbol{S}_{j})=\mathbb{P}_{H_{j}}(v_{j}\in A\mid\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\boldsymbol{X}_{-(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})=\mathbb{P}_{H_{j}}(v_{j}\in A\mid\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2}).$
 We suppress $\boldsymbol{S}_{j}$ and write $\mu_{j}(A)$ for simplicity of notation.
-Note the measure $\mu_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})$.
+Note the measure $\mu_{j}$ depends on $\boldsymbol{Y}$ only through $(\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})$.
 Note
-$f(v_{j})\in\left[1+\sum_{l\neq k}R_{l},\;b+\sum_{l\neq k}R_{l}\right]\subseteq\left[1\vee\left(-b+\sum_{k=1}^{K}R_{k}\right),\;b+\sum_{k=1}^{K}R_{k}\right].$
-By Lemma D.2, for some $\gamma_{j}\in[-b,b]$, we have
-$\widehat{\textnormal{FDR}}^{*}_{j}=\frac{\mu_{j}(j\in\mathcal{R}_{k})}{1\vee(R+\gamma_{j})}.$
+$f(v_{j})\in\left[1+\sum_{l\neq b}R_{l},\;m+\sum_{l\neq b}R_{l}\right]\subseteq\left[1\vee\left(-m+\sum_{b=1}^{B}R_{b}\right),\;m+\sum_{b=1}^{B}R_{b}\right].$
+By Lemma D.2, for some $\gamma_{j}\in[-m,m]$, we have
+$\widehat{\textnormal{FDR}}^{*}_{j}=\frac{\mu_{j}(j\in\mathcal{R}_{b})}{1\vee(R+\gamma_{j})}.$
 Therefore
-$\left|\widehat{\textnormal{FDR}}^{*}_{j}-\frac{\mu_{j}(j\in\mathcal{R}_{k})}{1\vee R}\right|=\left|\frac{\mu_{j}(j\in\mathcal{R}_{k})}{1\vee(R+\gamma_{j})}-\frac{\mu_{j}(j\in\mathcal{R}_{k})}{1\vee R}\right|\leq\frac{b}{1\vee R^{2}}\cdot\frac{1\vee R}{1\vee(R-b)}\leq\frac{4b^{2}}{1\vee R^{2}}.$
-For notational convenience, we suppress taking maximum with $1$ for $R$.
-Through out the section, we have $R$ represents $1\vee R$ and $\sum_{k}R_{k}$ represents $1\vee\sum_{k}R_{k}$.
-Then we can write
-$\widehat{\textnormal{FDR}}^{*}_{j}=\frac{\mu_{j}(j\in\mathcal{R}_{k})}{\sum_{k}R_{k}}+O\left(\frac{b^{2}}{R^{2}}\right).$
+$\left|\widehat{\textnormal{FDR}}^{*}_{j}-\frac{\mu_{j}(j\in\mathcal{R}_{b})}{1\vee R}\right|=\left|\frac{\mu_{j}(j\in\mathcal{R}_{b})}{1\vee(R+\gamma_{j})}-\frac{\mu_{j}(j\in\mathcal{R}_{b})}{1\vee R}\right|\leq\frac{m}{1\vee R^{2}}\cdot\frac{1\vee R}{1\vee(R-m)}\leq\frac{4m^{2}}{1\vee R^{2}}.$
+For notational convenience, we suppress taking maximum with $1$ for $R$. Throughout the section, we have $R$ represents $1\vee R$ and $\sum_{b}R_{b}$ represents $1\vee\sum_{b}R_{b}$. Then we can write
+$\widehat{\textnormal{FDR}}^{*}_{j}=\frac{\mu_{j}(j\in\mathcal{R}_{b})}{\sum_{b}R_{b}}+O\left(\frac{m^{2}}{R^{2}}\right).$
 Define
-$T_{k}=\sum_{j\in B_{k}}\mu_{j}(j\in\mathcal{R}_{k})\cdot\frac{\psi_{j}(p_{j})}{\mathbb{E}[\psi_{j}(p_{j})\mid\boldsymbol{S}_{j}]}.$
+$T_{b}=\sum_{j\in(b)}\mu_{j}(j\in\mathcal{R}_{b})\cdot\frac{\psi_{j}(p_{j})}{\mathbb{E}[\psi_{j}(p_{j})\mid\boldsymbol{S}_{j}]}.$
 We have
-$\sum_{j\in B_{k}}\widehat{\textnormal{FDR}}_{j}=\frac{T_{k}}{\sum_{k}R_{k}}+O\left(\frac{c_{\phi}b^{3}}{R^{2}}\right)$
+$\sum_{j\in(b)}\widehat{\textnormal{FDR}}_{j}=\frac{T_{b}}{\sum_{b}R_{b}}+O\left(\frac{c_{\phi}m^{3}}{R^{2}}\right)$
 and
-$\widehat{\textnormal{FDR}}=\sum_{k=1}^{K}\sum_{j\in B_{k}}\widehat{\textnormal{FDR}}_{j}=\frac{\sum_{k}T_{k}}{\sum_{k}R_{k}}+O\left(\frac{Kc_{\phi}b^{3}}{R^{2}}\right)$
-Note we can compute $(\boldsymbol{X}_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})$ from $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$ for any $j\in B_{k}$.
-Moreover, by Lemma D.1, for any $j\in B_{k}$, $p_{j}$ depends on $Y$ only through $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$.
-So $T_{k}$ depends on $Y$ only through $(\boldsymbol{X}_{B_{k}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$.
-Therefore, condition on RSS, both $\left\{T_{k}:\;k\in[K]\right\}$ and $\left\{R_{k}:\;k\in[K]\right\}$ are sets of conditional independent random variables respectively.
-We denote $T=\sum_{k}T_{k}$ for simplicity.
+$\widehat{\textnormal{FDR}}=\sum_{b=1}^{B}\sum_{j\in(b)}\widehat{\textnormal{FDR}}_{j}=\frac{\sum_{b}T_{b}}{\sum_{b}R_{b}}+O\left(\frac{Bc_{\phi}m^{3}}{R^{2}}\right)$
+Note we can compute $(\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})$ from $(\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$ for any $j\in(b)$.
+Moreover, by Lemma D.1, for any $j\in(b)$, $p_{j}$ depends on $Y$ only through $(\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$.
+So $T_{b}$ depends on $Y$ only through $(\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},\textnormal{RSS})$.
+Therefore, condition on RSS, both $\left\{T_{b}:\;b\in[B]\right\}$ and $\left\{R_{b}:\;b\in[B]\right\}$ are sets of conditional independent random variables respectively.
+We denote $T=\sum_{b}T_{b}$ for simplicity.
 Next we prove the theorem in three steps.
 Step 1.
-$\textnormal{Var}\left({\sum_{k}T_{k}}/{\sum_{k}R_{k}}\mid\textnormal{RSS}\right)\leq C(c_{\phi}^{2}b^{4}/Kq^{4}+c_{\phi}^{2}b^{2}K^{2}\rho)$ almost surely for some $C>0$ that does not depend on RSS.
-[…]
+$\textnormal{Var}\left({\sum_{b}T_{b}}/{\sum_{b}R_{b}}\mid\textnormal{RSS}\right)\leq C(c_{\phi}^{2}m^{4}/Bq^{4}+c_{\phi}^{2}m^{2}B^{2}\rho)$ almost surely for some $C>0$ that does not depend on RSS.
+For simplicity of notation, we denote $\textnormal{Var}(\cdot|\cdot,\textnormal{RSS})$ as $\textnormal{Var}_{\textnormal{RSS}}(\cdot|\cdot)$ and $\mathbb{E}(\cdot|\cdot,\textnormal{RSS})$ as $\mathbb{E}_{\textnormal{RSS}}(\cdot|\cdot)$.
+Let $q$ be the number in Proposition 3.1.
+We have
 $\textnormal{Var}_{\textnormal{RSS}}\left[\frac{T}{R}\right]=$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}$
-$=$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R,R^{\prime}>Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Kq)$
-$+\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R\wedge R^{\prime}\leq Kq),$
+$=$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R,R^{\prime}>Bq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Bq)$
+$+\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Bq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R\wedge R^{\prime}\leq Bq),$
 where $(T^{\prime},R^{\prime})$ is an IID copy of $(T,R)$ conditional on RSS. We will control each term separately. First,
-$\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R,R^{\prime}>Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Kq)$
-$=$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{TR^{\prime}-T^{\prime}R}{RR^{\prime}}\right)^{2}\mid R,R^{\prime}>Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Kq)$
-$\leq$ $\frac{1}{2K^{4}q^{4}}\cdot\mathbb{E}_{\textnormal{RSS}}\left[\left(TR^{\prime}-T^{\prime}R\right)^{2}\mid R,R^{\prime}>Kq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Kq)$
-$\leq$ $\frac{1}{2K^{4}q^{4}}\cdot\mathbb{E}_{\textnormal{RSS}}\left[\left(TR^{\prime}-T^{\prime}R\right)^{2}\right]$
-$=$ $\frac{1}{2K^{4}q^{4}}\cdot\left(2\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]-2\left(\mathbb{E}_{\textnormal{RSS}}^{2}\left[TR\right]-\mathbb{E}_{\textnormal{RSS}}^{2}\left[T\right]\mathbb{E}_{\textnormal{RSS}}^{2}\left[R\right]\right)\right)$
-$\leq$ $\frac{1}{K^{4}q^{4}}\cdot\left(\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]+2\mathbb{E}_{\textnormal{RSS}}\left[T\right]\mathbb{E}_{\textnormal{RSS}}\left[R\right]\cdot|\text{Cov}_{\textnormal{RSS}}(T,R)|\right)$
+$\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R,R^{\prime}>Bq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Bq)$
+$=$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{TR^{\prime}-T^{\prime}R}{RR^{\prime}}\right)^{2}\mid R,R^{\prime}>Bq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Bq)$
+$\leq$ $\frac{1}{2B^{4}q^{4}}\cdot\mathbb{E}_{\textnormal{RSS}}\left[\left(TR^{\prime}-T^{\prime}R\right)^{2}\mid R,R^{\prime}>Bq\right]\cdot\mathbb{P}_{\textnormal{RSS}}(R,R^{\prime}>Bq)$
+$\leq$ $\frac{1}{2B^{4}q^{4}}\cdot\mathbb{E}_{\textnormal{RSS}}\left[\left(TR^{\prime}-T^{\prime}R\right)^{2}\right]$
+$=$ $\frac{1}{2B^{4}q^{4}}\cdot\left(2\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]-2\left(\mathbb{E}_{\textnormal{RSS}}^{2}\left[TR\right]-\mathbb{E}_{\textnormal{RSS}}^{2}\left[T\right]\mathbb{E}_{\textnormal{RSS}}^{2}\left[R\right]\right)\right)$
+$\leq$ $\frac{1}{B^{4}q^{4}}\cdot\left(\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]+2\mathbb{E}_{\textnormal{RSS}}\left[T\right]\mathbb{E}_{\textnormal{RSS}}\left[R\right]\cdot|\text{Cov}_{\textnormal{RSS}}(T,R)|\right)$
 The last inequality is obtained by considering cases where $\text{Cov}_{\textnormal{RSS}}(T,R)$ is positive or negative separately.
@@
 $\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]=\textnormal{Var}_{\textnormal{RSS}}(T)\textnormal{Var}_{\textnormal{RSS}}(R^{\prime})+\textnormal{Var}_{\textnormal{RSS}}(T)(\mathbb{E}_{\textnormal{RSS}}R^{\prime})^{2}+\textnormal{Var}_{\textnormal{RSS}}(R^{\prime})(\mathbb{E}_{\textnormal{RSS}}T)^{2}$
-Note $T_{k}\in[0,c_{\phi}b]$, we have $\mathbb{E}_{\textnormal{RSS}}[T]\leq Kc_{\phi}b$ and $\textnormal{Var}_{\textnormal{RSS}}[T]=\sum_{k}\textnormal{Var}_{\textnormal{RSS}}[T_{k}]\leq Kc_{\phi}^{2}b^{2}$.
-Similarly, $\mathbb{E}_{\textnormal{RSS}}[R^{\prime}]\leq Kb$ and $\textnormal{Var}_{\textnormal{RSS}}[R^{\prime}]\leq Kb^{2}$.
+Note $T_{b}\in[0,c_{\phi}m]$, we have $\mathbb{E}_{\textnormal{RSS}}[T]\leq Bc_{\phi}m$ and $\textnormal{Var}_{\textnormal{RSS}}[T]=\sum_{b}\textnormal{Var}_{\textnormal{RSS}}[T_{b}]\leq Bc_{\phi}^{2}m^{2}$.
+Similarly, $\mathbb{E}_{\textnormal{RSS}}[R^{\prime}]\leq Bm$ and $\textnormal{Var}_{\textnormal{RSS}}[R^{\prime}]\leq Bm^{2}$.
 Thus
-$\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]\leq K^{2}c_{\phi}^{2}b^{4}+K^{3}c_{\phi}^{2}b^{4}=O(b^{4}K^{3})$
+$\textnormal{Var}_{\textnormal{RSS}}\left[TR^{\prime}\right]\leq B^{2}c_{\phi}^{2}m^{4}+B^{3}c_{\phi}^{2}m^{4}=O(m^{4}B^{3})$
 and
-$|\text{Cov}_{\textnormal{RSS}}(T,R)|\leq\sqrt{\textnormal{Var}_{\textnormal{RSS}}[T]\textnormal{Var}_{\textnormal{RSS}}[R]}\leq Kb^{2}c_{\phi}.$
-Therefore the first term is bounded by $O(b^{4}/K)$. For the second term,
-$\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Kq\right]$
-$\leq$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}+\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Kq\right]$
-$\leq$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(T+T^{\prime}\right)^{2}\mid R\wedge R^{\prime}\leq Kq\right]$
-$\leq$ $2K^{2}c_{\phi}^{2}b^{2}.$
-The last inequality is because $T\leq Kc_{\phi}b$. By Proposition 3.1, we have
-$\mathbb{P}_{\textnormal{RSS}}(R\wedge R^{\prime}\leq Kq)\leq 2\mathbb{P}_{\textnormal{RSS}}(R\leq Kq)\leq 2\rho.$
+$|\text{Cov}_{\textnormal{RSS}}(T,R)|\leq\sqrt{\textnormal{Var}_{\textnormal{RSS}}[T]\textnormal{Var}_{\textnormal{RSS}}[R]}\leq Bm^{2}c_{\phi}.$
+Therefore the first term is bounded by $O(m^{4}/B)$. For the second term,
+$\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}-\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Bq\right]$
+$\leq$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(\frac{T}{R}+\frac{T^{\prime}}{R^{\prime}}\right)^{2}\mid R\wedge R^{\prime}\leq Bq\right]$
+$\leq$ $\frac{1}{2}\,\mathbb{E}_{\textnormal{RSS}}\left[\left(T+T^{\prime}\right)^{2}\mid R\wedge R^{\prime}\leq Bq\right]$
+$\leq$ $2B^{2}c_{\phi}^{2}m^{2}.$
+The last inequality is because $T\leq Bc_{\phi}m$. By Proposition 3.1, we have
+$\mathbb{P}_{\textnormal{RSS}}(R\wedge R^{\prime}\leq Bq)\leq 2\mathbb{P}_{\textnormal{RSS}}(R\leq Bq)\leq 2\rho.$
 Combine the results above, we have
-$\textnormal{Var}_{\textnormal{RSS}}\left[\frac{T}{R}\right]\leq c_{\phi}^{2}\cdot\left[\frac{K^{2}b^{4}+5K^{3}b^{4}}{K^{4}q^{4}}+2K^{2}b^{2}\cdot 2\rho\right].$
-The bound is of $O\left({b^{4}}/{K}\right)$ if $\rho=O(K^{-3})$.
-Step 2. $\textnormal{Var}\left[\mathbb{E}\left({\sum_{k}T_{k}}/{\sum_{k}R_{k}}\mid\textnormal{RSS}\right)\right]=O(1/(n-d))$.
-We start by bounding the perturbation of $\mu_{j}(j\in\mathcal{R}_{k})$ when we perturb RSS while holding $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ fixed. We write
-[…]
-where $r_{j}=\sqrt{\textnormal{RSS}+u_{j}^{2}}$, $I_{j}(v_{j};X_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})=\mathbf{1}\left\{j\in\mathcal{R}_{k}\right\}$, and $q_{j}=d\mu_{j}/dv_{j}$ is the probability density.
+$\textnormal{Var}_{\textnormal{RSS}}\left[\frac{T}{R}\right]\leq c_{\phi}^{2}\cdot\left[\frac{B^{2}m^{4}+5B^{3}m^{4}}{B^{4}q^{4}}+2B^{2}m^{2}\cdot 2\rho\right].$
+The bound is of $O\left({m^{4}}/{B}\right)$ if $\rho=O(B^{-3})$.
+Step 2. $\textnormal{Var}\left[\mathbb{E}\left({\sum_{b}T_{b}}/{\sum_{b}R_{b}}\mid\textnormal{RSS}\right)\right]=O(1/(n-d))$.
+We start by bounding the perturbation of $\mu_{j}(j\in\mathcal{R}_{b})$ when we perturb RSS while holding $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$ fixed. We write
+$\mu_{j}(j\in\mathcal{R}_{b})=\mathbb{P}_{H_{j}}(v_{j}\in A\mid\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y},r_{j}^{2})=\int_{-r_{j}}^{r_{j}}I_{j}(v_{j};\boldsymbol{X}_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})\cdot q_{j}(v_{j})\,dv_{j},$
+where $r_{j}=\sqrt{\textnormal{RSS}+u_{j}^{2}}$, $I_{j}(v_{j};X_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})=\mathbf{1}\left\{j\in\mathcal{R}_{b}\right\}$, and $q_{j}=d\mu_{j}/dv_{j}$ is the probability density.
 Direct calculation shows
@@
 Note that $q_{j}(r_{j})=q_{j}(-r_{j})=0$. By the bounded convergence theorem,
-$\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{k})=\int_{-r_{j}}^{r_{j}}I_{j}(v_{j};X_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})\cdot\partial_{\textnormal{RSS}}\,q_{j}(v_{j})\,dv_{j},$
+$\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{b})=\int_{-r_{j}}^{r_{j}}I_{j}(v_{j};X_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})\cdot\partial_{\textnormal{RSS}}\,q_{j}(v_{j})\,dv_{j},$
 where
@@
 Note that $\partial_{\textnormal{RSS}}\,q_{j}\geq 0$ if and only if $v_{j}^{2}\geq\frac{r_{j}^{2}}{n-d-1}$ and
-[…]
-Since $I_{j}(v_{j};X_{B_{k}-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})\in\left\{0,1\right\}$, we have
-$|\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{k})|$ $\leq\max\left\{\int_{r_{j}^{2}\geq v_{j}^{2}\geq\frac{r_{j}^{2}}{n-d-1}}\partial_{\textnormal{RSS}}\,q_{j}\,dv_{j},\;-\int_{v_{j}^{2}\leq\frac{r_{j}^{2}}{n-d-1}}\partial_{\textnormal{RSS}}\,q_{j}\,dv_{j}\right\}$
+$\int_{-r_{j}}^{r_{j}}\partial_{\textnormal{RSS}}q_{j}dv_{j}=\partial_{\textnormal{RSS}}\int_{-r_{j}}^{r_{j}}q_{j}dv_{j}=\partial_{\textnormal{RSS}}1=0.$
+Since $I_{j}(v_{j};X_{(b);-j}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y})\in\left\{0,1\right\}$, we have
+$|\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{b})|$ $\leq\max\left\{\int_{r_{j}^{2}\geq v_{j}^{2}\geq\frac{r_{j}^{2}}{n-d-1}}\partial_{\textnormal{RSS}}\,q_{j}\,dv_{j},\;-\int_{v_{j}^{2}\leq\frac{r_{j}^{2}}{n-d-1}}\partial_{\textnormal{RSS}}\,q_{j}\,dv_{j}\right\}$
 $=\int_{r_{j}^{2}\geq v_{j}^{2}\geq\frac{r_{j}^{2}}{n-d-1}}\partial_{\textnormal{RSS}}\,q_{j}\,dv_{j}$
@@
 $p_{j}=2\cdot F_{t_{n-d}}\left(-\frac{|v_{j}|}{\sqrt{\textnormal{RSS}/(n-d)}}\right),$
-where $F_{t_{n-d}}$ is the cumulative distribution function of the t-distribution of degree of freedom $n-d$. Then
+where $F_{t_{n-d}}$ is the cumulative distribution function of the t-distribution of degree of freedom ndn-d. Then
 $\partial_{\textnormal{RSS}}\,p_{j}=c_{n-d}\cdot\left(1+\frac{v_{j}^{2}}{\textnormal{RSS}}\right)^{-\frac{n-d+1}{2}}\cdot\frac{|v_{j}|}{\sqrt{\textnormal{RSS}/(n-d)}}\cdot\frac{1}{\textnormal{RSS}}.$
@@
 Recall we assume $\sigma^{2}=1$.
-Since $\textnormal{RSS}\sim\chi^{2}_{n-d}$ is sub-exponential, we have $\mathbb{P}(A^{\mathsf{c}})\leq 2\exp(-(n-d)/32)$ by the Bernstein concentration equalities.
+Since $\textnormal{RSS}\sim\chi^{2}_{n-d}$ is sub-exponential, we have $\mathbb{P}(A^{\mathsf{c}})\leq 2\exp(-(n-d)/32)$ by the Bernstein concentration inequalities.
 On $A_{\textnormal{RSS}}$, we have
-$|\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{k})|\leq\frac{2}{n-d}$
+$|\partial_{\textnormal{RSS}}\,\mu_{j}(j\in\mathcal{R}_{b})|\leq\frac{2}{n-d}$
 and
@@
 $\leq\frac{1}{n-d}.$
-Note both $\mu_{j}(j\in\mathcal{R}_{k})$ and $\phi_{j}(p_{j})$ are bounded. Therefore, $\mu_{j}(j\in\mathcal{R}_{k})\cdot\phi_{j}(p_{j})$ is Lipschitz in RSS with a Lipschitz constant
-$\left(2c_{\phi}+L_{\psi}\right)\cdot\frac{1}{n-d}.$
-Moreover, $\sum_{k}T_{k}=\sum_{j=1}^{Kb}\mu_{j}(j\in\mathcal{R}_{k})\cdot\phi_{j}(p_{j})$ is then Lipschitz in RSS with a Lipschitz constant $Kb\left(2c_{\phi}+L_{\psi}\right)/(n-d)$.
-A key observation is that ${\sum_{k}T_{k}}/{R}$ is Lipschitz in RSS since $R$ is not a function of RSS.
-Now consider the event $A_{R}=\left\{R>Kq\right\}$.
+Note both $\mu_{j}(j\in\mathcal{R}_{b})$ and $\phi_{j}(p_{j})$ are bounded. Therefore, $\mu_{j}(j\in\mathcal{R}_{b})\cdot\phi_{j}(p_{j})$ is Lipschitz in RSS with a Lipschitz constant
+$\left(2c_{\phi}+L_{\phi}\right)\cdot\frac{1}{n-d}.$
+Moreover, $\sum_{b}T_{b}=\sum_{j=1}^{d}\mu_{j}(j\in\mathcal{R}_{b})\cdot\phi_{j}(p_{j})$ is then Lipschitz in RSS with a Lipschitz constant $Bm\left(2c_{\phi}+L_{\phi}\right)/(n-d)$.
+A key observation is that ${\sum_{b}T_{b}}/{R}$ is Lipschitz in RSS since $R$ is not a function of RSS.
+Now consider the event $A_{R}=\left\{R>Bq\right\}$.
 Since RSS is independent of $\boldsymbol{X}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}$, $A_{\textnormal{RSS}}$ and $A_{R}$ are independent.
 In addition, we have $\mathbb{P}(A_{R}^{\mathsf{c}})\leq\rho$.
-On $A_{\textnormal{RSS}}\cap A_{R}$, we have ${\sum_{k}T_{k}}/{R}$ is $L$-Lipschitz in RSS with
-$L=\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)\cdot\frac{1}{n-d}.$
+On $A_{\textnormal{RSS}}\cap A_{R}$, we have ${\sum_{b}T_{b}}/{R}$ is $L$-Lipschitz in RSS with
+$L=\left(\frac{2mc_{\phi}+mL_{\phi}}{q}\right)\cdot\frac{1}{n-d}.$
 Therefore, for any $\textnormal{RSS}\in A_{\textnormal{RSS}}$,
-$\mathbb{E}\left[\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS},A_{R}\right]$
+$\mathbb{E}\left[\frac{\sum_{b}T_{b}}{R}\mid\textnormal{RSS},A_{R}\right]$
 is $L$-Lipschitz in RSS.
 For notational convenience, we denote
-$E_{R}(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS},A_{R}\right],\quad E_{R^{\mathsf{c}}}(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS},A_{R}^{\mathsf{c}}\right],$
+$E_{R}(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{b}T_{b}}{R}\mid\textnormal{RSS},A_{R}\right],\quad E_{R^{\mathsf{c}}}(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{b}T_{b}}{R}\mid\textnormal{RSS},A_{R}^{\mathsf{c}}\right],$
 and
-$E(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS}\right]=E_{R}(\textnormal{RSS})\cdot\mathbb{P}(A_{R})+E_{R^{\mathsf{c}}}(\textnormal{RSS})\cdot\mathbb{P}(A_{R}^{\mathsf{c}}).$
-Note all $E_{R}(\textnormal{RSS})$, $E_{R^{\mathsf{c}}}(\textnormal{RSS})$, and $E(\textnormal{RSS})$ are bounded by $Kc_{\phi}b$.
+$E(\textnormal{RSS})=\mathbb{E}\left[\frac{\sum_{b}T_{b}}{R}\mid\textnormal{RSS}\right]=E_{R}(\textnormal{RSS})\cdot\mathbb{P}(A_{R})+E_{R^{\mathsf{c}}}(\textnormal{RSS})\cdot\mathbb{P}(A_{R}^{\mathsf{c}}).$
+Note all $E_{R}(\textnormal{RSS})$, $E_{R^{\mathsf{c}}}(\textnormal{RSS})$, and $E(\textnormal{RSS})$ are bounded by $Bc_{\phi}m$.
 Let $\textnormal{RSS}^{\prime}$ be an independent copy of RSS. Then
@@
 $\quad 2(E_{R}(\textnormal{RSS})-E_{R}(\textnormal{RSS}^{\prime}))(E_{R^{\mathsf{c}}}(\textnormal{RSS})-E_{R^{\mathsf{c}}}(\textnormal{RSS}^{\prime}))\cdot\mathbb{P}(A_{R})\mathbb{P}(A_{R}^{\mathsf{c}})$
-$=$ $(E_{R}(\textnormal{RSS})-E_{R}(\textnormal{RSS}^{\prime}))^{2}+O(c_{\phi}^{2}b^{2}\cdot K^{2}\rho)$
+$=$ $(E_{R}(\textnormal{RSS})-E_{R}(\textnormal{RSS}^{\prime}))^{2}+O(c_{\phi}^{2}m^{2}\cdot B^{2}\rho)$
 Therefore,
@@
 $=$ $\frac{1}{2}\mathbb{E}\left[(E_{R}(\textnormal{RSS})-E_{R}(\textnormal{RSS}^{\prime}))^{2}\mid A_{\textnormal{RSS}}\right]\cdot\mathbb{P}(A_{\textnormal{RSS}})$
-$\quad+O(c_{\phi}^{2}b^{2}\cdot K^{2}\rho)+O\left(c_{\phi}^{2}b^{2}\cdot K^{2}e^{-{(n-d)}/{32}}\right)$
+$\quad+O(c_{\phi}^{2}m^{2}\cdot B^{2}\rho)+O\left(c_{\phi}^{2}m^{2}\cdot B^{2}e^{-{(n-d)}/{32}}\right)$
 Since $E_{R}(\textnormal{RSS})$ is $L$-Lipschitz in RSS,
@@
 $\leq$ $L^{2}\textnormal{Var}(\textnormal{RSS})$
-$=$ $\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{2}{n-d},$
-where the last line uses the fact that $(n-d)\textnormal{RSS}\sim\chi^{2}(n-d)$ and hence $\textnormal{Var}(\textnormal{RSS})=2/(n-d)$. Then we have
-$\textnormal{Var}\left[E(\textnormal{RSS})\right]\leq\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}b^{2}\cdot K^{2}\rho)+O\left(c_{\phi}^{2}b^{2}\cdot K^{2}e^{-(n-d)/32}\right).$
+$=$ $\left(\frac{2mc_{\phi}+mL_{\phi}}{q}\right)^{2}\cdot\frac{2}{n-d},$
+where the last line uses the fact that $\textnormal{RSS}\sim\chi^{2}(n-d)$ and hence $\textnormal{Var}(\textnormal{RSS})=2(n-d)$. Then we have
+$\textnormal{Var}\left[E(\textnormal{RSS})\right]\leq\left(\frac{2mc_{\phi}+mL_{\phi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}m^{2}\cdot B^{2}\rho)+O\left(c_{\phi}^{2}m^{2}\cdot B^{2}e^{-(n-d)/32}\right).$
 That is
-$\textnormal{Var}\left[\mathbb{E}\left(\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS}\right)\right]\leq\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}b^{2}\cdot K^{2}(\rho\vee e^{-(n-d)/32})).$
+$\textnormal{Var}\left[\mathbb{E}\left(\frac{\sum_{b}T_{b}}{R}\mid\textnormal{RSS}\right)\right]\leq\left(\frac{2mc_{\phi}+mL_{\phi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}m^{2}\cdot B^{2}(\rho\vee e^{-(n-d)/32})).$
 Combining Step 1 and Step 2, we have
-$\textnormal{Var}\left[\frac{\sum_{k}T_{k}}{R}\right]$ $=\mathbb{E}\left[\textnormal{Var}_{\textnormal{RSS}}\left(\frac{T}{R}\right)\right]+\textnormal{Var}\left[\mathbb{E}\left(\frac{\sum_{k}T_{k}}{R}\mid\textnormal{RSS}\right)\right]$
-$\leq O\left(\frac{c_{\phi}^{2}b^{4}}{Kq^{4}}\right)+\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}b^{2}\cdot K^{2}(\rho\vee e^{-(n-d)/32}))$
+$\textnormal{Var}\left[\frac{\sum_{b}T_{b}}{R}\right]$ $=\mathbb{E}\left[\textnormal{Var}_{\textnormal{RSS}}\left(\frac{T}{R}\right)\right]+\textnormal{Var}\left[\mathbb{E}\left(\frac{\sum_{b}T_{b}}{R}\mid\textnormal{RSS}\right)\right]$
+$\leq O\left(\frac{c_{\phi}^{2}m^{4}}{Bq^{4}}\right)+\left(\frac{2mc_{\phi}+mL_{\phi}}{q}\right)^{2}\cdot\frac{2}{n-d}+O(c_{\phi}^{2}m^{2}\cdot B^{2}(\rho\vee e^{-(n-d)/32}))$
 Finally,
-$\textnormal{Var}\left[\widehat{\textnormal{FDR}}\right]=$ $\textnormal{Var}\left[\frac{\sum_{k}T_{k}}{\sum_{k}R_{k}}+O\left(\frac{Kc_{\phi}b^{3}}{R^{2}}\right)\right]$
-$\leq$ $2\textnormal{Var}\left[\frac{\sum_{k}T_{k}}{\sum_{k}R_{k}}\right]+O\left(2\mathbb{E}\left[\frac{K^{2}c_{\phi}^{2}b^{6}}{R^{4}}\right]\right)$
-$\leq$ $2\textnormal{Var}\left[\frac{\sum_{k}T_{k}}{\sum_{k}R_{k}}\right]+O\left(\frac{c_{\phi}^{2}b^{6}}{K^{2}q^{4}}\right)+O(K^{2}c_{\phi}^{2}b^{4}\rho)$
-$\leq$ $O\left(c_{\phi}^{2}\left(\frac{b^{4}}{q^{4}}+\frac{b^{6}}{Kq^{4}}\right)\cdot\frac{1}{K}\right)+\left(\frac{2bc_{\phi}+bL_{\psi}}{q}\right)^{2}\cdot\frac{4}{n-d}$
-$+O(c_{\phi}^{2}b^{4}\cdot K^{2}(\rho\vee e^{-(n-d)/32})).$
+$\textnormal{Var}\left[\widehat{\textnormal{FDR}}\right$]= $\textnormal{Var}\left[\frac{\sum_{b}T_{b}}{\sum_{b}R_{b}}+O\left(\frac{Bc_{\phi}m^{3}}{R^{2}}\right)\right]$
+$\leq$ $2\textnormal{Var}\left[\frac{\sum_{b}T_{b}}{\sum_{b}R_{b}}\right]+O\left(2\mathbb{E}\left[\frac{B^{2}c_{\phi}^{2}m^{6}}{R^{4}}\right]\right)$
+$\leq$ $2\textnormal{Var}\left[\frac{\sum_{b}T_{b}}{\sum_{b}R_{b}}\right]+O\left(\frac{c_{\phi}^{2}m^{6}}{B^{2}q^{4}}\right)+O(B^{2}c_{\phi}^{2}m^{6}\rho)$
+$\leq$ $O\left(c_{\phi}^{2}\left(\frac{m^{4}}{q^{4}}+\frac{m^{6}}{Bq^{4}}\right)\cdot\frac{1}{B}\right)+\left(\frac{2mc_{\phi}+mL_{\phi}}{q}\right)^{2}\cdot\frac{4}{n-d}$
+$+O(c_{\phi}^{2}m^{6}\cdot B^{2}(\rho\vee e^{-(n-d)/32})).$
 The big-O notation only hides universal constant.
 Proof of Corollary 3.1.
-Result (1): The setting satisfies the conditions in Part 2 of Proposition 3.1 with $\xi=K^{-\varepsilon}$ and $q=\xi/3=O(K^{-\varepsilon})$.
-Then $\rho=\exp\left(-K^{1-\varepsilon}/36\right)$.
+Result (1): The setting satisfies the conditions in Part 2 of Proposition 3.1 with $\xi=B^{-\varepsilon}$ and $q=\xi/3=O(B^{-\varepsilon})$.
+Then $\rho=\exp\left(-B^{1-\varepsilon}/36\right)$.
 Plugging them into the bound in Theorem 3.1, we have
-$\textnormal{Var}(\widehat{\textnormal{FDR}})=O\left(K^{-1+4\varepsilon}+\frac{K^{2\varepsilon}}{n-d}+K^{2}\exp\left(-K^{1-\varepsilon}/36\right)\right)=O(K^{-1+4\varepsilon}).$
-Result (2): The setting satisfies the conditions in Part 2 of Proposition 3.1 with $q=\xi/3$, and $\rho=\exp\left(-K\xi/36\right)$. Plugging them into the bound in Theorem 3.1, we have
-$\textnormal{Var}(\widehat{\textnormal{FDR}})=O\left(K^{-1}+\frac{1}{n-d}+K^{2}\exp\left(-(n-d)/32\right)\right)=O(K^{-1}).$
-Appendix E Extended numerical results
+$\textnormal{Var}(\widehat{\textnormal{FDR}})=O\left(B^{-1+4\varepsilon}+\frac{B^{2\varepsilon}}{n-d}+B^{2}\exp\left(-B^{1-\varepsilon}/36\right)\right)=O(B^{-1+4\varepsilon}).$
+Result (2): The setting satisfies the conditions in Part 2 of Proposition 3.1 with $q=\xi/3$, and $\rho=\exp\left(-B\xi/36\right)$. Plugging them into the bound in Theorem 3.1, we have
+$\textnormal{Var}(\widehat{\textnormal{FDR}})=O\left(B^{-1}+\frac{1}{n-d}+B^{2}\exp\left(-(n-d)/32\right)\right)=O(B^{-1}).$
+Appendix E Consistency of Bootstrap variance estimation
+E.1 Setup and notation
+We consider the Gaussian linear model with block-orthogonal design as in Assumption 3.1.
+In particular, the $d$ columns of $\boldsymbol{X}$ are partitioned into $B$ blocks, specified by the index sets $(b)\subseteq[d]$ for $b=1,\ldots,B$.
+The columns in different blocks are orthogonal.
+Moreover, we assume that $\sigma^{2}$ is known throughout the section.
+We consider an asymptotic regime in which the number of blocks $B$ tends to infinity. Hence so does the number of variables:
+$d\;=\;\sum_{b=1}^{B}m_{b}\;\to\;\infty.$
+Formally, consider a sequence of such models indexed by $d$, with design matrices $\boldsymbol{X}=\boldsymbol{X}_{d}\in\mathbb{R}^{n_{d}\times d}$ satisfying $n_{d}>d$ for all $d$ and $n_{d}\to\infty$ as $d\to\infty$, and we suppress the index $d$ from the notation when there is no ambiguity.
+For each block $b\in[B]:=\{1,\ldots,B\}$, let $\boldsymbol{X}_{(b)}\in\mathbb{R}^{n\times m_{b}}$ denote the submatrix of $\boldsymbol{X}$ formed by the columns in block $b$, let $\boldsymbol{\theta}_{(b)}\in\mathbb{R}^{m_{b}}$ be the corresponding coefficient vector, and let
+$\Sigma_{(b)}=\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{(b)}$
+be the within-block Gram matrix. We write the sufficient statistics within block $b$ as
+$\boldsymbol{Z}_{(b)}=\boldsymbol{X}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Y}\;\sim\;\mathcal{N}\big(\Sigma_{(b)}\boldsymbol{\theta}_{(b)},\;\sigma^{2}\Sigma_{(b)}\big).$
+In this section, the notation becomes simpler if we index variables within a given block.
+Thus, we make a slight notational adjustment from the main text, where $j$ denotes a global index across all variables.
+Here, we fix a block $b$ and let $j\in[m_{b}]$ denote the within-block index.
+For example, $\boldsymbol{Z}_{(b);j}$ refers to the $j$th coordinate of the vector $\boldsymbol{Z}_{(b)}$.
+Similarly, we write $\Sigma_{(b);-j,j}\in\mathbb{R}^{m_{b}-1}$ for the $j$th column of $\Sigma_{(b)}$ with its $j$th entry removed, and $\Sigma_{(b);j,-j}=\Sigma_{(b);-j,j}^{\top}$ for the corresponding row.
+We also use
+$\Sigma_{(b);-j,-j}\in\mathbb{R}^{(m_{b}-1)\times(m_{b}-1)}$
+to denote the principal submatrix obtained by removing the $j$th row and $j$th column of $\Sigma_{(b)}$.
+We make the following standing assumptions. In particular, the Assumption A1 is the same as Assumption 3.1 in the main text. We state it here to make the discussion self-contained.
+  1. A1)
+    (Column normalization and block orthogonality) $\|\boldsymbol{X}_{(b),j}\|=1$, and hence $\Sigma_{(b);jj}=1$, for all $b\in[B],j\in[m_{b}]$ for any $b_{1}\neq b_{2}$,
+    $\boldsymbol{X}_{(b_{1})}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{X}_{(b_{2})}=\boldsymbol{0}_{m_{b_{1}}\times m_{b_{2}}}.$
+  2. A2)
+    (Bounded block sizes) $m_{b}\;\leq\;m_{\max}$ for all $b\in[B]$ and some constant $m_{\max}=O(1)$.
+  3. A3)
+    (Within-block regularity) There exist constants $\tau_{\min}^{2},\tau_{\max}^{2},\rho_{1},\rho_{\infty}\in(0,1)$ such that, for every block $b\in[B]$ and every $j\in[m_{b}]$,
+    $\lambda_{\min}(\Sigma_{(b)})>0,$
+    $\tau_{\min}^{2}\;<\;1-\Sigma_{(b);j,-j}\,\Sigma_{(b);-j,-j}^{-1}\,\Sigma_{(b);-j,j}\;<\;\tau_{\max}^{2},$
+    and
+    $\max_{b\in[B]}\;\max_{j\in(b)}\big\|\Sigma_{(b);-j,-j}^{-1}\Sigma_{(b);-j,j}\big\|_{1}\;<\;\rho_{1},\qquad\max_{b\in[B]}\;\max_{\begin{subarray}{c}j,k\in[m_{b}]\\ j\neq k\end{subarray}}\big|\Sigma_{(b);j,k}\big|\;<\;\rho_{\infty}.$
+The next lemma provides a simple example in which Assumption A3 holds.
+Lemma E.1.
+If $\Sigma_{(b)}=(1-\rho_{b})I_{m_{b}}+\rho_{b}1_{m_{b}}1_{m_{b}}^{\mkern-1.5mu\mathsf{T}}$ for some $|\rho_{b}|<1/(m_{b}-1)$, then Assumption A3 holds with any
+$\tau_{\min}^{2}<\min_{b\in[B]}(1-\rho_{b})\frac{(m_{b}-1)\rho_{b}+1}{(m_{b}-2)\rho_{b}+1}<\max_{b\in[B]}(1-\rho_{b})\frac{(m_{b}-1)\rho_{b}+1}{(m_{b}-2)\rho_{b}+1}<\tau_{\max}^{2}.$
+and
+$\rho_{1}>\max_{b\in[B]}\bigg|\frac{(m_{b}-1)\rho_{b}}{(m_{b}-2)\rho_{b}+1}\bigg|,\quad\rho_{\infty}>\max_{b\in[B]}|\rho_{b}|.$
+We now define the Lasso estimator. Let $\lambda=\lambda_{d}\to\infty$ be a regularity level that may depend on $d$. The (global) Lasso solution is
+$\hat{\boldsymbol{\theta}}^{\lambda}=\operatorname*{argmin}_{\boldsymbol{\gamma}\in\mathbb{R}^{d}}\sum_{b=1}^{B}\bigg\{\frac{1}{2}\,\boldsymbol{\gamma}_{(b)}^{\mkern-1.5mu\mathsf{T}}\Sigma_{(b)}\boldsymbol{\gamma}_{(b)}-\boldsymbol{\gamma}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Z}_{(b)}\bigg\}+\lambda\|\boldsymbol{\gamma}\|_{1}.$
+Equivalently, by block orthogonality,
+$\hat{\boldsymbol{\theta}}^{\lambda}=\operatorname*{argmin}_{\boldsymbol{\gamma}\in\mathbb{R}^{d}}\sum_{b=1}^{B}\bigg\{\frac{1}{2}\,\boldsymbol{\gamma}_{(b)}^{\mkern-1.5mu\mathsf{T}}\Sigma_{(b)}\boldsymbol{\gamma}_{(b)}-\boldsymbol{\gamma}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Z}_{(b)}+\lambda\|\boldsymbol{\gamma}_{(b)}\|_{1}\bigg\},$
+so that the problem separates across blocks:
+$\hat{\boldsymbol{\theta}}^{\lambda}=\big(\hat{\boldsymbol{\theta}}_{(b)}^{\lambda}\big)_{b=1}^{B},\qquad\hat{\boldsymbol{\theta}}_{(b)}^{\lambda}=\operatorname*{argmin}_{\boldsymbol{\gamma}_{(b)}\in\mathbb{R}^{m_{b}}}\bigg\{\frac{1}{2}\,\boldsymbol{\gamma}_{(b)}^{\mkern-1.5mu\mathsf{T}}\Sigma_{(b)}\boldsymbol{\gamma}_{(b)}-\boldsymbol{\gamma}_{(b)}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Z}_{(b)}+\lambda\|\boldsymbol{\gamma}_{(b)}\|_{1}\bigg\}.$
+We denote the total number of selected variables by
+$R=\|\hat{\boldsymbol{\theta}}^{\lambda}\|_{0}=\sum_{b=1}^{B}R_{(b)},\qquad R_{(b)}=\|\hat{\boldsymbol{\theta}}_{(b)}^{\lambda}\|_{0}=\sum_{j\in[m_{b}]}\mathbf{1}\{\hat{\theta}^{\lambda}_{(b);j}\neq 0\}.$
+For any coordinate $j$ in block $b$, let $\boldsymbol{Z}_{(b);-j}\in\mathbb{R}^{m_{b}-1}$ denote the vector $\boldsymbol{Z}_{(b)}$ with its $j$th entry removed.
+Let $\boldsymbol{Z}_{-(b)}=\left\{\boldsymbol{Z}_{(b^{\prime})}\right\}_{b^{\prime}\neq b}$ denote the $\boldsymbol{Z}$ blocks except for $b$.
+Then the null sufficient statistic ($\boldsymbol{S}_{j}$ in the main text) for the variable $j$ in block $b$ is
+$\boldsymbol{S}_{(b);j}=(\boldsymbol{Z}_{(b);-j},\boldsymbol{Z}_{-(b)}).$
+We write $p_{(b);j}$ for a $p$-value constructed so that, under the null $\theta_{(b),j}=0$, $p_{(b);j}$ is (conditionally) independent of $\boldsymbol{S}_{(b);-j}$, taking the usual $z$-test p-value as an example.
+In addition, Assumption A1 implies that, for any $j$ in block $b$,
+$Z_{(b);j}\;\perp\;\boldsymbol{Z}_{-(b)}\;\big|\;\boldsymbol{Z}_{(b);-j}.$
+We further define, for a fixed threshold $\zeta\in(0,1)$,
+$\phi_{(b);j}(p_{(b);j})$ $=\frac{\mathbf{1}\{p_{(b);j}>\zeta\}}{1-\zeta},$
+$\pi_{(b);j}\big(\boldsymbol{S}_{(b);-j}\big)$ $=\mathbb{P}_{\theta_{(b),j}=0}\!\left\{\hat{\theta}_{(b),j}^{\lambda}\neq 0\;\middle|\;\boldsymbol{S}_{(b);-j}\right\}.$
+Finally, we categorize blocks into three types. Fix any constant $\xi>0$ and define:
+  • Null blocks $\mathcal{B}_{0}$: $b\in\mathcal{B}_{0}$ if and only if $\boldsymbol{\theta}_{(b)}=\boldsymbol{0}$.
+  • “Good” non-null blocks $\mathcal{B}_{1g}$: $b\in\mathcal{B}_{1g}$ if and only if there exists $j\in[m_{b}]$ such that $|\theta_{(b);j}|>(1+\xi)\lambda$ and $\theta_{(b);k}=0$ for all $k\neq j$ in $(b)$.
+  • “Bad” non-null blocks $\mathcal{B}_{1b}$: blocks that are neither null nor good non-null, i.e. $\mathcal{B}_{1b}=[B]\setminus(\mathcal{B}_{0}\cup\mathcal{B}_{1g})$.
+Let $B_{0}$, $B_{1g}=B_{1g}(\xi)$, and $B_{1b}=B_{1b}(\xi)$ denote the number of blocks of these three types, respectively. We impose the following additional assumptions on the asymptotic regime.
+  1. A4)
+    (Composition of three types of blocks)
+    $\frac{B_{1g}}{B_{0}}\;=\;o(1),\quad\frac{B_{1b}}{B_{1g}}\;=\;o(1),\quad\text{and}\quad B_{1g}\to\infty\qquad\text{as }B\to\infty.$
+  2. A5)
+    (Choice of the penalty level) There exists $\chi>0$ such that, for each $d$,
+    $\lambda=\lambda_{d}=\Phi^{-1}\!\left(1-\frac{v_{d}}{2d}\right),$
+    for some $v$ satisfying
+    $v_{d}\in\left[\frac{\chi}{1-\chi}\,B_{1g},\;\frac{1-\chi}{\chi}\,B_{1g}\right],$
+    where $\Phi$ is the standard normal distribution function.
+E.2 Main results
+Theorem E.1.
+Let
+$\alpha_{d}\;=\;\frac{v_{d}}{v_{d}+B_{1g}}.$
+Under Assumptions A1–A5,
+$d_{K}\left(\sqrt{\frac{v_{d}}{\alpha_{d}^{3}}}\;(\widehat{\textnormal{FDR}}-\tilde{\alpha}_{d}),\mathcal{N}(0,1)\right)=o(1).$
+where $d_{K}$ denote s the Kolmogorov distance and $\tilde{\alpha}_{d}$ is deterministic and satisfies $\tilde{\alpha}_{d}=\alpha_{d}(1+o(1))$ as $d\to\infty$.
+Theorem E.2.
+Suppose we can write the number of “good” non-null blocks
+$B_{1g}=O\!\bigl(d^{\nu}\,\Gamma_{d}\bigr)$
+for some $\nu\in[0,1)$ and a sequence $\Gamma_{d}$ with $\Gamma_{d}=o(d^{a})$ for any $a>0$.
+Consider the parametric bootstrap estimator $\widehat{\textnormal{FDR}}^{(*)}$ sampled from the OLS-after-Lasso model, where the Lasso penalty to define the OLS model is
+$\bar{\lambda}=\bar{c}\,\lambda$
+for some constant $\bar{c}>1$, with $\lambda$ definded as in Assumption A5. Assume that the signal strength parameter $\xi$ in the definition of “good” non-null blocks satisfies
+$\xi>\sqrt{\frac{\nu}{1-\nu}}+\bar{c}-1.$
+Then, in the setting of Theorem E.1, there exists a (data–dependent) centering constant $\alpha_{d}^{(*)}$ such that
+$\sup_{t\in\mathbb{R}}\biggl|\mathbb{P}\!\biggl(\sqrt{\frac{v_{d}}{\alpha_{d}^{3}}}\,\bigl(\widehat{\textnormal{FDR}}^{(*)}-\alpha_{d}^{(*)}\bigr)\leq t\,\Big|\,\boldsymbol{X},\boldsymbol{Y}\biggr)-\Phi(t)\biggr|\;\xrightarrow{P}\;0.$
+In particular, conditional on the data, the distribution of $\sqrt{v_{d}/\alpha_{d}^{3}}\,(\widehat{\textnormal{FDR}}^{(*)}-\alpha_{d}^{(*)})$ converges to $\mathcal{N}(0,1)$, with convergence in probability over the data-generating process.
+We prove the theorems in the following subsections. We start with supporting lemmas and then present the proofs of Theorems E.1 and E.2.
+E.3 Preparation: single-block analysis
+In this section, we focus on a single block. Fix a block $b$ and, for notational convenience, we suppress the subscript $(b)$ and the superscript $\lambda$. We write
+$m=m_{b},\quad\Sigma=\Sigma_{(b)},\quad\boldsymbol{\theta}=\boldsymbol{\theta}_{(b)},\quad\hat{\boldsymbol{\theta}}=\hat{\boldsymbol{\theta}}_{(b)},\quad\boldsymbol{Z}=\boldsymbol{Z}_{(b)}=\boldsymbol{X}_{(b)}^{\top}\boldsymbol{Y}.$
+Fix any constant $\eta>0$. Then, with probability $1$,
+$|Z_{j}|\not\in\{(1-\eta)\lambda,\lambda,(1+\eta)\lambda\}\quad\text{for all }j\in[m],$
+and we assume this event throughout the subsection.
+E.3.1 Deterministic properties of $\hat{\theta}$
+Lemma E.2.
+Under Assumption A3, $\hat{\boldsymbol{\theta}}$ is the unique solution to the KKT conditions:
+$Z_{j}-\sum_{k=1}^{m}\Sigma_{jk}\hat{\theta}_{k}=\lambda\,{\textnormal{sign}}(\hat{\theta}_{j}),\quad\text{if }\hat{\theta}_{j}\neq 0,$ (17)
+and
+$\bigg|Z_{j}-\sum_{k=1}^{m}\Sigma_{jk}\hat{\theta}_{k}\bigg|<\lambda,\quad\text{if }\hat{\theta}_{j}=0.$ (18)
+Proof.
+The KKT conditions are necessary for optimality.
+Under Assumption A3, $\Sigma$ is positive definite, so the Lasso objective is strongly convex.
+Hence any solution to the KKT conditions is the unique global minimizer.
+∎
+Lemma E.3.
+Under Assumption A3, $\hat{\boldsymbol{\theta}}=\boldsymbol{0}$ if and only if $\|\boldsymbol{Z}\|_{\infty}<\lambda$.
+Proof.
+($\Rightarrow$) If $\hat{\boldsymbol{\theta}}=\boldsymbol{0}$, then (18) gives
+$|Z_{j}|=\Big|Z_{j}-\sum_{k=1}^{m}\Sigma_{jk}\hat{\theta}_{k}\Big|<\lambda\quad\text{for all }j,$
+so $\|\boldsymbol{Z}\|_{\infty}<\lambda$.
+($\Leftarrow$) If $|Z_{j}|<\lambda$ for all $j$, then $\boldsymbol{\theta}=\boldsymbol{0}$ satisfies (18) for every $j$.
+By Lemma E.2 and strong convexity, $\hat{\boldsymbol{\theta}}$ must be this unique solution, so $\hat{\boldsymbol{\theta}}=\boldsymbol{0}$.
+∎
+Lemma E.4.
+Let $\hat{\boldsymbol{\theta}}^{(-j)}$ denote the leave-one-out Lasso estimate that omits the $j$th coordinate:
+$\hat{\boldsymbol{\theta}}^{(-j)}=\operatorname*{argmin}_{\boldsymbol{\gamma}\in\mathbb{R}^{m-1}}\Big\{\boldsymbol{\gamma}^{\top}\boldsymbol{Z}_{-j}-\tfrac{1}{2}\boldsymbol{\gamma}^{\top}\Sigma_{-j,-j}\boldsymbol{\gamma}+\lambda\|\boldsymbol{\gamma}\|_{1}\Big\}.$
+Under Assumption A3,
+$\hat{\theta}_{j}=0\quad\Longleftrightarrow\quad\big|Z_{j}-\Sigma_{j,-j}\hat{\boldsymbol{\theta}}^{(-j)}\big|<\lambda.$
+Proof.
+Under Assumption A3, $\Sigma_{-j,-j}$ is positive definite, so $\hat{\boldsymbol{\theta}}^{(-j)}$ is unique.
+($\Rightarrow$) If $\hat{\theta}_{j}=0$, then the subvector $\hat{\boldsymbol{\theta}}_{-j}$ solves the (m1)(m-1)-dimensional Lasso with data $(\boldsymbol{Z}_{-j},\Sigma_{-j,-j})$, hence $\hat{\boldsymbol{\theta}}_{-j}=\hat{\boldsymbol{\theta}}^{(-j)}$ by uniqueness.
+Then (18) gives
+$\big|Z_{j}-\Sigma_{j,-j}\hat{\boldsymbol{\theta}}^{(-j)}\big|=\bigg|Z_{j}-\sum_{k=1}^{m}\Sigma_{jk}\hat{\theta}_{k}\bigg|<\lambda.$
+($\Leftarrow$) Now assume $|Z_{j}-\Sigma_{j,-j}\hat{\boldsymbol{\theta}}^{(-j)}|<\lambda$, and define $\tilde{\boldsymbol{\theta}}$ by
+$\tilde{\theta}_{j}=0,\qquad\tilde{\boldsymbol{\theta}}_{-j}=\hat{\boldsymbol{\theta}}^{(-j)}.$
+By construction, $\tilde{\boldsymbol{\theta}}_{-j}$ satisfies the KKT conditions for coordinates $k\neq j$, and the inequality for coordinate $j$ holds by assumption.
+Thus $\tilde{\boldsymbol{\theta}}$ satisfies the full KKT system (17)–(18).
+By Lemma E.2 and uniqueness, $\hat{\boldsymbol{\theta}}=\tilde{\boldsymbol{\theta}}$, so $\hat{\theta}_{j}=0$.
+∎
+Lemma E.5.
+Under Assumption A3, If $|Z_{j}|>\lambda$ and $\|\boldsymbol{Z}_{-j}\|_{\infty}<\lambda$, then $\hat{\theta}_{j}\neq 0$.
+Proof.
+By Lemma E.3, $\|\boldsymbol{Z}_{-j}\|_{\infty}<\lambda$ implies $\hat{\boldsymbol{\theta}}^{(-j)}=\boldsymbol{0}$. Then Lemma E.4 says
+$\hat{\theta}_{j}=0\Longleftrightarrow|Z_{j}-\Sigma_{j,-j}\hat{\boldsymbol{\theta}}^{(-j)}|<\lambda\Longleftrightarrow|Z_{j}|<\lambda,$
+which contradicts $|Z_{j}|>\lambda$. Hence $\hat{\theta}_{j}\neq 0$. ∎
+Lemma E.6.
+Fix any $\eta>0$. Under Assumption A3, if
+$|Z_{j}|\in(\lambda,(1+\eta)\lambda)\quad\text{and}\quad\|\boldsymbol{Z}_{-j}\|_{\infty}<(1-\eta\rho_{\infty})\lambda,$
+then $\hat{\theta}_{j}\neq 0$ and $\hat{\boldsymbol{\theta}}_{-j}=\boldsymbol{0}$.
+Proof.
+Consider the candidate solution $\tilde{\boldsymbol{\theta}}$ with
+$\tilde{\theta}_{j}=(|Z_{j}|-\lambda)\,{\textnormal{sign}}(Z_{j}),\quad\tilde{\theta}_{k}=0,\ \forall k\neq j.$
+Then
+$Z_{j}-\Sigma_{jj}\tilde{\theta}_{j}=Z_{j}-\tilde{\theta}_{j}=\lambda\,{\textnormal{sign}}(\tilde{\theta}_{j}),$
+since $\Sigma_{jj}=1$. For $k\neq j$, Assumption A3 implies $|\Sigma_{kj}|\leq\rho_{\infty}$, and we have $|\tilde{\theta}_{j}|=|Z_{j}|-\lambda<\eta\lambda$, so
+$\big|Z_{k}-\Sigma_{kj}\tilde{\theta}_{j}\big|$ $\leq|Z_{k}|+|\Sigma_{kj}|\,|\tilde{\theta}_{j}|<(1-\eta\rho_{\infty})\lambda+\rho_{\infty}\eta\lambda=\lambda.$
+Thus $\tilde{\boldsymbol{\theta}}$ satisfies the KKT conditions (17)–(18).
+By Lemma E.2, $\hat{\boldsymbol{\theta}}=\tilde{\boldsymbol{\theta}}$, so $\hat{\theta}_{j}\neq 0$ and $\hat{\theta}_{k}=0$ for all $k\neq j$.
+∎
+Lemma E.7.
+Fix any $\eta>0$. Under Assumption A3, if $\|\hat{\boldsymbol{\theta}}\|_{0}\geq 2$, then one of the following holds:
+  1. (a)
+    $\|\boldsymbol{Z}\|_{\infty}>(1+\eta)\lambda$;
+  2. (b)
+    there exist $k_{1}<k_{2}$ such that
+    $|Z_{k_{1}}|,|Z_{k_{2}}|\in\big((1-\eta)\lambda,(1+\eta)\lambda\big).$
+Proof.
+Assume neither (a) nor (b) holds. Let $j=\operatorname*{argmax}_{k\in[m]}|Z_{k}|$, so
+$|Z_{j}|<(1+\eta)\lambda,\qquad\|\boldsymbol{Z}_{-j}\|_{\infty}<(1-\eta)\lambda.$
+If $|Z_{j}|<\lambda$, Lemma E.3 implies $\|\hat{\boldsymbol{\theta}}\|_{0}=0$, contrary to the assumption $\|\hat{\boldsymbol{\theta}}\|_{0}\geq 2$. Thus $|Z_{j}|>\lambda$. Since $\rho_{\infty}<1$,
+$\|\boldsymbol{Z}_{-j}\|_{\infty}<(1-\eta)\lambda<(1-\eta\rho_{\infty})\lambda.$
+By Lemma E.6, $\hat{\theta}_{j}\neq 0$ and $\hat{\theta}_{k}=0$ for all $k\neq j$, so $\|\hat{\boldsymbol{\theta}}\|_{0}=1$, again a contradiction. Hence at least one of (a)–(b) must hold. ∎
+E.3.2 Stochastic properties of $\hat{\theta}$
+Throughout this subsection, let $\check{Z}\sim\mathcal{N}(0,\boldsymbol{I})$ denote a generic standard Gaussian random vector.
+Proposition E.1 (Gordon (1941)).
+For any $z>0$,
+$\frac{1}{\sqrt{2\pi}}\,\frac{z}{z^{2}+1}\exp\Big(-\frac{z^{2}}{2}\Big)\;\leq\;1-\Phi(z)\;\leq\;\frac{1}{\sqrt{2\pi}}\,\frac{1}{z}\exp\Big(-\frac{z^{2}}{2}\Big).$
+Lemma E.8.
+For any $z,\tau>0$ and $\mu\in\mathbb{R}$,
+$\mathbb{P}(|\tau\check{Z}+\mu|>z)\;\leq\;2\Big\{1-\Phi\Big(\frac{z-|\mu|}{\tau}\Big)\Big\}.$
+Proof.
+We have
+$\mathbb{P}(|\tau\check{Z}+\mu|>z)$ $=\mathbb{P}\Big(\check{Z}>\frac{z-\mu}{\tau}\Big)+\mathbb{P}\Big(\check{Z}<\frac{-z-\mu}{\tau}\Big)$
+$=1-\Phi\Big(\frac{z-\mu}{\tau}\Big)+\Phi\Big(\frac{-z-\mu}{\tau}\Big).$
+If $\mu\geq 0$, then $z-\mu\leq z-|\mu|$ and $z-\mu\leq-z-|\mu|$, so
+$1-\Phi\Big(\frac{z-\mu}{\tau}\Big)\leq 1-\Phi\Big(\frac{z-|\mu|}{\tau}\Big),\quad\Phi\Big(\frac{-z-\mu}{\tau}\Big)\leq 1-\Phi\Big(\frac{z-|\mu|}{\tau}\Big),$
+and similarly for $\mu<0$ by symmetry. Thus each term is bounded by $1-\Phi((z-|\mu|)/\tau)$ and the result follows. ∎
+Lemma E.9.
+Under Assumption A3, for any $j\in[m]$ and any $\boldsymbol{z}\in\mathbb{R}^{m-1}$,
+$\pi_{j}(\boldsymbol{z})$ $\leq\mathbf{1}\{\|\boldsymbol{z}\|_{\infty}<\lambda\}\cdot 2\Big\{1-\Phi\Big(\frac{\lambda-\big|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\boldsymbol{z}\big|}{\tau_{j}}\Big)\Big\}$
+$\quad+\mathbf{1}\{\|\boldsymbol{z}\|_{\infty}\geq\lambda\}\cdot 2\Big\{1-\Phi\Big(\frac{\lambda(1-\rho_{1})}{\tau_{j}}\Big)\Big\},$
+where
+$\tau_{j}^{2}=1-\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\Sigma_{-j,j}.$
+Proof.
+By Lemma E.4 and the definition of $\pi_{j}(\cdot)$,
+$\pi_{j}(\boldsymbol{z})=\mathbb{P}_{\theta_{j}=0}\big(|Z_{j}-\Sigma_{j,-j}\hat{\boldsymbol{\theta}}^{(-j)}|>\lambda\mid\boldsymbol{Z}_{-j}=\boldsymbol{z}\big).$
+Let $\hat{\boldsymbol{\theta}}^{(-j)}(\boldsymbol{z})$ denote the leave-one-out Lasso estimator when $\boldsymbol{Z}_{-j}=\boldsymbol{z}$.
+By Lemma E.3, if $\|\boldsymbol{z}\|_{\infty}<\lambda$ then $\hat{\boldsymbol{\theta}}^{(-j)}(\boldsymbol{z})=\boldsymbol{0}$, so
+$\pi_{j}(\boldsymbol{z})=\mathbb{P}_{\theta_{j}=0}(|Z_{j}|>\lambda\mid\boldsymbol{Z}_{-j}=\boldsymbol{z}).$
+Conditional on $\boldsymbol{Z}_{-j}=\boldsymbol{z}$,
+$Z_{j}\mid\boldsymbol{Z}_{-j}=\boldsymbol{z}\overset{\theta_{j}=0}{\sim}\mathcal{N}\left(\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\boldsymbol{z},\,\tau_{j}^{2}\right),$
+so
+$\pi_{j}(\boldsymbol{z})=\mathbb{P}\Big(|\tau_{j}\check{Z}+\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\boldsymbol{z}|>\lambda\Big),$
+and Lemma E.8 yields
+$\pi_{j}(\boldsymbol{z})\leq 2\Big\{1-\Phi\Big(\frac{\lambda-\big|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\boldsymbol{z}\big|}{\tau_{j}}\Big)\Big\}.$
+Now consider the case $\|\boldsymbol{z}\|_{\infty}\geq\lambda$. By Lemma E.4, the same conditional law implies
+$\pi_{j}(\boldsymbol{z})=\mathbb{P}\Big(\big|\tau_{j}\check{Z}+\Sigma_{j,-j}\big(\Sigma_{-j,-j}^{-1}\boldsymbol{z}-\hat{\boldsymbol{\theta}}^{(-j)}(\boldsymbol{z})\big)\big|>\lambda\Big).$
+By Lemma E.2, there exists $v\in[-1,1]^{m-1}$ such that
+$\boldsymbol{z}-\Sigma_{-j,-j}\hat{\boldsymbol{\theta}}^{(-j)}(\boldsymbol{z})=\lambda v,$
+so $\Sigma_{-j,-j}^{-1}\boldsymbol{z}-\hat{\boldsymbol{\theta}}^{(-j)}(\boldsymbol{z})=\lambda\Sigma_{-j,-j}^{-1}v$, and hence
+$\Sigma_{j,-j}\big(\Sigma_{-j,-j}^{-1}\boldsymbol{z}-\hat{\boldsymbol{\theta}}^{(-j)}(\boldsymbol{z})\big)=\lambda\,\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}v.$
+Assumption A3 gives
+$\big|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}v\big|\leq\|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\|_{1}\|v\|_{\infty}\leq\rho_{1}.$
+Thus, by Lemma E.8,
+$\pi_{j}(\boldsymbol{z})\leq 2\left\{1-\Phi\left(\frac{\lambda(1-\rho_{1})}{\tau_{j}}\right)\right\}.$
+Combining the two cases gives the stated bound. ∎
+Lemma E.10.
+Under Assumption A3, if $\lambda>1$, then for any $j\in[m]$ and any $L>0$,
+$\mathbb{E}\big[\pi_{j}^{L}(\boldsymbol{Z}_{-j})\big]\leq\left(\frac{2}{(1-\rho_{1})\lambda}\right)^{L}\exp\left\{-\frac{L(1-\rho_{1})^{2}}{2\,\tau_{\max}^{2}}\lambda^{2}\right\}.$
+If, further, $\boldsymbol{\theta}_{-j}=\boldsymbol{0}$, then
+$\mathbb{E}[\pi_{j}(\boldsymbol{Z}_{-j})]\;\leq\;(1-\Phi(\lambda))\cdot\frac{12m}{1-\rho_{1}},$
+and
+$\mathbb{E}[\pi_{j}^{2}(\boldsymbol{Z}_{-j})]\;\leq\;(1-\Phi(\lambda))\cdot\frac{20m}{(1-\rho_{1})^{2}\lambda}\exp\left\{-\left(\frac{(1-\rho_{1})^{2}}{\tau_{\max}^{2}}\wedge\frac{\tau_{\min}^{2}}{2(2-\tau_{\min}^{2})}\right)\lambda^{2}\right\}.$
+Proof.
+First, consider the almost-sure bound. For any $z$ with $\|z\|_{\infty}<\lambda$,
+$\big|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}z\big|\leq\|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\|_{1}\|z\|_{\infty}<\rho_{1}\lambda.$ (19)
+Thus Lemma E.9 and Proposition E.1 imply
+$\pi_{j}(z)$ $\leq 2\left\{1-\Phi\left(\frac{\lambda(1-\rho_{1})}{\tau_{j}}\right)\right\}$
+$\leq\frac{2\tau_{j}}{(1-\rho_{1})\lambda}\exp\Big\{-\frac{(1-\rho_{1})^{2}}{2\tau_{j}^{2}}\lambda^{2}\Big\}<\frac{2}{(1-\rho_{1})\lambda}\exp\Big\{-\frac{(1-\rho_{1})^{2}}{2\tau_{j}^{2}}\lambda^{2}\Big\}.$
+By Assumption A3, $\tau_{j}^{2}\leq\tau^{2}_{\max}$, so
+$\pi_{j}(z)\leq\frac{2}{(1-\rho_{1})\lambda}\exp\Big\{-\frac{(1-\rho_{1})^{2}}{2\,\tau_{\max}^{2}}\lambda^{2}\Big\}.$
+Raising to the power $L$ and taking expectations gives the first bound.
+Now assume $\boldsymbol{\theta}_{-j}=\boldsymbol{0}$. Since $\boldsymbol{\theta}_{-j}=\boldsymbol{0}$, we have
+$\boldsymbol{Z}_{-j}\sim\mathcal{N}\big(\boldsymbol{0},\,\Sigma_{-j,-j}\big),$
+and therefore
+[…]
+Let $\check{Z}\sim\mathcal{N}(0,1)$ be a generic standard Gaussian. Then
+$\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\boldsymbol{Z}_{-j}\;\stackrel{{\scriptstyle d}}{{=}}\;\sqrt{1-\tau_{j}^{2}}\,\check{Z}.$
+We now derive refined bounds for the first and second moments of $\pi_{j}(\boldsymbol{Z}_{-j})$ using the decomposition from Lemma E.9. Recall that
+$\pi_{j}(\boldsymbol{z}_{-j})$ $\leq\mathbf{1}\big(\|\boldsymbol{z}_{-j}\|_{\infty}<\lambda\big)\cdot 2\Big\{1-\Phi\Big(\frac{\lambda-\big|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\boldsymbol{z}_{-j}\big|}{\tau_{j}}\Big)\Big\}$
+$\quad+\mathbf{1}\big(\|\boldsymbol{z}_{-j}\|_{\infty}>\lambda\big)\cdot 2\Big\{1-\Phi\Big(\frac{\lambda(1-\rho_{1})}{\tau_{j}}\Big)\Big\}.$
+Refined first moment bound under $\boldsymbol{\theta}_{-j}=\boldsymbol{0}$. On the event $\{\|\boldsymbol{Z}_{-j}\|_{\infty}<\lambda\}$, Assumption A3 implies
+$\Big|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\boldsymbol{Z}_{-j}\Big|\leq\big\|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\big\|_{1}\,\|\boldsymbol{Z}_{-j}\|_{\infty}<\rho_{1}\lambda.$ (20)
+Hence, from Lemma E.9,
+$\mathbb{E}\Big[\mathbf{1}\big(\|\boldsymbol{Z}_{-j}\|_{\infty}<\lambda\big)\cdot 2\Big\{1-\Phi\Big(\frac{\lambda-\big|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\boldsymbol{Z}_{-j}\big|}{\tau_{j}}\Big)\Big\}\Big]$
+$\leq\mathbb{E}\Big[\mathbf{1}\Big(\big|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\boldsymbol{Z}_{-j}\big|<\rho_{1}\lambda\Big)\cdot 2\Big\{1-\Phi\Big(\frac{\lambda-\big|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\boldsymbol{Z}_{-j}\big|}{\tau_{j}}\Big)\Big\}\Big]$
+$=\mathbb{E}\Big[\mathbf{1}\big(\sqrt{1-\tau_{j}^{2}}\,|\check{Z}|<\rho_{1}\lambda\big)\cdot 2\Big\{1-\Phi\Big(\frac{\lambda-\sqrt{1-\tau_{j}^{2}}\,|\check{Z}|}{\tau_{j}}\Big)\Big\}\Big]$
+$=4\,\mathbb{E}\Big[\mathbf{1}\big(0<\sqrt{1-\tau_{j}^{2}}\,\check{Z}<\rho_{1}\lambda\big)\cdot\Big\{1-\Phi\Big(\frac{\lambda-\sqrt{1-\tau_{j}^{2}}\,\check{Z}}{\tau_{j}}\Big)\Big\}\Big],$
+where we used symmetry of $\check{Z}$ in the last line.
+Let $L\geq 1$ be an integer (we will specialize to $L=1,2$ later).
+Applying the upper bound from the Mills ratio (Proposition E.1) to the tail, for any $z\in\mathbb{R}$ such that $0<\sqrt{1-\tau_{j}^{2}}z<\rho_{1}\lambda$, we have
+$1-\Phi\Big(\frac{\lambda-\sqrt{1-\tau_{j}^{2}}\,z}{\tau_{j}}\Big)\leq\frac{\tau_{j}}{\lambda-\sqrt{1-\tau_{j}^{2}}\,z}\cdot\frac{1}{\sqrt{2\pi}}\exp\Big\{-\frac{(\lambda-\sqrt{1-\tau_{j}^{2}}\,z)^{2}}{2\tau_{j}^{2}}\Big\}.$
+Using again that $\lambda-\sqrt{1-\tau_{j}^{2}}\,z\geq(1-\rho_{1})\lambda$ on this region,
+$1-\Phi\Big(\frac{\lambda-\sqrt{1-\tau_{j}^{2}}\,z}{\tau_{j}}\Big)\leq\frac{\tau_{j}}{(1-\rho_{1})\lambda}\cdot\frac{1}{\sqrt{2\pi}}\exp\Big\{-\frac{(\lambda-\sqrt{1-\tau_{j}^{2}}\,z)^{2}}{2\tau_{j}^{2}}\Big\}.$
+Therefore, for any integer $L\geq 1$,
+$\quad\frac{1}{2}\Big(\frac{\sqrt{2\pi}(1-\rho_{1})\lambda}{2\tau_{j}}\Big)^{L}\cdot\mathbb{E}\Big[\mathbf{1}\big(\|\boldsymbol{Z}_{-j}\|_{\infty}<\lambda\big)\cdot 2^{L}\Big\{1-\Phi\Big(\frac{\lambda-\big|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\boldsymbol{Z}_{-j}\big|}{\tau_{j}}\Big)\Big\}^{L}\Big]$
+$\leq\frac{1}{\sqrt{2\pi}}\int_{0}^{\frac{\rho_{1}}{\sqrt{1-\tau_{j}^{2}}}\lambda}\exp\Big\{-\frac{L(\lambda-\sqrt{1-\tau_{j}^{2}}z)^{2}}{2\tau_{j}^{2}}-\frac{z^{2}}{2}\Big\}\,dz$
+$=\frac{1}{\sqrt{2\pi}}\int_{0}^{\frac{\rho_{1}}{\sqrt{1-\tau_{j}^{2}}}\lambda}\exp\left\{-\frac{L\lambda^{2}-2L\sqrt{1-\tau_{j}^{2}}\lambda z+(L-(L-1)\tau_{j}^{2})z^{2}}{2\tau_{j}^{2}}\right\}dz$
+$=\frac{1}{\sqrt{2\pi}}\int_{0}^{\frac{\rho_{1}}{\sqrt{1-\tau_{j}^{2}}}\lambda}\exp\left\{-\frac{L-(L-1)\tau_{j}^{2}}{2\tau_{j}^{2}}\left(z-\frac{L\sqrt{1-\tau_{j}^{2}}}{L-(L-1)\tau_{j}^{2}}\lambda\right)^{2}\right\}dz$
+$\qquad\cdot\exp\left\{\frac{L^{2}(1-\tau_{j}^{2})}{2\tau_{j}^{2}\{L-(L-1)\tau_{j}^{2}\}}\lambda^{2}-\frac{L}{2\tau_{j}^{2}}\lambda^{2}\right\}$
+$=\frac{1}{\sqrt{2\pi}}\int_{-\frac{L\sqrt{1-\tau_{j}^{2}}}{L-(L-1)\tau_{j}^{2}}\lambda}^{\frac{\rho_{1}}{\sqrt{1-\tau_{j}^{2}}}\lambda-\frac{L\sqrt{1-\tau_{j}^{2}}}{L-(L-1)\tau_{j}^{2}}\lambda}\exp\left\{-\frac{L-(L-1)\tau_{j}^{2}}{2\tau_{j}^{2}}z^{2}\right\}dz$
+$\qquad\cdot\exp\left\{-\frac{L}{2\{L-(L-1)\tau_{j}^{2}\}}\lambda^{2}\right\}$
+$\leq\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{-\frac{((1-\rho_{1})L+\rho_{1})(1-\tau_{j}^{2})-\rho_{1}}{\sqrt{1-\tau_{j}^{2}}\{L-(L-1)\tau_{j}^{2}\}}\lambda}\exp\left\{-\frac{L-(L-1)\tau_{j}^{2}}{2\tau_{j}^{2}}z^{2}\right\}dz$
+$\qquad\cdot\exp\left\{-\frac{L}{2\{L-(L-1)\tau_{j}^{2}\}}\lambda^{2}\right\}$
+$=\frac{\tau_{j}}{\sqrt{L-(L-1)\tau_{j}^{2}}}\Bigg[1-\Phi\left(\frac{((1-\rho_{1})L+\rho_{1})(1-\tau_{j}^{2})-\rho_{1}}{\sqrt{(1-\tau_{j}^{2})\{L-(L-1)\tau_{j}^{2}\}}\;\tau_{j}}\,\lambda\right)\Bigg]$
+$\qquad\cdot\exp\left\{-\frac{L}{2\{L-(L-1)\tau_{j}^{2}\}}\lambda^{2}\right\}$
+$\leq\exp\left\{-\frac{L}{2\{L-(L-1)\tau_{j}^{2}\}}\lambda^{2}\right\},$
+where in the last line we used that $\tau<1$.
+Now take $L=1$. We obtain
+$\mathbb{E}\Big[\mathbf{1}\big(\|\boldsymbol{Z}_{-j}\|_{\infty}<\lambda\big)\cdot 2\Big\{1-\Phi\Big(\frac{\lambda-\big|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\boldsymbol{Z}_{-j}\big|}{\tau_{j}}\Big)\Big\}\Big]\;\leq\;\frac{4}{\sqrt{2\pi}(1-\rho_{1})\lambda}\exp\left\{-\frac{\lambda^{2}}{2}\right\}.$
+By the lower bound in the Mills ratio (again Proposition E.1),
+$\frac{\lambda}{\lambda^{2}+1}\,\frac{1}{\sqrt{2\pi}}\exp\left\{-\frac{\lambda^{2}}{2}\right\}\;\leq\;1-\Phi(\lambda).$
+For $\lambda>1$, we have
+$\frac{1}{\sqrt{2\pi}\lambda}\exp\left\{-\frac{\lambda^{2}}{2}\right\}\;\leq\;\frac{\lambda^{2}+1}{\lambda^{2}}\,\big(1-\Phi(\lambda)\big)\;\leq\;2\,\big(1-\Phi(\lambda)\big).$
+Thus,
+$\mathbb{E}\Big[\mathbf{1}\big(\|\boldsymbol{Z}_{-j}\|_{\infty}<\lambda\big)\cdot 2\Big\{1-\Phi\Big(\frac{\lambda-\big|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\boldsymbol{Z}_{-j}\big|}{\tau_{j}}\Big)\Big\}\Big]\;\leq\;\frac{8}{1-\rho_{1}}\,\big(1-\Phi(\lambda)\big).$
+The remaining contribution comes from the event $\{\|\boldsymbol{Z}_{-j}\|_{\infty}>\lambda\}$, for which Lemma E.9 yields
+$\pi_{j}(\boldsymbol{Z}_{-j})\leq 2\Big\{1-\Phi\Big(\frac{\lambda(1-\rho_{1})}{\tau_{j}}\Big)\Big\}$
+and hence
+$\mathbb{E}\big[\pi_{j}(\boldsymbol{Z}_{-j})\big]\leq\frac{8}{1-\rho_{1}}\,\big(1-\Phi(\lambda)\big)+2\,\mathbb{P}\big(\|\boldsymbol{Z}_{-j}\|_{\infty}>\lambda\big)\,\Big\{1-\Phi\Big(\frac{\lambda(1-\rho_{1})}{\tau_{j}}\Big)\Big\}.$
+By the union bound and symmetry,
+$\mathbb{P}\big(\|\boldsymbol{Z}_{-j}\|_{\infty}>\lambda\big)\leq\sum_{k\neq j}\mathbb{P}(|Z_{k}|>\lambda)=2(m-1)\big(1-\Phi(\lambda)\big).$
+Using again the (upper) Mills ratio on the factor $\big\{1-\Phi\big(\lambda(1-\rho_{1})/\tau_{j}\big)\big\}$, and combining constants, we find
+$\mathbb{E}\big[\pi_{j}(\boldsymbol{Z}_{-j})\big]\leq(1-\Phi(\lambda))\left\{\frac{8}{1-\rho_{1}}+4(m-1)\right\}\;\leq\;(1-\Phi(\lambda))\cdot\frac{12m}{1-\rho_{1}},$
+which yields the first refined bound in the lemma.
+Refined second moment bound under $\boldsymbol{\theta}_{-j}=\boldsymbol{0}$. Repeating the same calculation with $L=2$ in the display above, we obtain
+$\mathbb{E}\Big[\mathbf{1}\big(\|\boldsymbol{Z}_{-j}\|_{\infty}<\lambda\big)\cdot 4\Big\{1-\Phi\Big(\frac{\lambda-\big|\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\boldsymbol{Z}_{-j}\big|}{\tau_{j}}\Big)\Big\}^{2}\Big]\;\leq\;\frac{4}{\pi(1-\rho_{1})^{2}\lambda^{2}}\exp\left\{-\frac{\lambda^{2}}{2-\tau_{j}^{2}}\right\}.$
+Using Lemma E.9, we have
+$\mathbb{E}\big[\pi_{j}^{2}(\boldsymbol{Z}_{-j})\big]\leq\frac{4}{\pi(1-\rho_{1})^{2}\lambda^{2}}\exp\left\{-\frac{\lambda^{2}}{2-\tau_{j}^{2}}\right\}+4\,\mathbb{P}\big(\|\boldsymbol{Z}_{-j}\|_{\infty}>\lambda\big)\,\Big\{1-\Phi\Big(\frac{\lambda(1-\rho_{1})}{\tau_{j}}\Big)\Big\}^{2}.$
+Again, the Mills ratio bound gives
+$\Big\{1-\Phi\Big(\frac{\lambda(1-\rho_{1})}{\tau_{j}}\Big)\Big\}^{2}\leq\frac{\tau_{j}^{2}}{(1-\rho_{1})^{2}\lambda^{2}}\exp\left\{-\frac{(1-\rho_{1})^{2}}{\tau_{j}^{2}}\lambda^{2}\right\}.$
+and
+$\frac{1}{\lambda}\exp\left\{-\frac{\lambda^{2}}{2-\tau_{j}^{2}}\right\}=\frac{1}{\lambda}\exp\left\{-\frac{\lambda^{2}}{2}\right\}\cdot\exp\left\{-\frac{\tau_{j}^{2}\lambda^{2}}{2(2-\tau_{j}^{2})}\right\}\leq 6\,\big(1-\Phi(\lambda)\big)\cdot\exp\left\{-\frac{\tau_{j}^{2}\lambda^{2}}{2(2-\tau_{j}^{2})}\right\}$
+Combining these bounds, for $\lambda>1$, we obtain
+$\mathbb{E}\big[\pi_{j}^{2}(\boldsymbol{Z}_{-j})\big]$ $\leq\frac{2}{(1-\rho_{1})^{2}\lambda^{2}}\exp\left\{-\frac{\lambda^{2}}{2-\tau_{j}^{2}}\right\}+\frac{8(m-1)\big(1-\Phi(\lambda)\big)\,\tau_{j}^{2}}{(1-\rho_{1})^{2}\lambda^{2}}\exp\left\{-\frac{(1-\rho_{1})^{2}}{\tau_{j}^{2}}\lambda^{2}\right\}$
+$\leq\left\{\frac{12}{(1-\rho_{1})^{2}\lambda}+\frac{8(m-1)\tau_{j}^{2}}{(1-\rho_{1})^{2}\lambda^{2}}\right\}\big(1-\Phi(\lambda)\big)\,\exp\left\{-\left(\frac{\tau_{j}^{2}}{2(2-\tau_{j}^{2})}\wedge\frac{(1-\rho_{1})^{2}}{\tau_{j}^{2}}\right)\lambda^{2}\right\}$
+$\leq\frac{20m}{(1-\rho_{1})^{2}\lambda}\big(1-\Phi(\lambda)\big)\,\exp\left\{-\left(\frac{\tau_{j}^{2}}{2(2-\tau_{j}^{2})}\wedge\frac{(1-\rho_{1})^{2}}{\tau_{j}^{2}}\right)\lambda^{2}\right\}.$
+Finally, using $\tau_{\min}^{2}\leq\tau_{j}^{2}\leq\tau_{\max}^{2}$ in Assumption A3, we have
+$\mathbb{E}[\pi_{j}^{2}(\boldsymbol{Z}_{-j})]\;\leq\;(1-\Phi(\lambda))\cdot\frac{20m}{(1-\rho_{1})^{2}\lambda}\exp\left\{-\left(\frac{(1-\rho_{1})^{2}}{\tau_{\max}^{2}}\wedge\frac{\tau_{\min}^{2}}{2(2-\tau_{\min}^{2})}\right)\lambda^{2}\right\}.$
+Lemma E.11.
+Assume $\boldsymbol{\theta}=\boldsymbol{0}$. Under Assumption A3, for any $j\neq k$,
+$\mathbb{P}(|Z_{j}|>\lambda,|Z_{k}|>\lambda)\leq 4\Big(1-\Phi\Big(\sqrt{\frac{2}{1+\rho_{\infty}}}\,\lambda\Big)\Big).$
+Proof.
+When $\boldsymbol{\theta}=\boldsymbol{0}$,
+[…]
+Note
+$\mathbb{P}(|Z_{j}|>\lambda,|Z_{k}|>\lambda)\leq\mathbb{P}(|Z_{j}+Z_{k}|>2\lambda)+\mathbb{P}(|Z_{j}-Z_{k}|>2\lambda).$
+Since $Z_{j}+Z_{k}\sim\mathcal{N}(0,2(1+\Sigma_{jk}))$, by symmetry, we have
+$\mathbb{P}(|Z_{j}+Z_{k}|>2\lambda)=2\left(1-\Phi\Big(\sqrt{\tfrac{2}{1+\Sigma_{jk}}}\,\lambda\Big)\right).$
+Similarly $Z_{j}-Z_{k}\sim\mathcal{N}(0,2(1-\Sigma_{jk}))$, we have
+$\mathbb{P}(|Z_{j}-Z_{k}|>2\lambda)=2\left(1-\Phi\Big(\sqrt{\tfrac{2}{1-\Sigma_{jk}}}\,\lambda\Big)\right).$
+Assumption A3 gives $|\Sigma_{jk}|\leq\rho_{\infty}$. Thus
+$\mathbb{P}(|Z_{j}|>\lambda,|Z_{k}|>\lambda)\leq 4\Big(1-\Phi\Big(\sqrt{\tfrac{2}{1+\rho_{\infty}}}\,\lambda\Big)\Big).$
+Lemma E.12.
+Assume $\boldsymbol{\theta}=\boldsymbol{0}$ and $\lambda>1$. Under Assumption A3,
+$\mathbb{P}(\|\hat{\boldsymbol{\theta}}\|_{0}\geq 2)\leq(1-\Phi(\lambda))\cdot 12m^{2}\exp\big\{-\zeta(\rho_{\infty})\lambda^{2}\big\},$
+where
+$\zeta(\rho_{\infty})=\eta(\rho_{\infty})+\tfrac{1}{2}\eta^{2}(\rho_{\infty}),\qquad\eta(\rho_{\infty})=\frac{\sqrt{2}-\sqrt{1+\rho_{\infty}}}{\sqrt{2}+\sqrt{1+\rho_{\infty}}}.$
+Proof.
+By Lemma E.7,
+$\mathbb{P}(\|\hat{\boldsymbol{\theta}}\|_{0}\geq 2)\leq\mathbb{P}(\|\boldsymbol{Z}\|_{\infty}>(1+\eta)\lambda)+\sum_{k_{1}<k_{2}}\mathbb{P}\Big(|Z_{k_{1}}|,|Z_{k_{2}}|\in\big((1-\eta)\lambda,(1+\eta)\lambda\big)\Big).$
+Using the union bound,
+$\mathbb{P}(\|\boldsymbol{Z}\|_{\infty}>(1+\eta)\lambda)\leq\sum_{k=1}^{m}\mathbb{P}(|Z_{k}|>(1+\eta)\lambda)=2m\big\{1-\Phi((1+\eta)\lambda)\big\}.$
+By Lemma E.11,
+$\mathbb{P}\Big(|Z_{k_{1}}|,|Z_{k_{2}}|\in\big((1-\eta)\lambda,(1+\eta)\lambda\big)\Big)$
+$\quad\leq\mathbb{P}\big(|Z_{k_{1}}|>(1-\eta)\lambda,|Z_{k_{2}}|>(1-\eta)\lambda\big)$
+$\quad\leq 4\Big(1-\Phi\Big(\sqrt{\tfrac{2}{1+\rho_{\infty}}}(1-\eta)\lambda\Big)\Big).$
+The choice of $\eta=\eta(\rho_{\infty})$ yields
+$\sqrt{\tfrac{2}{1+\rho_{\infty}}}\,(1-\eta)=1+\eta.$
+Combining the bounds, we have
+$\mathbb{P}(\|\hat{\boldsymbol{\theta}}\|_{0}\geq 2)\leq 2m^{2}\Big\{1-\Phi\big((1+\eta)\lambda\big)\Big\}.$
+Applying Proposition E.1,
+$\mathbb{P}(\|\hat{\theta}\|_{0}\geq 2)$ $\leq\frac{2m^{2}}{(1+\eta)\lambda}\exp\left\{-\frac{(1+\eta)^{2}}{2}\lambda^{2}\right\}$
+$\leq 2m^{2}\cdot\frac{1}{\lambda}\exp\left\{-\frac{\lambda^{2}}{2}\right\}\exp\left\{-\left[\eta+\frac{1}{2}\eta^{2}\right]\lambda^{2}\right\}$
+$\leq(1-\Phi(\lambda))\cdot 12m^{2}\exp\left\{-\zeta\,\lambda^{2}\right\},$
+where the last line uses the condition $\lambda>1$ and the Mill’s ratio lower bound. ∎
+Lemma E.13.
+Under Assumption A3, for any $j$ with $\theta_{j}=0$,
+$\mathbb{P}(\hat{\theta}_{j}\neq 0)\leq\frac{2}{(1-\rho_{1})\lambda}\exp\left\{-\frac{(1-\rho_{1})^{2}}{2\,\tau_{\max}^{2}}\lambda^{2}\right\}.$
+Proof.
+When $\theta_{j}=0$,
+$\mathbb{P}(\hat{\theta}_{j}\neq 0)=\mathbb{E}[\pi_{j}(\boldsymbol{Z}_{-j})].$
+The result follows directly from the first bound in Lemma E.10 with $L=1$. ∎
+Lemma E.14.
+Assume $\boldsymbol{\theta}=\boldsymbol{0}$. Under Assumption A3, for any $j\in[m]$,
+$\left|\frac{\mathbb{P}(\hat{\theta}_{j}\neq 0)}{2(1-\Phi(\lambda))}-1\right|\leq 24m\,\exp\left\{-\frac{1-\rho_{\infty}}{2(1+\rho_{\infty})}\lambda^{2}\right\}+2m^{2}\exp\left\{-\zeta(\rho_{\infty})\lambda^{2}\right\}.$
+Proof.
+By Lemma E.5,
+$\mathbb{P}(\hat{\theta}_{j}\neq 0)$ $\geq\mathbb{P}(|Z_{j}|>\lambda,\|\boldsymbol{Z}_{-j}\|_{\infty}<\lambda)$
+$=\mathbb{P}(|Z_{j}|>\lambda)-\mathbb{P}(|Z_{j}|>\lambda,\|\boldsymbol{Z}_{-j}\|_{\infty}>\lambda)$
+$=2(1-\Phi(\lambda))-\mathbb{P}(|Z_{j}|>\lambda,\|\boldsymbol{Z}_{-j}\|_{\infty}>\lambda).$
+By the union bound, Lemma E.11, and Proposition E.1, for $\lambda>1$,
+$\mathbb{P}(|Z_{j}|>\lambda,\|\boldsymbol{Z}_{-j}\|_{\infty}>\lambda)\leq\sum_{k\neq j}\mathbb{P}(|Z_{j}|>\lambda,|Z_{k}|>\lambda)$
+$\leq 4m\left(1-\Phi\left(\sqrt{\frac{2}{1+\rho_{\infty}}}\lambda\right)\right)\leq 4m\sqrt{\frac{1+\rho_{\infty}}{2}}\cdot\frac{1}{\lambda}\exp\left\{-\frac{\lambda^{2}}{1+\rho_{\infty}}\right\}$
+$\leq 4m\cdot\frac{1}{\lambda}\exp\left\{-\frac{\lambda^{2}}{2}\right\}\cdot\exp\left\{-\frac{1-\rho_{\infty}}{2(1+\rho_{\infty})}\lambda^{2}\right\}$
+$\leq(1-\Phi(\lambda))\cdot 24m\cdot\exp\left\{-\frac{1-\rho_{\infty}}{2(1+\rho_{\infty})}\lambda^{2}\right\}.$
+Thus
+$\frac{\mathbb{P}(\hat{\theta}_{j}\neq 0)}{2(1-\Phi(\lambda))}-1\geq-12m\exp\left\{-\frac{1-\rho_{\infty}}{2(1+\rho_{\infty})}\lambda^{2}\right\}.$ (21)
+For the upper bound, Lemma E.4 gives
+$\mathbb{P}(\hat{\theta}_{j}\neq 0)=\mathbb{P}\big(|Z_{j}-\Sigma_{j,-j}\hat{\boldsymbol{\theta}}_{-j}|>\lambda\big).$
+If $|Z_{j}|<\lambda$, then $\hat{\boldsymbol{\theta}}_{-j}$ must be nonzero to have $|Z_{j}-\Sigma_{j,-j}\hat{\boldsymbol{\theta}}_{-j}|>\lambda$, so
+$\mathbb{P}(\hat{\theta}_{j}\neq 0)$ $\leq\mathbb{P}(|Z_{j}|>\lambda)+\mathbb{P}(\hat{\theta}_{j}\neq 0,|Z_{j}|<\lambda,\hat{\boldsymbol{\theta}}_{-j}\neq 0)$
+$\leq\mathbb{P}(|Z_{j}|>\lambda)+\mathbb{P}(\|\hat{\boldsymbol{\theta}}\|_{0}\geq 2)$
+$\leq 2(1-\Phi(\lambda))+(1-\Phi(\lambda))\cdot 4m^{2}\exp\big\{-\zeta(\rho_{\infty})\lambda^{2}\big\},$
+where the last step uses Lemma E.12. Thus
+$\frac{\mathbb{P}(\hat{\theta}_{j}\neq 0)}{2(1-\Phi(\lambda))}-1\leq 2m^{2}\exp\big\{-\zeta(\rho_{\infty})\lambda^{2}\big\}.$ (22)
+Combining (21) and (22) yields the claim. ∎
+Lemma E.15.
+Fix $\xi>0$. Assume $|\theta_{j}|>(1+\xi)\lambda$ for some $\lambda>1$ and $\boldsymbol{\theta}_{-j}=\boldsymbol{0}$. Under Assumption A3,
+$1-\mathbb{P}(\hat{\theta}_{j}\neq 0)\leq(1-\Phi(\xi\lambda))+2m(1-\Phi(\lambda)).$
+Proof.
+By Lemma E.5,
+$\mathbb{P}(\hat{\theta}_{j}\neq 0)$ $\geq\mathbb{P}(|Z_{j}|>\lambda,\|\boldsymbol{Z}_{-j}\|_{\infty}<\lambda)$
+$=\mathbb{P}(|Z_{j}|>\lambda)-\mathbb{P}(|Z_{j}|>\lambda,\|\boldsymbol{Z}_{-j}\|_{\infty}>\lambda)$
+$\geq\mathbb{P}(|Z_{j}|>\lambda)-\mathbb{P}(\|\boldsymbol{Z}_{-j}\|_{\infty}>\lambda).$
+By the union bound,
+$\mathbb{P}(\|\boldsymbol{Z}_{-j}\|_{\infty}>\lambda)\leq\sum_{k\neq j}\mathbb{P}(|Z_{k}|>\lambda)\leq 2m(1-\Phi(\lambda)).$
+Since $Z_{j}\sim\mathcal{N}(\theta_{j},1)$ and $|\theta_{j}|>(1+\xi)\lambda$,
+$\mathbb{P}(|Z_{j}|<\lambda)=\Phi(\lambda-\theta_{j})-\Phi(-\lambda-\theta_{j})\leq 1-\Phi(\xi\lambda).$
+Combining these bounds yields
+$1-\mathbb{P}(\hat{\theta}_{j}\neq 0)\leq(1-\Phi(\xi\lambda))+2m(1-\Phi(\lambda)).$
+E.3.3 Bias and variance analysis
+Theorem E.3.
+Let $W_{j}=\phi_{j}(p_{j})\,\pi_{j}(\boldsymbol{Z}_{-j})$ and $W=\sum_{j=1}^{m}W_{j}$.
+Further let $R=\sum_{j=1}^{m}\mathbf{1}\{\hat{\theta}_{j}\neq 0\}$.
+Assume $\lambda=\sqrt{c\log d}$, where $d$ is sufficiently large so that $\lambda>1$, and $m=O(d^{\omega})$ with
+$0\leq\omega<c\cdot\min\left\{\frac{1-\rho_{\infty}}{2(1+\rho_{\infty})},\frac{\zeta(\rho_{\infty})}{3},\frac{(1-\rho_{1})^{2}}{4\,\tau_{\max}^{2}},\frac{\tau_{\min}^{2}}{4(2-\tau_{\min}^{2})}\right\}.$
+Then:
+  • For a null block in $\mathcal{B}_{0}$,
+    $\mathbb{E}[W]=\mathbb{E}[R]=2m(1-\Phi(\lambda))(1+o(1)),$
+    and
+    $\textnormal{Var}(W)=o(\mathbb{E}[W]),\qquad\textnormal{Var}(R)=2m(1-\Phi(\lambda))(1+o(1)).$
+  • For a “good” non-null block in $\mathcal{B}_{1g}$,
+    $\mathbb{E}[W]=o(1),\quad\mathbb{E}[R]=1+o(1),$
+    and
+    $\textnormal{Var}(W)=o(1),\quad\textnormal{Var}(R)=o(1).$
+All $o(1)$ terms above are uniform in $\tau_{\min},\tau_{\max},\rho_{1},\rho_{\infty},c,$ and $\omega$.
+Proof.
+First note that for any $j$ with $\theta_{j}=0$, $p_{j}\sim\mathrm{Unif}([0,1])$ and $p_{j}$ is independent of $\boldsymbol{Z}_{-j}$. Thus
+$\mathbb{E}[W_{j}]=\mathbb{E}[\phi_{j}(p_{j})]\mathbb{E}[\pi_{j}(\boldsymbol{Z}_{-j})]=\mathbb{E}[\pi_{j}(\boldsymbol{Z}_{-j})]=\mathbb{P}(\hat{\theta}_{j}\neq 0),$ (23)
+where the last equality uses the iterated law of expectation and the definition of $\pi_{j}$. Also,
+$\mathbb{E}[W_{j}^{2}]=\mathbb{E}[\phi_{j}^{2}(p_{j})]\mathbb{E}[\pi^{2}_{j}(\boldsymbol{Z}_{-j})]=\frac{1}{1-\zeta}\,\mathbb{E}[\pi^{2}_{j}(\boldsymbol{Z}_{-j})],$ (24)
+since $\phi_{j}(p_{j})^{2}=\mathbf{1}\{p_{j}>\zeta\}/(1-\zeta)^{2}$ and $p_{j}\sim\mathrm{Unif}[0,1]$.
+Case 1. Null block. For a null block, Lemma E.14 and (23) give
+$\left|\frac{\mathbb{E}[W]}{2m(1-\Phi(\lambda))}-1\right|$ $=\left|\frac{\mathbb{E}[R]}{2m(1-\Phi(\lambda))}-1\right|$
+$\leq 24m\exp\left\{-\frac{1-\rho_{\infty}}{2(1+\rho_{\infty})}\lambda^{2}\right\}+2m^{2}\exp\left\{-\zeta(\rho_{\infty})\lambda^{2}\right\}.$
+Using $\lambda^{2}=c\log d$ and $m=O(d^{\omega})$, each term is of order $d^{\omega-c(1-\rho_{\infty})/2(1+\rho_{\infty})}$ and $d^{2\omega-c\zeta(\rho_{\infty})}$, which both tend to 0 under the condition on $\omega$.
+This yields
+$\mathbb{E}[W]=\mathbb{E}[R]=2m(1-\Phi(\lambda))(1+o(1)).$
+For the variance of $W$, by (24),
+$\textnormal{Var}(W)\leq m\sum_{j=1}^{m}\textnormal{Var}(W_{j})\leq m\sum_{j=1}^{m}\mathbb{E}(W_{j}^{2})\leq\frac{m}{1-\zeta}\sum_{j=1}^{m}\mathbb{E}[\pi_{j}^{2}(\boldsymbol{Z}_{-j})].$
+Lemma E.10 gives
+$\frac{\textnormal{Var}(W)}{2m(1-\Phi(\lambda))}$ $\preceq\frac{m^{2}}{\lambda}\exp\left\{-\left(\frac{(1-\rho_{1})^{2}}{\tau_{\max}^{2}}\wedge\frac{\tau_{\min}^{2}}{2(2-\tau_{\min}^{2})}\right)\lambda^{2}\right\}$
+$\preceq\exp\left\{-\left[c\left(\frac{(1-\rho_{1})^{2}}{\tau_{\max}^{2}}\wedge\frac{\tau_{\min}^{2}}{2(2-\tau_{\min}^{2})}\right)-2\omega\right]\log d\right\}=o(1),$
+again using the constraint on $\omega$. Thus $\textnormal{Var}(W)=o(\mathbb{E}[W])$.
+For $R$, we use
+$\textnormal{Var}(R)\leq\mathbb{E}[R^{2}]=\sum_{j=1}^{m}\mathbb{P}(\hat{\theta}_{j}\neq 0)+\sum_{j\neq k}\mathbb{P}(\hat{\theta}_{j}\neq 0,\hat{\theta}_{k}\neq 0)=\mathbb{E}[R]+\sum_{j\neq k}\mathbb{P}(\hat{\theta}_{j}\neq 0,\hat{\theta}_{k}\neq 0).$
+Lemma E.12 gives
+$\frac{\sum_{j\neq k}\mathbb{P}(\hat{\theta}_{j}\neq 0,\hat{\theta}_{k}\neq 0)}{2m(1-\Phi(\lambda))}\preceq m^{3}\exp\{-\zeta(\rho_{\infty})\lambda^{2}\}=d^{\,3\omega-c\zeta(\rho_{\infty})}=o(1),$
+again under the condition on $\omega$. Combining this with the expression for $\mathbb{E}[R]$ yields $\textnormal{Var}(R)=2m(1-\Phi(\lambda))(1+o(1))$.
+Case 2. Good non-null block. Assume, without loss of generality, that $\theta_{1}\neq 0$ with $|\theta_{1}|>(1+\xi)\lambda$ and $\theta_{j}=0$ for all $j>1$.
+For $j\geq 2$, Lemma E.13 and (23) give
+$\sum_{j=2}^{m}\mathbb{E}[W_{j}]=\sum_{j=2}^{m}\mathbb{P}(\hat{\theta}_{j}\neq 0)$ $\preceq\frac{m}{\lambda}\exp\left\{-\frac{(1-\rho_{1})^{2}}{2\,\tau_{\max}^{2}}\lambda^{2}\right\}$
+$\preceq\exp\left\{\left(c\frac{(1-\rho_{1})^{2}}{2\tau_{\max}^{2}}-\omega\right)\log d\right\}=o(1).$ (25)
+For $j=1$,
+$\mathbb{E}[W_{1}]\;\leq\;\mathbb{E}[\phi_{j}(p_{j})]\;\preceq\;\mathbb{P}(p_{1}>\zeta)\;=\;o(1)$
+by Mill’s bound. Furthermore, by Lemma E.15,
+$\big|1-\mathbb{P}(\hat{\theta}_{1}\neq 0)\big|\leq(1-\Phi(\xi\lambda))+2m(1-\Phi(\lambda))=o(1).$ (26)
+Thus
+$\mathbb{E}[W]=o(1),\qquad|\mathbb{E}[R]-1|=o(1).$
+For the variances, Lemma E.10 yields
+$\textnormal{Var}(W)$ $\leq\mathbb{E}[W^{2}]\leq m\sum_{j=1}^{m}\mathbb{E}[\pi_{j}^{2}(\boldsymbol{Z}_{-j})]$
+$\preceq\frac{m^{2}}{\lambda^{2}}\exp\left\{-\frac{2(1-\rho_{1})^{2}}{2\,\tau_{\max}^{2}}\lambda^{2}\right\}\preceq\exp\left\{2\left(c\frac{(1-\rho_{1})^{2}}{2\tau_{\max}^{2}}-\omega\right)\log d\right\}=o(1),$
+under the stated constraint on $\omega$.
+For $R$, write
+$\textnormal{Var}(R)\leq 2\textnormal{Var}(\mathbf{1}\{\hat{\theta}_{1}\neq 0\})+2\textnormal{Var}\Big(\sum_{j=2}^{m}\mathbf{1}\{\hat{\theta}_{j}\neq 0\}\Big).$
+By the same reasoning as in (25), one has
+$\textnormal{Var}\Big(\sum_{j=2}^{m}\mathbf{1}\{\hat{\theta}_{j}\neq 0\}\Big)=o(1),$
+and by (26), $\textnormal{Var}(\mathbf{1}\{\hat{\theta}_{1}\neq 0\})=o(1)$ as well. Hence $\textnormal{Var}(R)=o(1)$.
+This completes the proof. ∎
+E.4 Proof of Theorems E.1 and E.2
+Proof of Theorem E.1.
+We reintroduce block indices and write
+$R=\sum_{b=1}^{B}R_{(b)},\qquad W=\sum_{b=1}^{B}W_{(b)},$
+where $R_{(b)}$ and $W_{(b)}$ are, respectively, the blockwise model size and blockwise contribution to $W$ for block $b$.
+In this section, all $o(1)$, $o_{\mathbb{P}}(1)$, $O(1)$, and $O_{\mathbb{P}}(1)$ terms are uniform given the constants in Assumptions A1-A5.
+Step 1: Mean and variance of $R$.
+By Theorem E.3, for each block $b$ we have
+$\mathbb{E}[R_{(b)}]=\begin{cases}\displaystyle\frac{v_{d}\,m_{b}}{d}\,(1+o(1)),&b\in\mathcal{B}_{0},\\[6.99997pt] 1+o(1),&b\in\mathcal{B}_{1g},\\[2.5pt] O(1),&b\in\mathcal{B}_{1b},\end{cases}\qquad\textnormal{Var}(R_{(b)})=\begin{cases}\displaystyle\frac{v_{d}\,m_{b}}{d}\,(1+o(1)),&b\in\mathcal{B}_{0},\\[6.99997pt] o(1),&b\in\mathcal{B}_{1g},\\[2.5pt] O(1),&b\in\mathcal{B}_{1b}.\end{cases}$
+Here the $o(1)$ terms are uniform over blocks.
+Since $R=\sum_{b=1}^{B}R_{(b)}$, we obtain
+$\mathbb{E}[R]$ $=\sum_{b\in\mathcal{B}_{0}}\frac{v_{d}m_{b}}{d}\,(1+o(1))+\sum_{b\in\mathcal{B}_{1g}}(1+o(1))+O(B_{1b}).$
+Note that
+$\sum_{b\in\mathcal{B}_{0}}m_{b}=d-\sum_{b\in\mathcal{B}_{1g}\cup\mathcal{B}_{1b}}m_{b}.$
+By bounded block size and Assumption A4,
+$\Bigg|\sum_{b\in\mathcal{B}_{0}}m_{b}-d\Bigg|\;\leq\;m_{\max}(B_{1g}+B_{1b})\;=\;o(d),$
+so
+$\sum_{b\in\mathcal{B}_{0}}m_{b}=d\,(1+o(1)).$
+Furthermore, $B_{1b}=o(B_{1g})$ by Assumption A4, and $B_{1g}=\Theta(v_{d})$ by Assumption A5. Hence
+$\mathbb{E}[R]=v_{d}\,(1+o(1))+B_{1g}\,(1+o(1))=(v_{d}+B_{1g})(1+o(1)).$ (27)
+By block orthogonality (Assumption A1), $R_{(b)}$ are independent across $b$, so
+$\textnormal{Var}(R)=\sum_{b=1}^{B}\textnormal{Var}(R_{(b)}).$
+From the blockwise bounds above and the same counting argument as for the mean,
+$\textnormal{Var}(R)=v_{d}\,(1+o(1))+O(B_{1g}+B_{1b})=(v_{d}+B_{1g})(1+o(1)).$ (28)
+Assumption A5 states that $v_{d}$ is of the same order as $B_{1g}$, so
+$\textnormal{Var}(R)=\Theta(B_{1g}),\qquad v_{d}=\Theta(B_{1g}).$
+Step 2: CLT for $R$.
+By the Berry–Esseen theorem for sums of independent variables,
+$d_{K}\!\left(\frac{R-\mathbb{E}[R]}{\sqrt{\textnormal{Var}(R)}},\ \mathcal{N}(0,1)\right)\;\leq\;\frac{\sum_{b=1}^{B}\mathbb{E}|R_{(b)}-\mathbb{E}[R_{(b)}]|^{3}}{\textnormal{Var}(R)^{3/2}}.$
+Since $0\leq R_{(b)}\leq m_{b}\leq m_{\max}$,
+$|R_{(b)}-\mathbb{E}[R_{(b)}]|^{3}\;\leq\;m_{\max}\,(R_{(b)}-\mathbb{E}[R_{(b)}])^{2},$
+so
+$\sum_{b=1}^{B}\mathbb{E}|R_{(b)}-\mathbb{E}[R_{(b)}]|^{3}\;\leq\;m_{\max}\sum_{b=1}^{B}\textnormal{Var}(R_{(b)})=m_{\max}\textnormal{Var}(R).$
+Hence
+$d_{K}\!\left(\frac{R-\mathbb{E}[R]}{\sqrt{\textnormal{Var}(R)}},\ \mathcal{N}(0,1)\right)\;\leq\;\frac{m_{\max}\textnormal{Var}(R)}{\textnormal{Var}(R)^{3/2}}=\frac{m_{\max}}{\sqrt{\textnormal{Var}(R)}}=O\!\left(\frac{1}{\sqrt{B_{1g}}}\right)=o(1).$ (29)
+where the last step uses Assumption A4 which implies $B_{1g}\rightarrow\infty$.
+Step 3: Mean and variance of $W$.
+By Theorem E.3, for each block $b$,
+$\mathbb{E}[W_{(b)}]=\begin{cases}\displaystyle\frac{v_{d}m_{b}}{d}\,(1+o(1)),&b\in\mathcal{B}_{0},\\[6.99997pt] o(1),&b\in\mathcal{B}_{1g},\\[2.5pt] O(1),&b\in\mathcal{B}_{1b},\end{cases}\qquad\textnormal{Var}(W_{(b)})=\begin{cases}\displaystyle\frac{v_{d}m_{b}}{d}\cdot o(1),&b\in\mathcal{B}_{0},\\[6.99997pt] o(1),&b\in\mathcal{B}_{1g},\\[2.5pt] O(1),&b\in\mathcal{B}_{1b}.\end{cases}$
+Summing over blocks and using the same counting argument as for $R$,
+$\mathbb{E}[W]=v_{d}(1+o(1)),\qquad\textnormal{Var}(W)=o(v_{d})=o(B_{1g}),$ (30)
+where the last equality follows from $v_{d}=\Theta(B_{1g})$.
+Step 4: Approximating $\widehat{\textnormal{FDR}}$ by $W/R$.
+Recall $\boldsymbol{Z}_{-(b)}=\left\{\boldsymbol{Z}_{(b^{\prime})}\right\}_{b^{\prime}\neq b}$ is the subfficient statistic except for block $b$. Write
+$\widehat{\textnormal{FDR}}=\sum_{b=1}^{B}\sum_{j\in[m_{b}]}\widehat{\textnormal{FDR}}_{(b),j},$
+where for $j\in[m_{b}]$,
+$\widehat{\textnormal{FDR}}_{(b),j}$ $=\phi_{(b),j}(p_{(b),j})\cdot\mathbb{E}_{\theta_{(b),j}=0}\!\left[\frac{\mathbf{1}\{\hat{\theta}_{(b),j}\neq 0\}}{R}\;\middle|\;\boldsymbol{Z}_{(b),-j},\,\boldsymbol{Z}_{-(b)}\right]$
+$=\phi_{(b),j}(p_{(b),j})\cdot\mathbb{P}_{\theta_{(b),j}=0}\!\left(\hat{\theta}_{(b),j}\neq 0\mid\boldsymbol{Z}_{(b),-j}\right)\cdot\mathbb{E}_{\theta_{(b),j}=0}\!\left[\frac{1}{R}\;\middle|\;\hat{\theta}_{(b),j}\neq 0,\ \boldsymbol{Z}_{(b),-j},\,\boldsymbol{Z}_{-(b)}\right]$
+$=\frac{\phi_{(b),j}(p_{(b),j})\,\pi_{(b),j}(\boldsymbol{Z}_{(b),-j})}{R+U_{(b),j}}$
+$=\frac{W_{(b),j}}{R+U_{(b),j}},$
+where $|U_{(b),j}|\leq m_{b}\leq m_{\max}$; the perturbation $U_{(b),j}$ accounts for the fact that perturbating $Z_{(b),j}$ changes $R$ by at most $m_{b}$.
+Define
+$\widetilde{\mathrm{FDR}}=\frac{W}{R}.$
+Then
+$\widehat{\textnormal{FDR}}-\widetilde{\mathrm{FDR}}$ $=\sum_{b=1}^{B}\sum_{j\in[m_{b}]}\left(\frac{W_{(b),j}}{R+U_{(b),j}}-\frac{W_{(b),j}}{R}\right)$
+$=-\sum_{b=1}^{B}\sum_{j\in[m_{b}]}\frac{W_{(b),j}U_{(b),j}}{R(R+U_{(b),j})},$
+so, on the event $\{R>2m_{\max}\}$,
+$\bigl|\widehat{\textnormal{FDR}}-\widetilde{\mathrm{FDR}}\bigr|\leq\sum_{b,j}\frac{2m_{\max}W_{(b),j}}{R^{2}}=\frac{2m_{\max}W}{R^{2}}.$
+Since $\mathbb{E}[R]$, $\textnormal{Var}(R)$, and $\mathbb{E}[W]$ are $\Theta(v_{d})$, by Markov inequality,
+$\bigl|\widehat{\textnormal{FDR}}-\widetilde{\mathrm{FDR}}\bigr|=O_{\mathbb{P}}\!\left(\frac{1}{v_{d}}\right).$ (31)
+Step 5: Centering and linearization.
+Let
+$\tilde{\alpha}_{d}=\frac{\mathbb{E}[W]}{\mathbb{E}[R]}.$
+By (27) and (30),
+$\tilde{\alpha}_{d}=\frac{v_{d}(1+o(1))}{v_{d}+B_{1g}}(1+o(1))=\alpha_{d}(1+o(1)).$
+We now expand
+$\widetilde{\text{FDR}}-\tilde{\alpha}_{d}=\frac{W}{R}-\frac{\mathbb{E}[W]}{\mathbb{E}[R]}=\frac{W-\mathbb{E}[W]}{R}-\frac{(R-\mathbb{E}[R])\,\mathbb{E}[W]}{R\,\mathbb{E}[R]}.$
+For the first term, by Chebyshev’s inequality, (30), and the fact that $R$ is of order $v_{d}$,
+$\frac{W-\mathbb{E}[W]}{R}=O_{\mathbb{P}}\!\left(\frac{\sqrt{\textnormal{Var}(W)}}{R}\right)=o_{\mathbb{P}}\!\left(\frac{1}{\sqrt{v_{d}}}\right).$
+For the second term, by (27), (28), and the Markov’s inequality,
+$\frac{\mathbb{E}[R]}{R}=1-\frac{R-\mathbb{E}[R]}{R}=1+o_{\mathbb{P}}(1).$
+Thus
+$\frac{1}{R}=\frac{1}{\mathbb{E}[R]}(1+o_{\mathbb{P}}(1))$
+and
+$\frac{(R-\mathbb{E}[R])\,\mathbb{E}[W]}{R\,\mathbb{E}[R]}=\tilde{\alpha}_{d}\;\frac{R-\mathbb{E}[R]}{R}=\tilde{\alpha}_{d}\;\frac{R-\mathbb{E}[R]}{\mathbb{E}[R]}\,(1+o_{\mathbb{P}}(1)).$
+Putting this together,
+$\sqrt{v_{d}}\,(\widetilde{\text{FDR}}-\tilde{\alpha}_{d})=-\tilde{\alpha}_{d}\,\sqrt{v_{d}}\,\frac{R-\mathbb{E}[R]}{\mathbb{E}[R]}+o_{\mathbb{P}}(1).$
+Multiplying and dividing by $\sqrt{\textnormal{Var}(R)}$ and using (27)–(28), we obtain
+$\sqrt{v_{d}}\,(\widetilde{\text{FDR}}-\tilde{\alpha}_{d})$ $=-\tilde{\alpha}_{d}\,\frac{\sqrt{v_{d}}\,\sqrt{\textnormal{Var}(R)}}{\mathbb{E}[R]}\cdot\frac{R-\mathbb{E}[R]}{\sqrt{\textnormal{Var}(R)}}+o_{\mathbb{P}}(1)$
+$=-\tilde{\alpha}_{d}\,\sqrt{\alpha_{d}}\,\frac{R-\mathbb{E}[R]}{\sqrt{\textnormal{Var}(R)}}+o_{\mathbb{P}}(1),$
+where the second equation uses
+$\frac{\sqrt{v_{d}}\,\sqrt{\textnormal{Var}(R)}}{\mathbb{E}[R]}=\frac{\sqrt{v_{d}}\,\sqrt{v_{d}+B_{1g}}}{v_{d}+B_{1g}}(1+o(1))=\sqrt{\frac{v_{d}}{v_{d}+B_{1g}}}(1+o(1))=\sqrt{\alpha_{d}}\,(1+o(1)).$
+By Assumption A5,
+$\alpha_{d}\in\left[\chi,1-\chi\right]=\Theta(1),\quad\tilde{\alpha}_{d}=\Theta(1).$ (32)
+Therefore,
+$\sqrt{\frac{v_{d}}{\alpha_{d}^{3}}}\;(\widetilde{\text{FDR}}-\tilde{\alpha}_{d})=-\frac{\tilde{\alpha}_{d}}{\alpha_{d}}\;\frac{R-\mathbb{E}[R]}{\sqrt{\textnormal{Var}(R)}}+o_{\mathbb{P}}(1).$
+Recalling that all $o_{\mathbb{P}}(1)$ terms are uniform and $\Phi$ is Lipschitz, (29) implies that
+$d_{K}\left(\sqrt{\frac{v_{d}}{\alpha_{d}^{3}}}\;(\widetilde{\text{FDR}}-\tilde{\alpha}_{d}),\mathcal{N}(0,1)\right)=o(1).$
+Finally, by (31),
+$\sqrt{\frac{v_{d}}{\alpha_{d}^{3}}}\;(\widehat{\textnormal{FDR}}-\tilde{\alpha}_{d})=\sqrt{\frac{v_{d}}{\alpha_{d}^{3}}}\;(\widetilde{\text{FDR}}-\tilde{\alpha}_{d})+o_{\mathbb{P}}\left(\frac{1}{\sqrt{v_{d}}}\right),$
+so the same limit holds:
+$d_{K}\left(\sqrt{\frac{v_{d}}{\alpha_{d}^{3}}}\;(\widehat{\textnormal{FDR}}-\tilde{\alpha}_{d}),\mathcal{N}(0,1)\right)=o(1).$
+Proof of Theorem E.2.
+We work throughout under the assumptions of Theorem E.1, with the only change being that the Lasso penalty defining the OLS-after-Lasso model is $\bar{\lambda}=\bar{c}\lambda$ with $\bar{c}>1$.
+We will show that, with high probability, the estimated OLS-after-Lasso model satisfies the same structural assumptions (A1–A5) with $\hat{\boldsymbol{\theta}}^{(*)}$ (the OLS-after-Lasso coefficients) playing the role of the true coefficients.
+Conditional on such a “good” data event, we can then apply Theorem E.1 to the bootstrap procedure.
+Step 1: Null blocks contribute negligibly under the larger penalty.
+By Theorem E.3, applied with penalty level $\bar{\lambda}$, for any null block $b\in\mathcal{B}_{0}$ we have
+$\mathbb{E}[R_{(b)}]=2m_{b}\bigl(1-\Phi(\bar{\lambda})\bigr)\,(1+o(1)).$
+For sufficiently large $d$ we have $\bar{\lambda}=\bar{c}\lambda>\lambda>1$. By the upper Mill’s ratio bound (Proposition E.1),
+$1-\Phi(\bar{\lambda})\;\leq\;\frac{1}{\bar{\lambda}}\exp\!\left(-\frac{\bar{\lambda}^{2}}{2}\right)=\frac{1}{\bar{c}\lambda}\exp\!\left(-\frac{\bar{c}^{2}\lambda^{2}}{2}\right).$
+Write
+$1-\Phi(\lambda)=\frac{v_{d}}{2d},$
+as in Assumption A5. Using again the Mill’s ratio approximation, we have
+$\lambda^{2}=2\log(d/v_{d})+O(\log(\lambda)).$
+Therefore,
+$\exp\!\left(-\frac{\bar{c}^{2}\lambda^{2}}{2}\right)=\exp\!\bigl(-\bar{c}^{2}\log(d/v_{d})\bigr)\cdot O(\lambda)=(d/v_{d})^{-\bar{c}^{2}}\cdot O(\lambda).$
+Therefore,
+$1-\Phi(\bar{\lambda})=O\!\left(\Bigl(\frac{v_{d}}{d}\Bigr)^{\bar{c}^{2}}\right).$
+Summing over all null blocks and using $\sum_{b\in\mathcal{B}_{0}}m_{b}\asymp d$, we obtain
+$\mathbb{E}\Biggl[\;\sum_{b\in\mathcal{B}_{0}}R_{(b)}\Biggr]=O\!\left(d\cdot\Bigl(\frac{v_{d}}{d}\Bigr)^{\bar{c}^{2}}\right)=O\!\left(\frac{v_{d}^{\bar{c}^{2}}}{d^{\bar{c}^{2}-1}}\right)=o(v_{d}),$
+since $v_{d}/d\to 0$ and $\bar{c}^{2}>1$.
+Step 2: Good blocks are recovered (support-wise) with high probability.
+Without loss of generality, for each good non-null block $b\in\mathcal{B}_{1g}$ we relabel coordinates so that
+$\theta_{(b),1}\neq 0,\qquad\theta_{(b),-1}=\boldsymbol{0}.$
+Note that
+$|\theta_{j}|>(1+\xi)\lambda=\frac{1+\xi}{\bar{c}}\bar{\lambda}>\bar{\lambda}.$
+Apply the single-block analysis (Theorem E.3 and its lemmas) with penalty $\bar{\lambda}$; the same structure carries through since the within-block regularity assumption A3 is unchanged and $\bar{\lambda}\to\infty$ as $d\to\infty$.
+In particular, by equations (25) and (26) (applied with $\lambda$ replaced by $\bar{\lambda}$), we have, for each such good block,
+$\mathbb{P}\bigl(\hat{\theta}_{(b),-1}^{\bar{\lambda}}\neq 0\bigr)=o(1),\qquad\mathbb{P}\bigl(\hat{\theta}_{(b),1}^{\bar{\lambda}}=0\bigr)=o(1),$
+where $\hat{\boldsymbol{\theta}}^{\bar{\lambda}}$ is the Lasso estimator at level $\bar{\lambda}$. Hence
+$\mathbb{E}\Biggl[\sum_{b\in\mathcal{B}_{1g}}\mathbf{1}\Bigl(\hat{\theta}_{(b),-1}^{\bar{\lambda}}\neq 0\ \text{or}\ \hat{\theta}_{(b),1}^{\bar{\lambda}}=0\Bigr)\Biggr]=o(B_{1g}).$
+Combining this with the bound from Step 1 and using Markov’s inequality, there exists a deterministic sequence $\iota_{d}\to 0$ and an event $\mathcal{E}_{d}$ with
+$\mathbb{P}(\mathcal{E}_{d})=1-o(1),$
+such that on $\mathcal{E}_{d}$,
+$\sum_{b\in\mathcal{B}_{0}}R_{(b)}^{\bar{\lambda}}\leq\iota_{d}v_{d},\qquad\sum_{b\in\mathcal{B}_{1g}}\mathbf{1}\Bigl(\hat{\theta}_{(b),-1}^{\bar{\lambda}}\neq 0\ \text{or}\ \hat{\theta}_{(b),1}^{\bar{\lambda}}=0\Bigr)\leq\iota_{d}B_{1g},$
+where $R_{(b)}^{\bar{\lambda}}$ is the blockwise model size under Lasso with penalty $\bar{\lambda}$.
+Define the set of “error blocks” as
+$\mathcal{B}_{\mathrm{err}}=\bigl\{b\in\mathcal{B}_{0}:\ R_{(b)}^{\bar{\lambda}}>0\bigr\}\;\cup\;\Bigl\{b\in\mathcal{B}_{1g}:\ \hat{\theta}_{(b),-1}^{\bar{\lambda}}\neq 0\ \text{or}\ \hat{\theta}_{(b),1}^{\bar{\lambda}}=0\Bigr\}.$
+Then on $\mathcal{E}_{d}$,
+$|\mathcal{B}_{\mathrm{err}}|\;\leq\;\iota_{d}(v_{d}+B_{1g})\;=\;o(B_{1g}),$
+using Assumption A5 to relate $v_{d}$ and $B_{1g}$.
+We now define a perturbed block partition:
+$\tilde{\mathcal{B}}_{0}=\mathcal{B}_{0}\setminus\mathcal{B}_{\mathrm{err}},\qquad\tilde{\mathcal{B}}_{1g}=\mathcal{B}_{1g}\setminus\mathcal{B}_{\mathrm{err}},\qquad\tilde{\mathcal{B}}_{1b}=\mathcal{B}_{1b}\cup\mathcal{B}_{\mathrm{err}}.$
+On $\mathcal{E}_{d}$, we then have
+$|\tilde{\mathcal{B}}_{0}|=B_{0}+o(B_{1g})=B_{0}(1+o(1)),$
+$|\tilde{\mathcal{B}}_{1g}|=B_{1g}+o(B_{1g})=B_{1g}(1+o(1)),$
+$|\tilde{\mathcal{B}}_{1b}|=B_{1b}+o(B_{1g})=o(B_{1g}),$
+so the sparsity/purity structure of Assumption A4 is preserved up to $o(B_{1g})$ blocks.
+On $\mathcal{E}_{d}$, the OLS-after-Lasso estimator $\hat{\boldsymbol{\theta}}^{(*)}$ has the same support as $\hat{\boldsymbol{\theta}}^{\bar{\lambda}}$ across the null and good non-null blocks.
+Step 3: Lower bound on the estimated signal in good blocks.
+For any good block $b\in\tilde{\mathcal{B}}_{1g}$, on $\mathcal{E}_{d}$, the Lasso at level $\bar{\lambda}$ selects only coordinate 1 in block $b$.
+The OLS-after-Lasso coefficient on that coordinate is simply the least squares coefficient in a univariate regression, and hence
+$\hat{\theta}_{(b),1}^{(*)}=Z_{(b),1},$
+where $Z_{(b),1}=\boldsymbol{X}_{(b),1}^{\top}\boldsymbol{Y}$.
+Fix any $\iota>0$. Using a union bound and the Gaussian tail bound, we have
+$\mathbb{P}\Biggl(\max_{b\in\mathcal{B}_{1g}}\bigl|\hat{\theta}_{(b),1}^{(*)}-\theta_{(b),1}\bigr|>\sqrt{2(1+\iota)\log B_{1g}}\Biggr)$ $\leq B_{1g}\exp\bigl(-(1+\iota)\log B_{1g}\bigr)$
+$=B_{1g}^{-\iota}\;=\;o(1),$
+since $B_{1g}\to\infty$. Thus there exists an event $\tilde{\mathcal{E}}_{d}$ with
+$\mathbb{P}(\tilde{\mathcal{E}}_{d})=1-o(1),$
+such that on $\tilde{\mathcal{E}}_{d}$,
+$\max_{b\in\mathcal{B}_{1g}}\bigl|\hat{\theta}_{(b),1}^{(*)}-\theta_{(b),1}\bigr|\;\leq\;\sqrt{2(1+\iota)\log B_{1g}}.$
+Therefore, on $\tilde{\mathcal{E}}_{d}$,
+$\min_{b\in\mathcal{B}_{1g}}|\hat{\theta}_{(b),1}^{(*)}|\;\geq\;(1+\xi)\lambda-\sqrt{2(1+\iota)\log B_{1g}}.$
+We now relate $\lambda$ and $\log B_{1g}$. By the Mill’s ratio expansion, we have
+$\lambda^{2}=2\log\Bigl(\frac{d}{v_{d}}\Bigr)\,(1+o(1)).$
+By Assumption A5, $v_{d}=\Theta(B_{1g})$. For sufficiently large $d$, it follows that
+$\lambda^{2}\in\left[(2-2\iota)\log\Bigl(\frac{d}{B_{1g}}\Bigr),\ (2+2\iota)\log\Bigl(\frac{d}{B_{1g}}\Bigr)\right].$
+Moreover, recall
+$B_{1g}=O\!\bigl(d^{\nu}\Gamma_{d}\bigr),\qquad\Gamma_{d}=o(d^{a})\ \text{for all }a>0,$
+we write
+$\log B_{1g}=\nu\log d+\log\Gamma_{d}+o(\log d),\qquad\log({d}/{B_{1g}})=(1-\nu)\log d-\log\Gamma_{d}+o(\log d).$
+Therefore
+$\quad\frac{\min_{b\in\mathcal{B}_{1g}}|\hat{\theta}_{(b),1}^{(*)}|}{\bar{\lambda}}$
+$>\frac{(1+\xi)\sqrt{2-2\iota}\cdot\sqrt{\log(d/B_{1g})}-\sqrt{2(1+\iota)\log B_{1g}}}{\bar{c}\sqrt{2+2\iota}\sqrt{\log(d/B_{1g})}}$
+$=\frac{(1+\xi)\sqrt{1-\iota}\cdot\sqrt{(1-\nu)\log d-\log\Gamma_{d}+o(\log d)}-\sqrt{(1+\iota)(\nu\log d+\log\Gamma_{d}+o(\log d))}}{\bar{c}\sqrt{1+\iota}\sqrt{(1-\nu)\log d-\log\Gamma_{d}+o(\log d)}}.$
+Since $\log\Gamma_{d}=o(\log d)$, for a sufficiently large $d$,
+$\frac{\min_{b\in\mathcal{B}_{1g}}|\hat{\theta}_{(b),1}^{(*)}|}{\bar{\lambda}}$ $>\;\frac{(1+\xi)\sqrt{1-\nu}\,\sqrt{1-\iota}-\sqrt{\nu}\,(1+\iota)}{\bar{c}\,\sqrt{1+\iota}\,\sqrt{1-\nu}}.$
+We now choose $\iota>0$ small enough so that
+$1+\xi>\frac{1+\iota}{\sqrt{1-\iota}}\left(\sqrt{\frac{\nu}{1-\nu}}+\bar{c}\right),$
+which is possible precisely because
+$\xi>\sqrt{\frac{\nu}{1-\nu}}+\bar{c}-1$
+by assumption. For such an $\iota$, the right-hand side of the previous display exceeds $\sqrt{1+\iota}>1$, and therefore, on $\tilde{\mathcal{E}}_{d}$,
+$\frac{\min_{b\in\mathcal{B}_{1g}}|\hat{\theta}_{(b),1}^{(*)}|}{\bar{\lambda}}>1.$
+Combining with the support-error control from Step 2, we obtain that on $\mathcal{E}_{d}\cap\tilde{\mathcal{E}}_{d}$ (an event with probability 1o(1)1-o(1)), every good block in $\tilde{\mathcal{B}}_{1g}$ has a single non-zero coordinate whose magnitude exceeds $\bar{\lambda}$, and all other coordinates in those blocks are zero.
+Moreover, the total number of “bad” blocks $\tilde{\mathcal{B}}_{1b}$ remains $o(B_{1g})$.
+Step 4: Applying the CLT in the bootstrap world.
+On the event $\mathcal{E}_{d}\cap\tilde{\mathcal{E}}_{d}$, the OLS-after-Lasso model with coefficients $\hat{\boldsymbol{\theta}}^{(*)}$ and block partition $(\tilde{\mathcal{B}}_{0},\tilde{\mathcal{B}}_{1g},\tilde{\mathcal{B}}_{1b})$ satisfies Assumptions A1–A5 (with unchanged constants $\tau_{\min},\tau_{\max},\rho_{1},\rho_{\infty}$, and with the same growth rates for $B_{1g}$ and the number of bad blocks).
+Therefore, if we now treat $\hat{\boldsymbol{\theta}}^{(*)}$ as the “true” parameter vector in a parametric bootstrap model, Theorem E.1 applies to the bootstrap estimator $\widehat{\textnormal{FDR}}^{(*)}$ with some deterministic centering
+$\alpha_{d}^{(*)}(\boldsymbol{Z},\boldsymbol{X})\in(0,1),$
+which depends on the data only through $\hat{\boldsymbol{\theta}}^{(*)}$ and the design.
+Concretely, Theorem E.1, together with the uniformity of $o(1)$ terms mentioned at the beginning of the section, implies that there exists $u_{d}=o(1)$ such that
+$d_{K}\left(\mathcal{L}\left(\sqrt{\frac{v_{d}}{\alpha_{d}^{3}}}\;\bigl(\widehat{\textnormal{FDR}}^{(*)}-\alpha_{d}^{(*)}\bigr)\mid\boldsymbol{X},\boldsymbol{Y}\right),\mathcal{N}(0,1)\right)\leq u_{d},\quad\text{if }(\boldsymbol{X},\boldsymbol{Y})\in\mathcal{E}_{d}\cap\tilde{\mathcal{E}}_{d},$
+where $\mathcal{L}(\cdot\mid\cdot)$ denote the conditional law. The proof is then completed by noting that $\mathbb{P}(\mathcal{E}_{d}\cap\tilde{\mathcal{E}}_{d})=1-o(1)$. ∎
+Proof of Lemma E.1.
+A standard computation shows that
+$\Sigma_{(b);-j,-j}^{-1}=\frac{1}{1-\rho_{b}}\left(\boldsymbol{I}_{m_{b}-1}-\frac{\rho_{b}}{(m_{b}-2)\rho_{b}+1}\,\boldsymbol{1}_{m_{b}-1}\boldsymbol{1}_{m_{b}-1}^{\mkern-1.5mu\mathsf{T}}\right).$
+Therefore
+$1-\Sigma_{(b);j,-j}\Sigma_{(b);-j,-j}^{-1}\Sigma_{(b);-j,j}=1-\frac{(m_{b}-1)\rho_{b}^{2}}{(m_{b}-2)\rho_{b}+1}=(1-\rho_{b})\,\frac{(m_{b}-1)\rho_{b}+1}{(m_{b}-2)\rho_{b}+1},$
+and
+$\Sigma_{(b);-j,-j}^{-1}\Sigma_{(b);-j,j}=\frac{\rho_{b}}{(m_{b}-2)\rho_{b}+1}\,\boldsymbol{1}_{m_{b}-1}.$
+Appendix F Extended numerical results
 Figure 13 shows the extended results of $\widehat{\textnormal{FDR}}$ along with estimated one-standard-error bars in Section 4.1 and Figure 14 shows the extended results for Section 4.2.
@@
 Figure 14: Extended results of $\widehat{\textnormal{FDR}}$ along with estimated one-standard-error bars in Section 4.2. From top-left to bottom-right are stimulation settings 1 to 14.
-Appendix F Running time of $\widehat{\textnormal{FDR}}$ in simulation problems
+Appendix G Running time of $\widehat{\textnormal{FDR}}$ in simulation problems
 Figure 15 shows a box plot of the running time of evaluating $\widehat{\textnormal{FDR}}$ (without estimating its s.e.